### v1

# Multiple Choice Learning of Low Rank Adapters for Language Modeling

Abstract
We propose LoRA-MCL, a training scheme that extends next-token prediction in language models with a method designed to decode diverse, plausible sentence continuations at inference time. Traditional language modeling is an intrinsically ill-posed problem: given a context, multiple futures may be equally plausible. Our approach leverages Multiple Choice Learning (MCL) and the Winner-Takes-All (WTA) loss to efficiently handle ambiguity through Low-Rank Adaptation (LoRA). We provide a theoretical interpretation of applying Multiple Choice Learning to Language Modeling, assuming the data is generated from a mixture of distributions. To illustrate the proposed approach, we use data sampled from mixtures of Markov chains. We then demonstrate with extensive experiments on real-world visual and audio captioning tasks that our method achieves high diversity and relevance in generated outputs.
**footnotetext: Equal Contribution.
1 Introduction
Predicting what a person will say next or describing the content of an audio or visual scene with text is difficult, if not impossible, to do with perfect accuracy. When the context is not informative enough, external factors may lead to different scenarios or modes of plausible text continuations yang2018breaking ; dieng2020topic ; mei2022diverse . In such ambiguous tasks, the conditional distribution over the space of output sentences given the input context may be multi-modal due to the underlying inherent uncertainty malinin2020uncertainty .
Initially proposed for text processing, transformer-based language models have quickly become a general framework for modeling streams of tokens, which can also represent, for instance, images and audio signals touvron2023llama ; chu2024qwen2 ; liu2023visual . Such models are trained as next-token predictors and allow, by nature, for addressing this uncertainty, by generating plausible output sentences through a wide body of maximum a posteriori (MAP) or sampling-based decoding approaches welleckdecoding . While sampling allows exploration and diversity, it may lead to unreliable responses and requires truncation to avoid unexpected answers, partly due to the overestimation of low-probability tokens zhang2021trading ; hewitt2022truncation . When seeking reliable and expected answers, maximum a posteriori estimation techniques, like Beam Search beamSearch , look for sentences that maximize the model’s likelihood. However, these alternatives have drawbacks as they may lack diversity, be prone to repetition loops keskar2019ctrl , and may sound unnatural holtzmancurious . Some approaches, e.g., Diverse Beam Search (DBS) vijayakumar2018diverse , were therefore proposed to artificially increase the diversity at inference, e.g., through a diversity penalty parameter $\lambda$, to find a tradeoff between generation quality and sample diversity sucontrastive . In contrast with these methods, our approach aims to predict diverse sentences reflecting the ambiguity of the input context.
Multiple Choice Learning (MCL) guzman2012multiple ; lee2016stochastic has emerged as a paradigm for addressing ambiguous tasks. It generally consists of a network with a shared backbone and multiple output heads. During training, it utilizes the winner-takes-all loss for adaptively updating the head that performs the best for each example. This is a competitive training scheme that specializes each model to subsets of the conditional output distribution rupprecht2017learning ; letzelter2024winner . In this paper, we propose incorporating this idea for language model finetuning, leveraging multiple Low-Rank Adapters hu2022lora instead of multiple heads, which may be impractical due to computation requirements and architectural constraints. Our method natively generates diverse and plausible sequences in a single forward pass, aiming to best approximate the conditional output distribution.
Our contributions are as follows:
We propose a new paradigm that adapts MCL for token sequence modeling, with a method, LoRA-MCL, that is particularly suited for efficient finetuning of language models.
We provide a theoretical analysis of our approach, which applies MCL to language modeling. Assuming the sequences are sampled from a mixture of distributions, we explain in particular why LoRA-MCL should enable capturing the modes of the data distribution. We then consider the case of Markov chains, for which we validate our claims with a well-designed toy example.
We undertake an extensive experimental study to validate our approach, on audio and machine vision captioning tasks, demonstrating its wide applicability in challenging real-world scenarios and showing an excellent diversity–quality trade-off. The code will be made publicly available.
2 Problem setup
Let $x\triangleq(x_{t})_{t=1}^{T}\in\mathcal{V}^{T}$ be a sequence of $T$ tokens belonging to a finite vocabulary $\mathcal{V}=\{1,\dots,|\mathcal{V}|\}$, and $c\triangleq(c_{t})_{t=1}^{\tau}\in(\mathbb{R}^{d})^{\tau}$ be a sequence of $\tau$ context vector embeddings of dimension $d$. Language modeling aims at learning the law $p(x\,|\,c)=\prod_{t=1}^{T}p(x_{t}\,|\,x_{<t},c)$ using a model $p_{\theta}$ with parameters $\theta$, by minimizing the following negative log-likelihood loss, which is equivalent to the maximum likelihood estimation (MLE):
$\mathcal{L}(\theta)=-\mathbb{E}_{c,x}[\log p_{\theta}(x\,|\,c)]=\mathbb{E}_{c,x}\Big{[}-\sum_{t=1}^{T}\log p_{\theta}(x_{t}\,|\,x_{<t},c)\Big{]}\;,$ (1)
where $x_{<t}$ denotes the sequence of tokens prior to time $t$.
Optimizing (1) is referred to as teacher-forcing williams1989learning , where $p_{\theta}$ is fed with target (instead of predicted) tokens $x_{<t}$ during training sutskever2014sequence . When using a transformer architecture vaswani2017attention , this is implemented with causal attention modules, which allow for computing the conditional distributions in parallel through all time steps within a single forward pass.
During inference, decoding methods welleckdecoding proceed to generating sequences $\hat{x}$ from the trained model $p_{\theta}$ in an auto-regressive fashion. First, they start with a conditional distribution for the first token from the context: $p_{\theta}(x_{1}\,|\,c)$, which allows selecting $\hat{x}_{1}$. Then, for $t\geq 2$ they predict $p_{\theta}(x_{t}\,|\,\hat{x}_{<t},c)$, and select $\hat{x}_{t}$, until reaching either the sequence length limit or an end-of-sentence (EOS) token. The choice of the decoding method used to generate $K$ candidate sequences $\hat{x}^{1},\dots,\hat{x}^{K}$ depends on the purpose of the task, but the general goal is $(i)$ to get highly likely sentences, i.e., ones that maximize $p_{\theta}(\hat{x})$; and $(ii)$ to get sufficiently diverse sentences, as can be measured by $n$-gram similarity ippolito2019comparison .
Although this is a widely adopted paradigm, we show next how this training and decoding pipeline can be improved: instead of artificially generating diversity at inference time, we aim at learning to predict sequences $\hat{x}^{1},\dots,\hat{x}^{K}$ that cover well the modes of the target distribution $p(x\,|\,c)$.
3 Methodology
3.1 Motivation
In language modeling, topic models nigam2000text ; blei2003latent ; dieng2020topic are data-generating processes in which the ground-truth probability distribution of sequences is modeled as a mixture of latent components or topics. For example, the sequence “I am eating …” may have multiple plausible continuations, but the likelihood of each depends heavily on contextual factors such as the speaker’s location, which influences their culinary habits. Each location (or context) can thus be associated with a distinct word distribution. In topic models, data generation proceeds by first sampling a topic $z\in\mathcal{Z}$ for each sentence (usually referred to as a document in the literature), and then sampling words (or $n$-grams) from the distribution associated with that topic.
With this in mind, MLE in (1) may not be suitable yang2018breaking . While MLE is effective for estimating the overall distribution $p(x)$, it does not capture the individual components when $p(x)$ is expressed as a mixture, i.e., $p(x)=\sum_{k}p(z_{k})p(x\,|\,z_{k})$. In such cases, MLE tends to model the aggregate rather than distinguish topic-specific distributions $p(x\,|\,z_{k})$.
3.2 Applying Multiple Choice Learning to language modeling
Our approach is inspired by the multiple choice learning (MCL) literature guzman2012multiple ; lee2016stochastic . We propose the following training scheme, intending to enable the recovery of the different topics $z_{k}$. Instead of a single model, we consider a set of models $(\theta_{1},\dots,\theta_{K})$. Then the objective (1) is replaced by one consisting of iterating between the following two steps:
  1. 1.
    For each training sample $(c,x)$ in the batch $\mathcal{B}$: Compute $p(x\,|\,c;\theta_{k})$ for $k\in\{1,\dots,K\}$, and choose the best model $k^{\star}(x,c)=\mathrm{argmax}_{k}\;p(x\,|\,c;\theta_{k})$.
  2. 2.
    Compute the winner-takes-all (WTA) loss as:
    $\mathcal{L}^{\mathrm{WTA}}(\theta_{1},\dots,\theta_{K})=-\mathbb{E}_{c,x}\Big{[}\max_{k=1,\dots,K}\log p(x\,|\,c;\theta_{k})\Big{]}\,,$ (2)
    where $\log p(x\,|\,c;\theta_{k})=\sum_{t=1}^{T}\log p(x_{t}\,|\,x_{<t},c;\theta_{k})$, and perform an optimization step.
In practice when using a finite batch size, (2) is approximated with
$\mathcal{L}^{\mathrm{WTA}}(\theta_{1},\dots,\theta_{K})\simeq-\sum_{k=1}^{K}\sum_{\{(x,c)\in\mathcal{B}\;\,|\,\;k^{\star}(x,c)=k\}}\log p(x\,|\,c;\theta_{k})\;,$ (3)
where each hypothesis $k$ is trained only on the subset of examples in which it is the winner, according to the assignment $k^{\star}(x,c)$.
This training procedure, similar to a hard-EM style optimization min2019discrete ; wen2023an , is a competitive training scheme that encourages the different models to explore different areas of the data distribution. However, it is subject to two main issues: First, using $K$ models instead of a single one drastically increases the training time and memory cost, which may be intractable for large language models. Second, the optimization may be subject to collapse, where the same models are chosen as winners through the iterations, leaving the other models untrained. In the next section, we describe how we solve these issues with our approach LoRA-MCL.
3.3 LoRA-MCL method
Multiple choice learning typically alleviates the high training cost issue of $K$ models by training a single model with several heads lee2016stochastic ; lee2017confident . However, we argue that such an approach is not well-suited for fine-tuning language models. First, heads of most language models are quite large (for example in Qwen2-Audio chu2024qwen2 the lm_head has $d\times|\mathcal{V}|=4096\times 156032\simeq 640\mathrm{M}$ params), and standard MCL would not scale easily with the number of heads. Second, the initialization of the heads poses several challenges. Initializing each head with the parameters of the head of the pretrained model requires special care, as the collapse of the predictions is likely given the very similar hypotheses (same parameters). A complete re-initialization of the heads is detrimental to performance, as numerous training iterations would be necessary to reach the same level of knowledge as in the pretrained head. For these reasons, we consider a Low Rank Adapter (LoRA) approach hu2022lora due to its excellent trade-off between performance and computational requirements, as well as its wide adoption in the context of large model fine-tuning.
Let $\theta$ be the parameters of the pretrained base model. At each layer $\ell$ where LoRA is enabled, we use a family of adapters $(A^{k}_{\ell},B^{k}_{\ell})\in\mathbb{R}^{d\times r}\!\times\!\mathbb{R}^{r\times d}$ for $k\in\{1,\dots,K\}$. Let
$\theta_{k}=\theta\cup\left\{(A_{\ell}^{k},B_{\ell}^{k})\;|\;\ell=1,\dots,L\right\}\;,$ (4)
be the set of parameters that are involved in hypothesis $k$, with $L$ being the total number of layers where LoRA is used.
Training in the WTA fashion involves computing $p(x\,|\,c;\theta_{k})$ for $k=1,\dots,K$. To avoid collapse into a single head, we may use a (relaxed) winner-takes-all loss as a training objective:
$\mathcal{L}^{\mathrm{WTA}}(\theta)=-\mathbb{E}_{c,x}\Big{[}\sum_{k=1}^{K}q_{k}\log p(x\,|\,c;\theta_{k})\Big{]}\,,$ (5)
where $\{q_{k}\}$ is a set of coefficients that sum to $1$ and that give more weight to the winner index $q_{k^{\star}}$ than to the others, but still provides gradient weight to non-winner heads in order to avoid collapse. In practice, we use $q_{k^{\star}}=1-\varepsilon$ and $q_{k}=\frac{\varepsilon}{K-1}$ for $k\neq k^{\star}$ as in rupprecht2017learning .
3.4 Accelerating LoRA-MCL training with parallelization over the hypotheses
A naive implementation of MCL for winner selection, as in Section 3.2, may require a loop over the $K$ hypotheses in the batch to determine the winner associated with each index of the batch, which would drastically slow down training.
To alleviate this issue, we propose the following methodology. Let $\mathbf{x}\in\mathbb{R}^{b\times\mathcal{T}\times d}$ denote the input of the transformer architecture, where $b$ is the batch size, $\mathcal{T}\triangleq T+\tau$ is the total sequence length, and $d$ is the number of features. We duplicate $\mathbf{x}$, $K$ times along the batch size dimension to get $\mathbf{x}^{\prime}=(\mathbf{x}_{1},\dots,\mathbf{x}_{K})\in\mathbb{R}^{bK\times\mathcal{T}\times d}$.
LoRA units at layer $\ell$ are then computed as:
$\left[\begin{array}[]{c}\mathbf{x}_{1}\\ \mathbf{x}_{2}\\ \vdots\\ \mathbf{x}_{K}\end{array}\right]\leftarrow\left[\begin{array}[]{cccc}B^{1}_{\ell}A^{1}_{\ell}&0&0&0\\ 0&B^{2}_{\ell}A^{2}_{\ell}&0&0\\ \vdots&\vdots&\ddots&\vdots\\ 0&0&0&B^{K}_{\ell}A^{K}_{\ell}\end{array}\right]\left[\begin{array}[]{c}\mathbf{x}_{1}\\ \mathbf{x}_{2}\\ \vdots\\ \mathbf{x}_{K}\end{array}\right]+\left[\begin{array}[]{c}f_{\theta}^{\ell}(\mathbf{x}_{1})\\ f_{\theta}^{\ell}(\mathbf{x}_{2})\\ \vdots\\ f_{\theta}^{\ell}(\mathbf{x}_{K})\end{array}\right]\,,$ (6)
where $f_{\theta}^{\ell}$ is the base model (whose parameters remain frozen during training). In practice, this computation is conducted using a Group Convolution operation.111 This can be done by first reshaping $\mathbf{x}^{\prime}$ to shape $(b\times\mathcal{T}\times Kd)$ and applying a nn.Conv1d($Kd$, $Kd$, kernel_size=1, groups=$K$) following the PyTorch layer implementation, reshaping back $\mathbf{x}^{\prime}$ to shape $(Kb\times\mathcal{T}\times d)$ before adding the base model output, and then repeating at the next LoRA unit. While this duplication virtually multiplies the batch size by $K$, the memory overhead remains manageable, assuming that $r\ll d$. The next section is dedicated to the inner properties of the proposed approach.
4 Theoretical Analysis
We justify the use of MCL in the context of language modeling by assuming that the sequence distribution can be expressed as a mixture. We contrast this formulation with standard maximum likelihood estimation (1). In particular, Section 4.1 establishes a connection between our approach and the Expectation-Maximization (EM) algorithm and derives lower and upper bounds on the optimal achievable test loss. We then apply this analysis to the case of Markov chains in Section 4.2, where the dynamic of the method is simulated in a controlled setting in Section 4.3.
4.1 Training dynamics and optimality conditions
For the next-token prediction loss (1), one can show that: $\mathrm{min}_{\theta}\;\mathcal{L}(\theta)=\mathcal{H}(x\,|\,c)$, where $\mathcal{H}(x\,|\,c)\triangleq-\mathbb{E}_{c,x}[\mathrm{log}\;p(x\,|\,c)]$ is the entropy of $p$ mackay2003information . Following the rationale of Section 3.1, let us now assume the data distribution can be written as a mixture. In the case of the WTA loss under this setup, we have the following proposition.
Proposition 1.
Assume the data-generating process is $p(x\,|\,c)=\sum_{k=1}^{K}p(z_{k}\,|\,c)\;p(x\,|\,z_{k},c)$ (Asm. 2) with the number of modes equal to the number of hypotheses. Also assume perfect model expressiveness (Asm. 1), and large enough batch size to approximate the true risk well (Asm. 3). Then:
  • (i)
    The Winner-Takes-All two-step optimization in LoRA-MCL acts as a conditional form of the hard-EM algorithm.
  • (ii)
    Under Assumption 4, and assuming (with one permutation) that $p(x\,|\,z_{k},c)=p(x\,|\,c;\theta_{k})$ for each $k$, $\mathcal{L}^{\mathrm{WTA}}(\theta)=-\mathbb{E}_{x,c}\left[\underset{k=1,\dots,K}{\mathrm{max}}\mathrm{log}\,p(x\,|\,c,z_{k})\right]$. In this case, we also have:
    $\mathcal{L}^{\mathrm{WTA}}(\theta)=\mathcal{H}(x\,|\,c,z)\triangleq\mathbb{E}_{c}\left[\sum_{k=1}^{K}p(z_{k}\,|\,c)\mathcal{H}\big{(}x\,|\,c,z_{k}\big{)}\right]\,,$ (7)
    where $\mathcal{H}(x\,|\,c,z)$ is the conditional entropy given the random variable $z$.
  • (iii)
    We have the following inequalities:
    $\mathrm{min}_{\theta}\,\mathcal{L}(\theta)-\mathrm{log}\,K\overset{(a)}{\leq}\mathrm{min}_{\theta}\,\mathcal{L}^{\mathrm{WTA}}(\theta)\overset{(b)}{\leq}\mathcal{H}(x\,|\,c,z)\overset{(c)}{\leq}\mathrm{min}_{\theta}\,\mathcal{L}(\theta)\;,$ (8)
    where $\mathrm{min}_{\theta}\;\mathcal{L}(\theta)=\mathcal{H}(x\,|\,c)$.
Proof. See Appendix B.∎
$(i)$ in Proposition 1 describes the relationship between LoRA-MCL and the hard-EM algorithm. $(ii)$ provides an expression for the WTA loss as a conditional entropy, $\mathcal{H}(x\,|\,c,z)$, under the assumption of a perfect matching between the hypotheses and the modes. Proposition 1 also establishes in $(iii)$ both a lower and an upper bound on the optimal achievable loss for LoRA-MCL, given in $(a)$ and $(b)$, respectively. Note that the gap between these bounds is $\mathcal{H}(x\,|\,c,z)-\mathrm{min}_{\theta}\,\mathcal{L}(\theta)+\mathrm{log}\,K=-\mathbb{E}_{c}[\mathcal{I}(x,z\,|\,c)]+\mathrm{log}\,K$ where $\mathcal{I}$ denotes the mutual information. Finally, $(c)$ shows in particular that $\mathrm{min}_{\theta}\,\mathcal{L}^{\mathrm{WTA}}(\theta)\leq\mathrm{min}_{\theta}\,\mathcal{L}(\theta)$.
4.2 Case of Markov chains
To make the analysis more concrete, we consider the case where the sequence of tokens is generated from Markov chains. Formally, given a finite state space $\mathcal{V}$, a homogeneous Markov chain is defined by an initial distribution $\pi$ over $\mathcal{V}$ and a transition matrix $P\in[0,1]^{\mathcal{V}\times\mathcal{V}}$. A sequence $(x_{1},\dots,x_{T})$ is sampled from the Markov chain if $x_{1}\sim\pi$ and, for each $t\geq 1$, the conditional distribution of $x_{t+1}$ given $x_{t}$ is $x_{t+1}\,|\,x_{t}\sim P_{x_{t},\cdot}$, where $P_{i,j}=\mathbb{P}(x_{t+1}=j\,|\,x_{t}=i)$. In the following, we ignore the initial warm-up phase and we assume that $\pi$ is the stationary distribution of $P$, i.e., such that $\pi P=\pi$. In this case, we will denote $x\sim\mathrm{MC}(P)$.
While the study of the training dynamics of transformers on Markov Chain data has been investigated in previous works edelman2024evolution ; rajaraman2024analysis ; makkuva2024attention ; zekri2024large , our setup instead considers a mixture of Markov chains gupta2016mixtures ; kausik2023learning . Assuming a uniform mixture, the data generating process is $x\sim\frac{1}{K}\sum_{k=1}^{K}\mathrm{MC}(P_{k})$, which is equivalent to:
$z_{k}\sim\mathcal{U}(1,\dots,K),\quad x\,|\,z_{k}\sim\mathrm{MC}(P_{k})\;.$ (9)
When training a language model on such sequences, we have the following Corollary from Prop. 1, where the context $c$ is ignored for simplicity.
Corollary 1.
As per Assumption 5, let us assume that the data-generating process writes as a uniform mixture of Markov chains of order $n=1$. We denote as $\hat{P}(\theta)\triangleq(p(x_{t+1}=j\,|\,x_{t}=i))_{i,j}$ the predicted transition matrix when using a language model with parameters $\theta$. Under the same assumptions that in Proposition 1, we have that:
  • (i)
    Whenever the maximum likelihood estimator trained with next-token-prediction (12) reaches its optimal loss $\mathcal{L}(\theta)$, we have
    $\hat{P}(\theta)_{i,j}=\frac{1}{\sum_{s=1}^{K}(\pi_{s})_{i}}\sum_{k=1}^{K}(\pi_{k})_{i}(P_{k})_{i,j}\;,$ (10)
    where $\pi_{k}\in[0,1]^{\mathcal{V}}$ is the stationary distribution of $P_{k}$.
  • (ii)
    The inequality (8) holds in this context, where the conditional entropy $\mathcal{H}(x\,|\,z)$ can be computed by weighting the entropy rate of each Markov Chain:
    $\mathcal{H}(x\,|\,z)=-{\color[rgb]{0,0,0}T}\sum_{k=1}^{K}\sum_{i=1}^{|\mathcal{V}|}(\pi_{k})_{i}\Big{[}\sum_{j=1}^{|\mathcal{V}|}(P_{k})_{i,j}\;\mathrm{log}(P_{k})_{i,j}\Big{]}\,.$ (11)
Proof. See Appendix C. ∎
The entropy $\mathcal{H}(x)$, which is theoretical optimal loss of the MLE baseline, can be computed either exactly for short sequences, or approximated, e.g., through Monte-Carlo integration.
Note that we considered first-order Markov chains in our analysis. We expect the properties generalize to higher orders ($n>1$) by using transition matrices in the form $P\in[0,1]^{\mathcal{V}^{n+1}}$ where $P_{i_{1},\dots,i_{n+1}}=p(x_{t+1}=i_{n+1}\,|\,x_{t}=i_{n},\dots,x_{t-n+1}=i_{1})$. Please refer to Appendix C for further discussions.
4.3 Illustration with synthetic data
Figure 1: Comparison of LoRA-MCL with vanilla maximum likelihood estimation (MLE). The experimental setup is detailed in Section 4.2. (Left) Validation loss over training steps (averaged across three random seeds) for LoRA-MLE (blue) and LoRA-MCL (red). The theoretical optimal MLE loss, approximated via Monte Carlo sampling, is given by the entropy $\mathcal{H}(x)$. The gray shaded region represents the lower and upper bounds of the theoretical optimal MCL loss, as given by $(a)$ and $(b)$ in (8). (Right) Learned transition matrices (top) are compared with the references (bottom). MLE converges approximately toward the weighted average $\bar{P}$ defined as the right-hand side of (10). In contrast, LoRA-MCL successfully recovers the two underlying modes.
To illustrate our approach, let us evaluate our algorithm on a synthetic dataset, for which results are given in Figure 1. We used $\mathcal{V}=\{1,2\}$, and $(P_{1},P_{2})=(P(p_{1},q_{1}),P(p_{2},q_{2}))$, with $P(p,q)\triangleq\left[\begin{array}[]{cc}1-p&p\\ q&1-q\end{array}\right]\;$ and $p,q\in[0,1]$. We sampled data from a mixture of two Markov chains following (9): we generated a sequence by sampling first $k\sim\mathcal{U}\{1,2\}$ and then $P_{k}$. Once $P_{k}$ was set, we sampled uniformly the initial state, then sampled the Markov chain according to the transition matrix until reaching the maximum sequence length ($T=32$ here).
We considered a GPT-2-like architecture radford2019language ; gpt-neo using local-attention suggested by makkuva2024attention to improve convergence on Markov chain data. We then trained the model with $1$ and $2$ hypotheses, using LoRA adapters as described above, $1$ hypothesis corresponding to vanilla maximum likelihood estimation. Training details are provided in Appendix D. Figures 1 and 4 show both the evolution of the losses along training and the predicted transition matrix from the trained models. We see that the optimum is close to being global in this setup. While LoRA-MCL matches can capture the two modes of the mixture, we see that MLE tends to predict the weighted average of the transition matrices given by (10), which is consistent with Prop. 1.
5 Empirical evaluation
We evaluate LoRA-MCL on realistic datasets and large-scale models for audio and image captioning tasks. Predicting a textual description for images or audio signals is an ill-posed problem: from an input image or audio clip, multiple descriptions may be plausible. We argue this is a particularly relevant real-world case of ill-posed problems, where the conditional output distribution is inherently multi-modal. Our goal is to demonstrate that LoRA-MCL provides a competitive approach for capturing these distribution modes when fine-tuning either audio (Sec. 5.2) or vision-language models (Sec. 5.3). We present hereafter the experimental setup. Please refer to Appendix E.2 for further details.
5.1 Experimental Setup
Datasets. We experimented on both Clotho-V2 font2013freesound ; drossos2020clotho , and AudioCaps gemmeke2017audio ; kim2019audiocaps datasets for the audio captioning task, while we make use of the TextCaps dataset sidorov2020textcaps for a challenging task for vision language models coined as image captioning with reading comprehension. Table 4 of the Appendix describes the datasets sizes.
Architecture and training details with audio data. We used the instructed version of Qwen2-Audio chu2024qwen2 as the base model, which has ${\sim}8.4$ billion parameters and a vocabulary size $|\mathcal{V}|=156{,}032$. We used LoRA adapters applied to the $Q,K,V$ linear projections of the attention modules, and the upside and downside projections of the feedforward blocks, for all the layers of the transformers. We used a rank $r$ and scaling factor $\alpha$, with $r=\alpha=8$ unless otherwise stated. We trained for 1 epoch on AudioCaps and 10 epochs on Clotho with a validation loss computed at the end of each epoch.
Architecture and training details with visual data. We used LLaVA 1.6 liu2023visual with Vicuna-7B zheng2023judging for the LLM, as the base model which features ${\sim}7.1$ billion parameters and a vocabulary size $|\mathcal{V}|=32{,}000$. We used the official codebase for the implementation. Following standard practices zhou2024empirical , we applied LoRA adapters only for the LLM decoder. The adapters were applied to $Q,K,V$, upside and downside projections as in Qwen2-Audio, and we used $r=\frac{\alpha}{4}=8$ unless otherwise stated. Training was done over 1 epoch (without validation data), and the validation set of TextCaps was used for evaluation.
Metrics. We evaluate the approach using various quality metrics. First, we compute the test-loss (or perplexity), which evaluates the likelihood of the reference captions under the trained model. Second, we use natural language generation quality metrics (BLEU-$n$, ROUGE, METEOR) papineni2002bleu ; lin2004rouge ; banerjee2005meteor based on $n$-gram matching between predicted and target sentences. We also leverage the latter proposed CIDEr, SPICE and SPIDEr vedantam2015cider ; anderson2016spice ; liu2017improved metrics, which show higher correlation with human evaluation for captioning tasks. Finally, we use Sentence-BERT (sBERT) reimers2019sentence ; zhou2022can for embedding comparison. We consider sentence-based oracle evaluation for these metrics (see e.g., (lee2016stochastic, , Sec. 4.3) or labbe2022my ). To evaluate the diversity of the generated captions, we consider: Div-$n$, which compares the number of distinct $n$-grams against the number of $n$-grams for a set of generated captions as in mei2022diverse and Self-BLEU (mBLEU-4), which computes BLEU similarity scores between generated captions zhu2018texygen .
Baselines. To validate our approach, we compare it against a maximum likelihood baseline (LoRA-MLE) trained using (1), under the same conditions as the ones considered for our multi-hypothesis model. Specifically, both models use the same LoRA configuration, the same number of trainable parameters (the LoRA rank for the baseline is $K\times$ larger to this end), and the same number of iterations. At inference time, for each decoding method applied to the baseline that returns $K$ sentences, we decode the same number of candidates with LoRA-MCL. When evaluating Maximum A Posteriori methods –greedy, beam search beamSearch and diverse beam search vijayakumar2018diverse –, we ensure a consistent computational budget by aligning the number of forward passes. For example, if the MLE baseline uses a beam size of $B$, our model uses a beam size of $\frac{B}{K}$ per hypothesis.
5.2 Audio captioning
Figure 2: Quality vs. Diversity on Audio Captioning with $5$ candidates. Quality measured by SPIDEr ()($\uparrow$)( ↑ ) and Diversity by mBLEU-4 ()($\downarrow$)( ↓ ). Marker shape stands for the method, and size is proportional to the number of forwards performed per example at inference time. Note that LoRA-MLE was trained with two rank values, $r=8$ and $r=8K$, for fair comparison in terms of parameter count. Color corresponds to the decoding method (beam search or diverse beam search). Values of $\varepsilon$ and $\lambda$ for LoRA-MCL and diverse beam search, resp., are indicated in the plot, with the color shade proportional to the corresponding parameter value to better distinguish the markers.
Quality vs. diversity trade-off. Quality–diversity performance is shown in Figure 2. While a well-chosen value of $\lambda$ allows Diverse Beam Search applied to LoRA-MLE to be competitive, LoRA-MCL (circles) achieves the best trade-off between quality and diversity, appearing in the top-left corner of the plot. Although increasing the beam size generally improves performance for standard beam search, we observe that increasing the beam size within each group negatively impacts Diverse Beam Search. In theory, setting $\varepsilon=0$ should maximize both quality and diversity. However, using a small positive value of $\varepsilon$ ensures all heads are trained, albeit at the cost of reduced diversity and potentially quality when $\varepsilon$ is too large. Full results are given in Tables 5 and 6 of the Appendix.
Effect of the number of hypotheses. We include in Figures 6 and 5 (given in the appendix) the same plot as in Figure 2, but using 3 and 4 candidates, respectively. These results further reinforce the findings obtained with 5 hypotheses. In Table 5.2, we report the evolution of the test negative log-likelihood (NLL) loss as a function of the number of hypotheses, showing a monotonically decreasing trend. This provides additional evidence for Proposition 1, and shows that LoRA-MCL improves its coverage of the data distribution modes as $K$ increases.
Table 1: Test Loss ()($\downarrow$)( ↓ ) as a function of the number of hypotheses $K$. LoRA-MCL is trained with $\varepsilon=0.05$ and $r=8$.
Training $K$ AudioCaps Clotho
LoRA-MLE ($r=8$) 1 2.152 2.811
LoRA-MLE ($r=40$) 1 2.124 2.902
LoRA-MCL 2 2.071 2.700
LoRA-MCL 3 2.021 2.665
LoRA-MCL 5 1.927 2.629
LoRA-MCL 8 1.869 2.592
Table 2: Quality and Diversity Evaluation on TextCaps with $3$ candidates. Best scores are in bold, second-best are underlined. For each of the presented metrics, higher is better ()($\uparrow$)( ↑ ) except for mBLEU-4 ()($\downarrow$)( ↓ ). LoRA-MCL is trained with $\varepsilon=0.1$, $r=8$ and $\alpha=32$. LoRA-MLE is trained with $r=24$ and $\alpha=4\times r=96$, ensuring the same dynamics and number of parameters across models. For completeness, we also trained LoRA-MLE with $r=8$ and $\alpha=32$ in the rows marked with .
Training Decoding Beam mBLEU-4 BLEU-4 METEOR sBERT CIDEr-D SPICE SPIDEr
LoRA-MLE Beam Search 3 0.688 0.318 0.315 0.670 1.517 0.244 0.873
LoRA-MLE Beam Search 6 0.786 0.338 0.326 0.671 1.557 0.246 0.895
LoRA-MLE DBS ($\lambda=0.8$) 3 0.437 0.349 0.327 0.686 1.590 0.251 0.909
LoRA-MLE DBS ($\lambda=1.0$) 3 0.416 0.348 0.326 0.685 1.586 0.250 0.906
LoRA-MLE DBS ($\lambda=1.0$) 3 0.425 0.357 0.327 0.686 1.601 0.252 0.915
LoRA-MLE DBS ($\lambda=0.8$) 6 0.671 0.341 0.328 0.681 1.573 0.251 0.903
LoRA-MLE DBS ($\lambda=1.0$) 6 0.666 0.340 0.328 0.680 1.577 0.250 0.904
LoRA-MLE DBS ($\lambda=1.0$) 6 0.671 0.341 0.328 0.680 1.584 0.251 0.908
LoRA-MCL Beam Search 1 0.520 0.344 0.330 0.690 1.674 0.255 0.955
LoRA-MCL Beam Search 2 0.490 0.360 0.333 0.687 1.627 0.258 0.932
5.3 Image description with reading comprehension
5.3.1 Quality and Diversity Evaluation
Similarly to Figure 2, Table 2 confirms the trends observed in the audio captioning task. At an equal number of forward passes, LoRA-MCL outperforms LoRA-MLE, even using DBS decoding with a $\lambda$ diversity parameter specifically optimized for the task (SPIDEr of $0.955$ against $0.915$). However, we noticed that the DBS decoding of LoRA-MLE tends to generate more diverse outputs than the greedy decoding of LoRA-MCL (mBLEU of $0.416$ against $0.520$). Interestingly, increasing the number of beams in each group decreases diversity and does not improve the performance of LoRA-MLE, while with LoRA-MCL, increasing the beam size (from 1 to 2) not only slightly increases the performance but also improves the diversity of the generations, indicating promising further inference-time scaling. Additional results are provided in Appendix (Table 9).
5.3.2 Observing hypothesis specialization in bilingual image description
To highlight the behavior of LoRA-MCL in a realistic case where one can control the modes of the data-generating process, we simulated an artificial bi-modal distribution of the dataset, similarly to the setup of the toy experiment of Section 4.3. We did so by translating half of the captions of the data from English to French (using T5-small T5 ), while keeping the prompts in English.
We trained and evaluated a 2-hypothesis model in the LoRA-MCL setup, as well as a baseline trained in LoRA-MLE. We observed a clear specialization of each hypothesis towards a given language (i.e., one hypothesis learned French and the other English): at test-time, the winning head is the first one in $\sim 89\%$ of the French captions and the second one in $\sim 97\%$ of the English captions. Table 3 shows performance and diversity scores of both LoRA-MLE and LoRA-MCL on our synthetic test set. LoRA-MCL is decoded greedily while LoRA-MLE uses DBS with the $\lambda$ value maximizing its performance (0.8). Although the overall performances are similar, LoRA-MCL generates significantly more diverse captions (mBLEU-4 of 0.027 and 0.029 against 0.138 and 0.126 on the French and English subparts, respectively), and outperforms LoRA-MLE in the French part of the test set (SPIDEr of 0.464 against 0.411).
These results relate to the synthetic data experiments of Sec. 4.2: in the presence of two modes, LoRA-MLE learns an average of them. Here, we believe this is more shifted toward English as the original LLaVA model performs better in English than in French. In contrast, LoRA-MCL is expected to learn the two modes separately. We believe this is what explains the gain in performance in French, at the cost of a slight reduction in English performance.
Fig. 3 qualitatively illustrates the behavior of the models: LoRA-MLE learns a weighted average of the two modes, shifted towards English, and it is sometimes not able to output French captions within the 2 candidates as in the book example. We found LoRA-MLE to be particularly prone to errors when outputting French sentences: for instance, in the second generation of examples with an image of a bottle, LoRA-MLE enters a repetition loop. On the other hand, LoRA-MCL is less affected by those artifacts on French sentences, and successfully captures the two modes of the distribution, benefiting from the specialization of each hypothesis.
Table 3: SPIDEr ()($\uparrow$)( ↑ ) & mBLEU-4 ()($\downarrow$)( ↓ ) on different parts of synthetic test set.
Test subset Training SPIDEr mBLEU-4
French LoRA-MLE 0.411 0.138
LoRA-MCL 0.464 0.027
English LoRA-MLE 0.756 0.126
LoRA-MCL 0.722 0.029
LoRA-MLE.
{A bottle of Cerveza is on a table.}
{Une bouteille de vin de cidre de cidre de cidre […]}
LoRA-MCL.
{A bottle of beer with a label that says "Sel Maguet"}
{Une bouteille de vin est étiquetée avec le mot « Maguay ».}
LoRA-MLE.
{A book titled Papa Told Me is being held by a person.}
{A book called Papa told me is being held by a person.}
Lora-MCL.
{A book titled Papa Told Me is being held by a person}
{Un livre papier intitulé Papa Told Me.}
Figure 3: Observing specialization in bilingual image description. Quantitative (Left) and Qualitative (Right) analysis for LoRA-MLE and LoRA-MCL in the setup of Section 5.3.2.
6 Related Work
Multiple Choice Learning to predict diverse and plausible outputs. Multiple Choice Learning guzman2012multiple ; lee2016stochastic is a training paradigm that minimizes the oracle or winner-takes-all loss across a set of models, encouraging specialization rupprecht2017learning ; letzelter2024winner . While it has been noted that the collapse issue needs to be addressed with care rupprecht2017learning ; amcl , MCL has demonstrated broad applicability across machine learning tasks lee2017confident ; seo2020trajectory ; garcia2021distillation , typically using a shared backbone and multiple projection heads. However, using multiple heads may be impractical in large language models. Mixture-of-Experts jacobs1991adaptive ; shazeer2017outrageously can be an effective strategy to manage computational costs in this case. However, as primarily thought for model scalability and not diversity, only a small subset of experts are active at each forward pass via routing, making them incompatible with MCL, while still suffering from knowledge redundancy in the experts jiang2024mixtral . To the best of our knowledge, we are the first to adapt MCL to the setup of language modeling with next-token prediction, specifically, using internal models instantiated via multiple LoRA hu2022lora .
Generating Multiple Outputs with Language Models. Language models are commonly trained via next-token prediction, framed as Maximum Likelihood Estimation. This is arguably the most popular method for training large-scale language models shlegerislanguage ; radford2019language ; touvron2023llama , including those that take audio or images as input chu2024qwen2 ; liu2023visual , with much of its success attributed to tokenization makkuva2024attention ; rajaraman2024analysis . Generating diverse and plausible sequences at inference remains challenging: $(i)$ Sampling methods holtzmancurious ; fan2018hierarchical ; meister2023locally may be unreliable depending on the chosen parameters; $(ii)$ Exact MAP decoding is intractable due to the exponential search space eikema2020map ; $(iii)$ Strategies like Beam Search often yield repetitive or overly coherent outputs fan2018hierarchical ; holtzmancurious ; keskar2019ctrl . Diverse Beam Search mei2022diverse or contrastive decoding sucontrastive inject diversity through test-time parameters (e.g., penalty $\lambda$), but in contrast, our method infers diversity from the input’s inherent ambiguity.
Diversity in Audio and Visual Captioning. Audio drossos2017automated ; mei2021audio ; mei2022automated and image captioning aneja2018convolutional ; herdade2019image ; hossain2019comprehensive have traditionally relied on dataset- and task-specific models trained with Maximum Likelihood Estimation (MLE). A key challenge identified in both domains is the lack of diversity in generated captions, prompting the development of other training objectives specifically designed to address this issue mei2022diverse ; xu2022diversity ; xu2024towards ; zhang2024generating ; wang2019describing ; wang2020diversity ; mahajan2020diverse . These methods often require architectural modifications and are typically designed for models trained from scratch. However, recent advances have shifted the field toward general-purpose multimodal large language models chu2024qwen2 ; liu2023visual , which offer broader capabilities but pose new challenges to address the diversity-quality trade-off. In this work, we demonstrate that LoRA-MCL can be effectively applied at the fine-tuning stage of such models.
7 Conclusion
This paper proposes a new paradigm for training language models to generate diverse plausible predictions. This is done with LoRA-MCL, combining the Multiple Choice Learning paradigm and Low-rank Adaptation. We provide a theoretical analysis showing that when the target sequence is written as a mixture, LoRA-MCL is expected to capture the modes of the data distribution. We validate our analysis with experiments on Markov chains, and on the real-world audio and advanced image captioning tasks. Future work includes a careful interpretation of the concepts learned by each model as a function of the ambiguity of the input, e.g., for uncertainty estimation as an alternative to wang2023lora .
Limitations. In LoRA-MCL, setting $\varepsilon$ too high ensures that all hypotheses receive gradients, but it can also overly homogenize the models, thereby reducing diversity. While MCL performs well across a broad range of $\varepsilon$ values, identifying the optimal setting can be challenging. To do so, we plan to incorporate recent annealing MCL techniques, e.g., amcl , which employ a scheduling mechanism for $\varepsilon$ and may help improve the quality of the solution. Note that recent LoRA variants, such as hayou2024lora+ may also be an orthogonal axis of improvement for our method.
Acknowledgments
We thank Mickael Chen for his valuable advice, as well as Gilles Puy for his help in the implementation of the parallelization over the hypotheses. This work was funded by the French Association for Technological Research (ANRT CIFRE contract 2022-1854), and granted access to the HPC resources of IDRIS (allocation 2024-AD011014345) by GENCI.
Appendix A Notations and setup
In the following, $x\triangleq(x_{t})_{t=1}^{T}\in\mathcal{V}^{T}$ be a sequence of $T$ tokens belonging to a finite vocabulary $\mathcal{V}=\{1,\dots,|\mathcal{V}|\}$, and $c\triangleq(c_{t})_{t=1}^{\tau}\in(\mathbb{R}^{d})^{\tau}$ be a sequence of $\tau$ context of embeddings of dimension $d$. In the following $\mathcal{X}=\mathcal{V}^{T}$ and $\mathcal{C}=(\mathbb{R}^{d})^{\tau}$.
Language modeling aims at learning the law $p(x\,|\,c)=\prod_{t=1}^{T}p(x_{t}\,|\,x_{<t},c)$ using a model $p_{\theta}$ with parameters $\theta\in\Theta$, by maximum likelihood estimation, minimizing the negative log-likelihood loss
$\mathcal{L}(\theta)=-\mathbb{E}_{c,x}[\log p_{\theta}(x\,|\,c)]=\mathbb{E}_{c,x}\left[-\sum_{t=1}^{T}\log p_{\theta}(x_{t}\,|\,x_{<t},c)\right]\;,$ (12)
where $x_{<t}$ denotes the sequence of tokens prior to time $t$. In practice, we assume that $p_{\theta}=s_{\eta}\circ f_{\theta}$, where $f_{\theta}(x_{<t},c)\in\mathbb{R}^{\mathcal{V}}$ are the predicted logits and $s_{\eta}:z\in\mathbb{R}^{\mathcal{V}}\mapsto\left(\frac{\mathrm{exp}(z_{j}/\eta)}{\sum_{q=1}^{|\mathcal{V}|}\mathrm{exp}(z_{q}/\eta)}\right)\in[0,1]^{\mathcal{V}}$ is the $\mathrm{softmax}$ operator with temperature $\eta>0$.
In the following, we make the following assumption.
Assumption 1 (Expressiveness).
In the following, we assume that the model $p_{\theta}$ is perfectly expressive. Formally, let $\mathcal{F}_{\theta}\triangleq\{p_{\theta}:\mathcal{C}\rightarrow\mathcal{P}(\mathcal{X})\}$ be the family of conditional distributions realized by the model, where $\mathcal{P}(\mathcal{X})$ is the set of probability distributions on $\mathcal{X}$. We assume the family $\mathcal{F}_{\theta}$ is perfectly expressive, that is $\mathcal{F}_{\theta}=\{p^{\star}:\mathcal{C}\rightarrow\mathcal{P}(\mathcal{X})\}$.
First, note that we have the Proposition 2, a well-known result that is due to the Gibbs inequality (e.g., [42]), for which we provide a proof for completeness.
Proposition 2 ([42]).
Under Assumption 1, for the next-token prediction loss (12), one can show that
$\mathrm{min}_{\theta}\;\mathcal{L}(\theta)=\mathcal{H}(x\,|\,c)\,,$ (13)
where $\mathcal{H}(x\,|\,c)\triangleq-\mathbb{E}_{c,x}[\mathrm{log}\;p(x\,|\,c)]$ is the entropy of $p(\cdot\,|\,c)$.
Proof.
Let us denote $\mathscr{S}(p)=\{(x,c)\,|\,p(x,c)>0\}$ the support of $p$. We use the convention $p(x,c)\,\mathrm{log}\,p(x,c)=0$ for $(x,c)\in\mathcal{X}\times\mathcal{C}-\mathscr{S}(p)$. Because $\mathrm{log}$ is a concave function, we have under Jensen’s inequality:
$-\int_{\mathscr{S}(p)}\mathrm{log}\left[\frac{p_{\theta}(x\,|\,c)}{p(x,c)}\right]p(x,c)\;\mathrm{d}x\mathrm{d}c\geq-\mathrm{log}\left(\int_{\mathscr{S}(p)}\frac{p_{\theta}(x\,|\,c)}{p(x,c)}p(x,c)\;\mathrm{d}x\mathrm{d}c\right)\geq 0\;.$ (14)
However, because of the convention, the left-hand side of (14) is also equal to the integral over $\mathcal{X}\times\mathcal{C}$. This shows:
$-\int_{\mathcal{X}\times\mathcal{C}}p(x,c)\;\mathrm{log}\;p_{\theta}(x\,|\,c)\;\mathrm{d}x\mathrm{d}c\geq-\int_{\mathcal{X}\times\mathcal{C}}p(x,c)\;\mathrm{log}\;p(x\,|\,c)\;\mathrm{d}x\mathrm{d}c=\mathcal{H}(x\,|\,c)\;,$
where the equality is reached for parameter $\theta$ such that $p_{\theta}=p$, whose existence is guaranteed by Assumption 1. ∎
In the following, we denote the Kullback–Leibler divergence between two distributions $\alpha$ and $\beta$ as $\mathrm{KL}(\alpha\,\|\,\beta)\triangleq\int_{\mathscr{S}(\beta)}\alpha(x)\mathrm{log}\frac{\alpha(x)}{\beta(x)}\mathrm{d}x$, where $\mathscr{S}(\beta)\triangleq\{x\in\mathcal{X}\,|\,\beta(x)>0\}$. Note that we have the equality:
$-\mathbb{E}_{\alpha}[\mathrm{log}\,\beta(x)]=\mathrm{KL}(\alpha\,\|\,\beta)+\mathcal{H}(\alpha)\;,$ (15)
where the left-hand side is usually referred as the Cross-Entropy, and $\mathcal{H}(\alpha)\triangleq-\mathbb{E}_{\alpha}[\mathrm{log}\,\alpha(x)]$ is the entropy of $\alpha$. When the context is clear, we will also write the entropy of a distribution $\alpha$ as $\mathcal{H}(x)$ where $x\sim\alpha$.
Appendix B Proof of Proposition 1
Let us now consider the following assumptions.
Assumption 2 (Mixture of latent processes).
The data-generating process writes in form $p(x\,|\,c)=\sum_{k=1}^{K}p(z_{k}\,|\,c)\;p(x\,|\,z_{k},c)$. The Mixture is said to be uniform if $\forall k,p(z_{k}\,|\,c)=\frac{1}{K}$.
Assumption 3 (Minimization of the true risk).
The batch size is large enough so that the minimization of the empirical risk comes down to minimizing the true risk (12).
Remark 1.
Under Assumptions 2 and 3, the optimal reachable loss by maximum likelihood estimation is $\min_{\theta}\mathcal{L}(\theta)=\mathbb{E}_{c}\left[\mathcal{H}(x\,|\,c)\right]$, where $x\sim\frac{1}{K}\sum_{k=1}^{K}p(x\,|\,z_{k},c)$.
Assumption 4 (Disjoint components).
This assumption states that $p(x\,|\,c,z_{s})=0$ when $p(x\,|\,c,z_{k})>0$, for $s\neq k$.
Proposition 3.
Under Assumptions 1, 2, and 3, we have that:
  • (i)
    The Winner-Takes-All two-step optimization in LoRA-MCL acts as a conditional form of the hard-EM algorithm.
  • (ii)
    Under Assumption 4, and assuming (with one permutation) that $p(x\,|\,z_{k},c)=p(x\,|\,c;\theta_{k})$ for each $k$, $\mathcal{L}^{\mathrm{WTA}}(\theta)=-\mathbb{E}_{x,c}\left[\underset{k=1,\dots,K}{\mathrm{max}}\mathrm{log}\,p(x\,|\,c,z_{k})\right]$. In this case, we also have:
    $\mathcal{L}^{\mathrm{WTA}}(\theta)=\mathcal{H}(x\,|\,c,z)\triangleq\mathbb{E}_{c}\left[\sum_{k=1}^{K}p(z_{k}\,|\,c)\mathcal{H}\big{(}x\,|\,c,z_{k}\big{)}\right]\,,$ (16)
    where $\mathcal{H}(x\,|\,c,z)$ is the conditional entropy given the random variable $z$.
  • (iii)
    We have the following inequalities:
    $\mathrm{min}_{\theta}\,\mathcal{L}(\theta)-\mathrm{log}\,K\overset{(a)}{\leq}\mathrm{min}_{\theta}\,\mathcal{L}^{\mathrm{WTA}}(\theta)\overset{(b)}{\leq}\mathcal{H}(x\,|\,c,z)\overset{(c)}{\leq}\mathrm{min}_{\theta}\,\mathcal{L}(\theta)\;,$ (17)
    where $\mathrm{min}_{\theta}\;\mathcal{L}(\theta)=\mathcal{H}(x\,|\,c)$.
Proof of (i) First, let us remind that the hard-EM consists of fitting a distribution $p_{\theta}(x,z)$ to observed data $x\sim p(x)$ where $z$ are (unknown) hidden variables. The fitting starts from randomly initialized parameters $\theta$ and latent variables $z$. It consists of repeating the following operations at each iteration $t$ until convergence:
  1. 1.
    (Expectation) $z_{k}^{\star}=\mathrm{argmax}_{k}\;p(x,z_{k};\theta^{(t)})$
  2. 2.
    (Maximization) $\theta^{(t+1)}=\mathrm{argmax}_{\theta}\;p(x,z_{k}^{\star};\theta^{(t})$
Let us define:
$D(\theta,q)\triangleq\int_{\mathcal{X}}\sum_{k=1}^{K}q(k\,|\,x)\mathrm{log}\,p(x,z_{k};\theta)\;\mathrm{d}p(x),\leavevmode\nobreak\ D(\theta)\triangleq\int_{\mathcal{X}}\max_{k=1,\dots,K}\;\mathrm{log}\,p(x,z_{k};\theta)\;\mathrm{d}p(x)\;,$ (18)
where $q$ is a discrete distribution over $\{1,\dots,K\}$ with exactly one non-zero component that controls the assignment of each $x$ to a fixed $k^{\star}$. Let us define $q(\theta)$ as the discrete distribution defined so that $q(k\,|\,x;\theta)\triangleq\mathbf{1}[k=\mathrm{argmax}_{s}\;p(x,z_{s};\theta)]$. Note that then $D(\theta)=D\left(\theta,q(\theta)\right)$.
For the vanilla (or soft) EM algorithm, the complete data log-likelihood $\int_{\mathcal{X}}\mathrm{log}\,p(x;\theta)p(x)\mathrm{d}x$ is expected to increase at each iteration $t$. Similarly, for the hard-EM, we have that the $D(\theta)$ increases at each iteration.
Indeed, the expectation step comes down to computing $q(\theta^{(t)})$. For for the Maximization step, we have: $D\left(\theta^{(t+1)},q(\theta^{(t)})\right)\geq D\left(\theta^{(t)},q(\theta^{(t)})\right)$, by definition. At the next expectation step, we have: $D\left(\theta^{(t+1)},q(\theta^{(t+1)})\right)\geq D\left(\theta^{(t+1)},q(\theta^{(t)})\right)$, because $q(\theta^{(t+1)})$ computes the best assignment given the parameters $\theta^{(t+1)}$. This shows that $D(\theta^{(t+1})\geq D(\theta^{(t)})$.
First note that the main difference compared to the vanilla form of the (hard) EM algorithm, is that the goal is to fit here a conditional distribution $p(x\,|\,c)$ given pair $(c,x)\sim p(c,x)$. Furthermore, step 2 performs a gradient step (of the neural network weights) instead of a full maximization.
Note that under Assumption 2 the complete data log-likelihood writes as $\mathrm{log}\,p(x\,|\,c;\theta)=\mathrm{log}\left[\;\sum_{k=1}^{K}p(x,z_{k}\,|\,c;\theta)\right]$. However, the variables $z_{k}$ are not known in practice. Lora-MCL works by analogy with the Hard-EM algorithm, which consists, in the Expectation step, of picking for each pair $(x;c)$, $k^{\star}(x,c)=\mathrm{argmax}_{k}\;p(x\,|\,c;\theta_{k})$. Indeed, under Assumption 3, each training step of Lora-MCL writes as the optimization of
$\mathcal{L}^{\mathrm{WTA}}(\theta)=-\int_{\mathcal{X}\times\mathcal{C}}\max_{k=1,\dots,K}\mathrm{log}\;p(x\,|\,c;\theta_{k})\mathrm{d}p(c,x)\;,$ (19)
which is expected to decrease at each training iteration. Note that because the loss is bounded from below (by 00), the sequence of real numbers $\{\mathcal{L}^{\mathrm{WTA}}(\theta^{(t)})\}_{t\geq 0}$ is therefore expected to converge.
To conclude, we can view $(\theta_{1},\dots,\theta_{K})$ as the parameters involved in the estimation of the modes of the conditional distribution with $p(x\,|\,c)=\sum_{k=1}^{K}p(x\,|\,c;\theta_{k})p(\theta_{k}\,|\,c)$. Note that the current form of the algorithm does not estimate the weight of each mode $p(\theta_{k}\,|\,c)$, further work could include incorporating scoring heads to estimate $p(\theta_{k}\,|\,c)$ each $k$ as in [35]. ∎
We then expect to be able to recover the distributions (with one permutation) $\{p(x\,|\,c,z_{k})\}$ from estimated $\{p(x\,|\,c;\theta_{k})\}$, assuming identifiability of the data generating mixture, which we expect to be made easier if the components are enough separated (See e.g., [59, Par. 2.5] or [47, Sec. 2.2]).
Proof of (ii) Let us assume that (with one permutation) $p(x\,|\,z_{k},c)=p(x\,|\,c;\theta_{k})$ for $k\in\{1,\dots,K\}$. This is possible thanks to Assumption 1. Let us show that (16).
Let us define:
$\mathcal{X}_{k}(c,\theta)\triangleq\Big{\{}x\in\mathcal{X}\,|\,\mathrm{log}\;p(x\,|\,c,\theta_{k})\geq\mathrm{log}\;p(x\,|\,c,\theta_{s})\;\forall s\in\{1,\dots,K\}\Big{\}}\;.$ (20)
In this case, the WTA loss (2) writes as
$\mathcal{L}^{\mathrm{WTA}}(\theta)$ $=-\int_{\mathcal{C}}\sum_{k=1}^{K}\int_{\mathcal{X}_{k}(c,\theta)}\mathrm{log}\;p(x\,|\,c;\theta_{k})\;p(x\,|\,c)\;\mathrm{d}x\;p(c)\mathrm{d}c$
$=-\int_{\mathcal{C}}\sum_{k=1}^{K}\sum_{s=1}^{K}\int_{\mathcal{X}_{k}(c,\theta)}\mathrm{log}\;p(x\,|\,c;z_{k})\;p(x\,|\,c;z_{s})\;p(z_{s}\,|\,c)\;\mathrm{d}x\;p(c)\;\mathrm{d}c$
[…]
$=-\int_{\mathcal{C}}\sum_{k=1}^{K}\mathcal{H}(x\,|\,c;z_{k})\;p(z_{k}\,|\,c)\;p(c)\mathrm{d}x\;\mathrm{d}c\text{\quad as $\int_{\mathcal{X}_{k}(c,\theta)}p(x\,|\,c;z_{k})\mathrm{d}x=1$.}$
$=\mathbb{E}_{c}\left[\sum_{k=1}^{K}p(z_{k}\,|\,c)\mathcal{H}\big{(}x\,|\,c;z_{k}\big{)}\right]\;.$
Proof of (iii) Let us show that:
$\mathrm{min}_{\theta}\,\mathcal{L}(\theta)-\mathrm{log}\,K\overset{(a)}{\leq}\mathrm{min}_{\theta}\,\mathcal{L}^{\mathrm{WTA}}(\theta)\overset{(b)}{\leq}\mathcal{H}(x\,|\,c,z)\overset{(c)}{\leq}\mathrm{min}_{\theta}\,\mathcal{L}(\theta)\;.$
$(a)$: First, we have $\underset{k=1,\dots,K}{\mathrm{max}}\,p(x\,|\,c,z_{k})\leq\sum_{k=1}^{K}p(x\,|\,c,\theta_{k})$. Therefore,
$\mathcal{L}^{\mathrm{WTA}}(\theta)$ $\geq-\mathbb{E}_{x,c}\left[\mathrm{log}\frac{1}{K}\sum_{k=1}^{K}p(x\,|\,c,\theta_{k})\right]-\mathrm{log}\,K$
[…]
$\geq\mathcal{H}\left(x\,|\,c\right)-\mathrm{log}\,K\,.$
Because $\mathrm{min}_{\theta}\,\mathcal{L}(\theta)=\mathcal{H}\left(x\,|\,c\right)$, we have shown that $(a)$ occurs when the KL term vanishes, which is when $p(x\,|\,c)=\frac{1}{K}\sum_{k=1}^{K}p(x\,|\,c,\theta_{k})$.
$(b)$: We have
$\mathcal{L}^{\mathrm{WTA}}(\theta)$ $=-\mathbb{E}_{x}[\underset{k=1,\dots,K}{\mathrm{max}}\mathrm{log}\,p(x\,|\,c,\theta_{k})]$
$=-\mathbb{E}_{z}\mathbb{E}_{x\,|\,z}\left[\mathrm{log}\frac{\underset{k=1,\dots,K}{\mathrm{max}}\,p(x\,|\,c,\theta_{k})}{p(x\,|\,c,z)}\right]\underbrace{-\mathbb{E}_{z}\mathbb{E}_{x\,|\,z}\left[\mathrm{log}\,p(x\,|\,c,z)\right]}_{\mathcal{H}(x\,|\,c,z)}\,.$
Now let us leverage Assumption 1 to choose $\tilde{\theta}_{k}$ such that $p(x\,|\,c,\tilde{\theta}_{k})=p(x\,|\,c,z_{k})$ for each $k\in\{1,\dots,K\}$. In this case, $\mathrm{max}_{k}\,p(x\,|\,c,\tilde{\theta}_{k})\geq p(x\,|\,c,z)$ for each $z\in\{z_{1},\dots,z_{K}\}$, and $-\mathbb{E}_{z}\mathbb{E}_{x\,|\,z}\left[\mathrm{log}\frac{\mathrm{max}_{k}\,p(x\,|\,c,\tilde{\theta}_{k})}{p(x\,|\,c,z)}\right]\leq 0$. Then,
$\mathrm{min}_{\theta}\mathcal{L}^{\mathrm{WTA}}(\theta)\leq\mathcal{L}(\tilde{\theta}_{1},\dots,\tilde{\theta}_{K})\leq\mathcal{H}(x\,|\,c,z)\;,$
which proves $(b)$.
Finally $(c)$ can be directly deduced from the inequality $\mathcal{H}(x\,|\,c,z)\leq\mathcal{H}(x\,|\,c)$.∎
Appendix C Proof of Corollary 1
Let us consider the following assumptions.
Assumption 5 (Markov Chain).
We assume that the data-generating process writes as a uniform mixture of Markov chains of order $n\in\mathbb{N}\setminus\{0\}$, that is, for each $t$ and each $k$, $p(x_{t}\,|\,x_{<t},c,z_{k})=p(x_{t}\,|\,x_{t-1},\dots,x_{t-n},c,z_{k})$.
Corollary 2.
As per Assumption 5, let us assume that the data-generating process writes as a uniform mixture of Markov chains of order $n=1$. We denote as $\hat{P}(\theta)\triangleq(p(x_{t+1}=j\,|\,x_{t}=i))_{i,j}$ the predicted transition matrix when using a language model with parameters $\theta$. Under the same assumptions that in Proposition 1, we have:
  • (i)
    Whenever the maximum likelihood estimator trained with next-token-prediction (12) reaches its optimal loss $\mathcal{L}(\theta)$, we have
    $\hat{P}(\theta)_{i,j}=\sum_{k=1}^{K}p(z=z_{k}\,|\,x_{t}=i)(P_{k})_{i,j}=\frac{1}{\sum_{s=1}^{K}(\pi_{s})_{i}}\sum_{k=1}^{K}(\pi_{k})_{i}(P_{k})_{i,j}\;,$
    where $\pi_{k}\in[0,1]^{\mathcal{V}}$ is the stationary distribution of $P_{k}$.
  • (ii)
    The inequality (8) holds in this context, where the conditional entropy $\mathcal{H}(x\,|\,z)$ can be computed by a weighted sum the entropy rate of each of the $K$ Markov Chains:
    $\mathcal{H}(x\,|\,z)=-{\color[rgb]{0,0,0}T}\sum_{k=1}^{K}\sum_{i=1}^{|\mathcal{V}|}(\pi_{k})_{i}\Big{[}\sum_{j=1}^{|\mathcal{V}|}(P_{k})_{i,j}\;\mathrm{log}(P_{k})_{i,j}\Big{]}\,.$
    The entropy $\mathcal{H}(x)$, which is the lower bound of the MLE baseline, can be computed either exactly for short sequences, or approximated, e.g., through Monte-Carlo integration.
Proof of (i) The Cross entropy of the maximum-likelihood model is optimal whenever $p_{\theta}=p$.
In this case, for each $i,j\in\{1,\dots,|\mathcal{V}|\}$, we have $p_{\theta}(x_{t+1}=j\,|\,x_{t}=i)=p(x_{t+1}=j\,|\,x_{t}=i)$ and
$p_{\theta}(x_{t+1}=j\,|\,x_{t}=i)=\sum_{k=1}^{K}p(z=z_{k}\,|\,x_{t}=i)\,p(x_{t+1}=j\,|\,x_{t}=i,z_{k})\,.$ (21)
Furthermore, by Bayes’ rule, we have:
$p(z=z_{k}\,|\,x_{t}=i)=\frac{p(x_{t}=i\,|\,z=z_{k})p(z=z_{k})}{\sum_{s=1}^{K}p(x_{t}=i\,|\,z=z_{s})p(z=z_{s})}=\frac{(\pi_{k})_{i}}{\sum_{s=1}^{K}(\pi_{s})_{i}}\;,$
because we have assumed a uniform prior over the mixture components, i.e., $p(z=z_{k})=\frac{1}{K}$, and we assumed stationary regime so that $p(x_{t}=i\,|\,z=z_{k})=(\pi_{k})_{i}$. ∎
Proof of (ii) Because we assumed first-order Markov Chains, we have
$\mathcal{H}(x)=\sum_{t=1}^{T}\mathcal{H}(x_{t}\,|\,x_{<t})\,.$
Because we assumed stationary Markov chains, we have that $\mathcal{H}(x_{t}\,|\,x_{t-1})$ doesn’t depend on $t$ and $\mathcal{H}(x_{t}\,|\,x_{t-1})=\mathcal{H}(x_{2}\,|\,x_{1})=-\sum_{i=1}^{|\mathcal{V}|}(\pi_{k})_{i}\Big{[}\sum_{j=1}^{|\mathcal{V}|}(P_{k})_{i,j}\;\mathrm{log}(P_{k})_{i,j}\Big{]}\,$ (see [7] page 66, Theorem 4.2.4). ∎
Note that we considered first-order Markov chains in our analysis. However we expect the properties to generalize to higher orders ($n>1$) by using transition matrices in the form $P\in[0,1]^{\mathcal{V}^{n+1}}$ where
$P_{i_{1},\dots,i_{n+1}}=p(x_{t+1}=i_{n+1}\,|\,x_{t}=i_{n},\dots,x_{t-n+1}=i_{1})\;.$
Appendix D Experimental details of Section 4.3
Figure 4: Comparison of LoRA-MCL with standard maximum likelihood estimation (MLE). The setup mirrors that of Figure 1, but uses different transition matrices. While the overall behavior remains consistent with Figure 1, we observe a distinct transition in the MLE loss. We interpret this as evidence that MLE increasingly incorporates contextual information as training progresses (see also [44]).
The results in Figure 1 were obtained using the setup described in this section. The same illustration with different transition matrices is shown in Figure 4. Note that in both figures, the loss and theoretical quantities are normalized by $T-1$, since the first token is excluded from the computation.
Dataset. We used a uniform mixture of two first-order homogeneous Markov chains with transition matrices $(P_{1},P_{2})=\big{(}P(p_{1},q_{1}),P(p_{2},q_{2})\big{)}$, with $P(p,q)\triangleq\left[\begin{array}[]{cc}1-p&p\\ q&1-q\end{array}\right]\;$ and $p,q\in[0,1]$. For the experiment in Figure 1, we used $(p_{1},q_{1})=(0.2,0.9)$ and $(p_{2},q_{2})=(0.8,0.25)$. Figure 4 uses $(p_{1},q_{1})=(0.7,0.8)$ and $(p_{2},q_{2})=(0.8,0.25)$. The first state of the sequences sampled according to the stationary distribution of the Markov chain given by $\pi_{k}=\frac{1}{p_{k}+q_{k}}(q_{k},p_{k})$ (see e.g., [44]). The sequences have a fixed length of $T=32$.
Architecture. We considered a GPT-2-like architecture [56] using the GPT-Neo implementation [4] using local-attention suggested by [44] to improve convergence on Markov chain data (with window size of $5$). The model has a hidden size of $64$, $2$ layers of transformer blocks with $2$-heads attention. LoRA adapters for the MLE baseline have rank $r=64$, $\alpha=64$, and dropout disabled. To align the numbers of parameters when $K=2$, we used $r=32$ (and $\alpha=32$) for LoRA-MCL. The models have a total of $65{,}536$ trainable parameters over a total of $230{,}912$. Note that aligning the ranks of LoRA-MCL and LoRA-MCL leads to the same conclusions. Weights of the base model were kept frozen (both for LoRA-MLE and LoRA-MCL) to mimic the dynamics on larger language models.
Training details. We used a cosine scheduler with learning rate of $41e-$ 4, weight decay of $31e-$ 3 with AdamW optimizer, with $(\beta_{1},\beta_{2})=(0.9,0.95)$ as in [44]. We used a batch size of $128$, and trained for $500$ iterations, with validation loss computed every $10$ steps. Only the first $450$ iterations are plotted in the Figures 1 and 4. For the LoRA-MCL runs, we trained with vanilla Winner-Takes-All update. Figure 1 (left) shows the mean and standard deviation of the Winner-Takes-All validation loss across three training seeds, for $K=1$ (LoRA-MLE) and $K=2$ (LoRA-MCL).
Theoretical quantities computations To verify the theoretical results, we computed the following quantities:
  • Theoretical Optimal Loss of the MLE model. It is expressed as $\mathcal{H}(x)$ where $x\sim\frac{1}{K}\sum_{k=1}^{K}p(x\,|\,z_{k})$. It can be approximated by Monte-Carlo sampling with samples $(z_{s},x_{s})\sim p(x,z)$ by first sampling modes $z_{s}\sim\mathcal{U}\{1,\dots,K\}$, then $x_{s}\,|\,z_{s}\sim p(x\,|\,z_{s})$, and by computing
    $-\frac{1}{N}\sum_{s=1}^{N}\sum_{t=2}^{T}\mathrm{log}\,(P_{s})_{x_{s,t},x_{s,t+1}}\,.$ (22)
    Here, $(P_{s})_{x_{s,t},x_{s,t+1}}$ denotes the entry of $P_{s}$ at row $x_{s,t}$ and column $x_{s,t+1}$, and $N$ is the total number of samples. We used $N=50{,}000$ here. Note that the term corresponding to $t=1$ in (22) is discarded, as the first token is typically excluded from the loss computation. For the illustration, (22) was normalized by $T-1$, since the first token is discarded.
  • Theoretical Optimal Loss of the MCL model. It is computed by subtracting $\mathrm{log}K$ to the Theoretical optimal loss (22) of the MLE model. To verify $(iii)$ in Proposition 1, we also computed the mixture of entropy rates given by (11), using a normalization factor of $T-1$.
Appendix E Experimental details on the Audio and Image Captioning tasks
E.1 Setup
Dataset statistics are provided in Table 4.
Audio Datasets. We conducted experiments on two audio captioning datasets: Clotho-V2 [14, 10] and AudioCaps [16, 30]. For both datasets, we used the official training, validation, and test splits. Clotho-V2 provides five reference captions per audio clip across all splits, whereas AudioCaps includes a single caption per clip in the training set and five captions per clip in the validation and test sets. During training on Clotho-V2, which is performed on $10$ epochs, at each epoch, one of the five reference captions is sampled uniformly at random for each audio clip.
Vision datasets. We conducted experiments on the TextCaps dataset. Specifically, we used the official training split for training and the official validation split as our test set. Since each image in the dataset is annotated with five different captions, we duplicated each image five times—associating each duplicate with a distinct caption—to ensure that all reference captions are seen during a single training epoch.
Preprocessing of audio data. Following the implementation of [31], for Clotho and AudioCaps raw audio files are resampled from $44.1\,\mathrm{kHz}$ and $32\,\mathrm{kHz}$ respectively to $16\,\mathrm{kHz}$. They are then cropped to a maximum length of 30 seconds for Clotho and 10 seconds for AudioCaps. The data is then fed to the Qwen-2-Audio pipeline, which includes conversion of the raw waveform into a $128$-channel mel-spectrogram, with a window size of $25$ ms and a hop size of $10$ ms.
Preprocessing of image data. Our image pre-processing pipeline follows the recipe of LLaVA (i.e., resizing to $(336,336)$ and normalization using CLIP mean and standard deviation).
For both modalities, we used the HuggingFace transformers [78] and PEFT [46] Python libraries as part of the implementation. Note that for each of the experiments, we set the repetition penalty [29] to $1.1$ for decoding.
Table 4: Statistics of the audio and image captioning datasets. Num. samples includes {train, validation, test} sets, except for AudioCaps, where we used only {train, validation} sets.
Dataset Num. samples Duration (h) Num. Captions Modality
AudioCaps [30] 48,286 134.1 54 K Audio
Clotho [10] 5,929 37.0 30 K Audio
TextCaps [65] 25,119 N/A 126 K Image
E.2 Audio Captioning Experiments
E.2.1 Experimental setup
Architecture and training. We used the Instructed version of Qwen-2-Audio [6] as the base model, which features $\sim 8.4$ billion parameters. We trained using bfloat16 precision. We used LoRA adapters applied to the $Q,K,V$ linear projections of the attention modules, and the upside and downside projections of the feedforward blocks, for all the transformer blocks, which include both the audio encoder and language model decoder. We used a rank $r$, with $r=\alpha=8$ unless otherwise stated, with dropout equal to $0.1$. We trained with a batch size of $2$, with AdamW optimizer [40] (with $\beta_{1}=0.9$, and $\beta_{2}=0.98$), weight decay of $0.05$, using a cosine scheduler with minimum learning rate of $10^{-6}$ and maximum learning rate of $10^{-5}$, with a warmup ratio of $0.1$. Gradient clipping is used with a maximum gradient norm of $1.0$. The validation loss was computed once every epoch.
E.2.2 Additional results
Figure 5: Quality vs. Diversity on Audio Captioning with $4$ candidates.
Figure 6: Quality vs. Diversity on Audio Captioning with $3$ candidates.
Varying the number $K$ of hypotheses. Figures 5 and 6 present the same scatter plots as Figure 2, but with 4 and 3 candidates respectively. The observed trends remain consistent: LoRA-MCL continues to demonstrate an excellent balance between quality and diversity. Notably, the case with $\varepsilon=0$ performs particularly well, which we attribute to improved stability in vanilla WTA when the number of heads is reduced. Note that the detailed results with $K=5$ associated with Figure 2 are given in Tables 5 and 6 for Clotho and AudioCaps, respectively.
Table 5: Results on Clotho-V2 with 5 candidates and MAP Decoding. ‘BS’ and ‘DBS’ stand for beam search and diverse beam search respectively.
Training Decoding Beam Div1 Div2 mBLEU-4 BLEU-1 BLEU-4 METEOR ROUGE-L sBERT CIDEr-D SPICE SPIDEr
LoRA-MLE ($r=8$) BS 5 0.298 0.360 0.827 0.651 0.133 0.229 0.443 0.575 0.621 0.175 0.391
LoRA-MLE ($r=40$) BS 5 0.300 0.365 0.819 0.619 0.105 0.213 0.422 0.559 0.556 0.166 0.354
LoRA-MLE ($r=8$) BS 10 0.310 0.386 0.787 0.650 0.139 0.233 0.450 0.576 0.634 0.181 0.400
LoRA-MLE ($r=40$) BS 10 0.314 0.393 0.780 0.623 0.113 0.219 0.426 0.561 0.564 0.170 0.360
LoRA-MLE ($r=8$) BS 25 0.316 0.402 0.764 0.653 0.143 0.236 0.453 0.580 0.639 0.184 0.404
LoRA-MLE ($r=40$) BS 25 0.324 0.411 0.750 0.629 0.121 0.222 0.433 0.563 0.580 0.171 0.368
LoRA-MLE ($r=8$) DBS ($\lambda=0.5$) 5 0.421 0.560 0.524 0.679 0.145 0.241 0.467 0.597 0.674 0.193 0.421
LoRA-MLE ($r=40$) DBS ($\lambda=0.5$) 5 0.422 0.558 0.524 0.652 0.122 0.224 0.445 0.582 0.601 0.176 0.378
LoRA-MLE ($r=8$) DBS ($\lambda=0.8$) 5 0.451 0.606 0.446 0.683 0.149 0.242 0.470 0.600 0.678 0.194 0.423
LoRA-MLE ($r=40$) DBS ($\lambda=0.8$) 5 0.454 0.607 0.444 0.659 0.125 0.227 0.450 0.585 0.612 0.179 0.384
LoRA-MLE ($r=8$) DBS ($\lambda=1.0$) 5 0.466 0.627 0.413 0.685 0.147 0.243 0.471 0.602 0.683 0.195 0.426
LoRA-MLE ($r=40$) DBS ($\lambda=1.0$) 5 0.470 0.628 0.414 0.660 0.124 0.228 0.451 0.588 0.614 0.180 0.386
LoRA-MLE ($r=8$) DBS ($\lambda=0.5$) 10 0.376 0.475 0.722 0.667 0.139 0.238 0.459 0.593 0.656 0.190 0.413
LoRA-MLE ($r=40$) DBS ($\lambda=0.5$) 10 0.377 0.480 0.716 0.642 0.115 0.223 0.439 0.575 0.589 0.179 0.374
LoRA-MLE ($r=8$) DBS ($\lambda=0.8$) 10 0.386 0.491 0.708 0.668 0.139 0.238 0.459 0.595 0.652 0.191 0.412
LoRA-MLE ($r=40$) DBS ($\lambda=0.8$) 10 0.390 0.495 0.704 0.640 0.115 0.223 0.437 0.577 0.585 0.179 0.371
LoRA-MLE ($r=8$) DBS ($\lambda=1.0$) 10 0.392 0.495 0.704 0.664 0.138 0.240 0.455 0.596 0.655 0.192 0.413
LoRA-MLE ($r=40$) DBS ($\lambda=1.0$) 10 0.395 0.498 0.704 0.641 0.112 0.224 0.437 0.578 0.583 0.182 0.372
LoRA-MLE ($r=8$) DBS ($\lambda=0.5$) 25 0.304 0.369 0.813 0.650 0.134 0.229 0.445 0.574 0.625 0.176 0.393
LoRA-MLE ($r=40$) DBS ($\lambda=0.5$) 25 0.305 0.370 0.812 0.619 0.106 0.214 0.423 0.558 0.555 0.166 0.354
LoRA-MLE ($r=8$) DBS ($\lambda=0.8$) 25 0.302 0.366 0.818 0.649 0.133 0.228 0.443 0.574 0.620 0.176 0.391
LoRA-MLE ($r=40$) DBS ($\lambda=0.8$) 25 0.304 0.369 0.814 0.619 0.106 0.214 0.423 0.558 0.556 0.166 0.354
LoRA-MLE ($r=8$) DBS ($\lambda=1.0$) 25 0.302 0.365 0.819 0.648 0.133 0.229 0.444 0.574 0.620 0.177 0.391
LoRA-MLE ($r=40$) DBS ($\lambda=1.0$) 25 0.305 0.369 0.814 0.619 0.105 0.214 0.423 0.558 0.556 0.167 0.355
LoRA-MCL ($\varepsilon=0.0$) BS 1 0.490 0.633 0.389 0.662 0.127 0.231 0.453 0.597 0.643 0.186 0.401
LoRA-MCL ($\varepsilon=0.0005$) BS 1 0.464 0.609 0.433 0.672 0.132 0.233 0.459 0.597 0.649 0.185 0.404
LoRA-MCL ($\varepsilon=0.05$) BS 1 0.471 0.623 0.405 0.677 0.141 0.238 0.463 0.600 0.672 0.191 0.419
LoRA-MCL ($\varepsilon=0.0$) BS 2 0.477 0.622 0.407 0.675 0.147 0.237 0.468 0.598 0.670 0.191 0.417
LoRA-MCL ($\varepsilon=0.0005$) BS 2 0.449 0.594 0.453 0.687 0.151 0.243 0.471 0.597 0.679 0.193 0.424
LoRA-MCL ($\varepsilon=0.05$) BS 2 0.448 0.598 0.450 0.691 0.157 0.246 0.478 0.600 0.700 0.197 0.436
LoRA-MCL ($\varepsilon=0.0$) BS 5 0.473 0.618 0.420 0.680 0.153 0.242 0.471 0.599 0.689 0.194 0.429
LoRA-MCL ($\varepsilon=0.0005$) BS 5 0.443 0.585 0.474 0.688 0.154 0.246 0.475 0.598 0.684 0.197 0.429
LoRA-MCL ($\varepsilon=0.05$) BS 5 0.445 0.592 0.465 0.691 0.165 0.249 0.482 0.601 0.708 0.199 0.442
Table 6: Results for AudioCaps with 5 candidates and MAP Decoding.
Training Decoding Beam Div1 Div2 mBLEU-4 BLEU-1 BLEU-4 METEOR ROUGE-L sBERT CIDEr-D SPICE SPIDEr
LoRA-MLE ($r=8$) BS 5 0.238 0.303 0.881 0.601 0.145 0.304 0.490 0.682 0.582 0.239 0.400
LoRA-MLE ($r=40$) BS 5 0.232 0.298 0.888 0.576 0.134 0.296 0.472 0.679 0.510 0.234 0.362
LoRA-MLE ($r=8$) BS 10 0.252 0.331 0.849 0.606 0.156 0.309 0.497 0.686 0.608 0.242 0.414
LoRA-MLE ($r=40$) BS 10 0.246 0.324 0.861 0.581 0.147 0.300 0.479 0.681 0.520 0.238 0.369
LoRA-MLE ($r=8$) BS 25 0.263 0.352 0.825 0.604 0.163 0.310 0.500 0.687 0.616 0.245 0.418
LoRA-MLE ($r=40$) BS 25 0.258 0.350 0.830 0.583 0.156 0.304 0.483 0.685 0.531 0.240 0.373
LoRA-MLE ($r=8$) DBS ($\lambda=0.2$) 5 0.312 0.454 0.684 0.620 0.166 0.317 0.508 0.697 0.622 0.257 0.421
LoRA-MLE ($r=40$) DBS ($\lambda=0.2$) 5 0.310 0.462 0.676 0.595 0.157 0.314 0.492 0.697 0.534 0.255 0.376
LoRA-MLE ($r=8$) DBS ($\lambda=0.5$) 5 0.358 0.545 0.531 0.636 0.188 0.327 0.521 0.707 0.664 0.270 0.445
LoRA-MLE ($r=40$) DBS ($\lambda=0.5$) 5 0.354 0.551 0.528 0.610 0.183 0.324 0.506 0.707 0.577 0.267 0.400
LoRA-MLE ($r=8$) DBS ($\lambda=0.8$) 5 0.377 0.576 0.483 0.640 0.193 0.330 0.521 0.708 0.672 0.270 0.447
LoRA-MLE ($r=40$) DBS ($\lambda=0.8$) 5 0.374 0.585 0.473 0.614 0.185 0.324 0.509 0.708 0.588 0.269 0.405
LoRA-MLE ($r=8$) DBS ($\lambda=1.0$) 5 0.386 0.590 0.460 0.638 0.190 0.329 0.521 0.709 0.668 0.269 0.444
LoRA-MLE ($r=40$) DBS ($\lambda=1.0$) 5 0.382 0.598 0.452 0.612 0.181 0.323 0.505 0.709 0.581 0.266 0.401
LoRA-MLE ($r=8$) DBS ($\lambda=0.5$) 10 0.314 0.439 0.777 0.624 0.180 0.318 0.510 0.699 0.636 0.262 0.431
LoRA-MLE ($r=40$) DBS ($\lambda=0.5$) 10 0.309 0.440 0.780 0.605 0.170 0.315 0.496 0.699 0.558 0.260 0.390
LoRA-MLE ($r=8$) DBS ($\lambda=0.8$) 10 0.320 0.448 0.770 0.624 0.176 0.318 0.507 0.700 0.626 0.263 0.426
LoRA-MLE ($r=40$) DBS ($\lambda=0.8$) 10 0.316 0.451 0.771 0.604 0.167 0.313 0.496 0.698 0.556 0.258 0.388
LoRA-MLE ($r=8$) DBS ($\lambda=1.0$) 10 0.325 0.455 0.764 0.621 0.175 0.317 0.506 0.702 0.632 0.262 0.427
LoRA-MLE ($r=40$) DBS ($\lambda=1.0$) 10 0.322 0.456 0.769 0.602 0.165 0.313 0.495 0.699 0.553 0.258 0.386
LoRA-MLE ($r=8$) DBS ($\lambda=0.5$) 25 0.247 0.319 0.864 0.604 0.152 0.306 0.493 0.682 0.597 0.243 0.408
LoRA-MLE ($r=40$) DBS ($\lambda=0.5$) 25 0.241 0.315 0.872 0.582 0.141 0.298 0.477 0.683 0.523 0.236 0.370
LoRA-MLE ($r=8$) DBS ($\lambda=0.8$) 25 0.246 0.316 0.867 0.603 0.148 0.305 0.492 0.682 0.592 0.242 0.406
LoRA-MLE ($r=40$) DBS ($\lambda=0.8$) 25 0.240 0.311 0.876 0.581 0.138 0.298 0.475 0.682 0.523 0.235 0.369
LoRA-MLE ($r=8$) DBS ($\lambda=1.0$) 25 0.245 0.314 0.870 0.603 0.150 0.306 0.493 0.682 0.592 0.242 0.406
LoRA-MLE ($r=40$) DBS ($\lambda=1.0$) 25 0.239 0.310 0.878 0.581 0.140 0.298 0.476 0.682 0.525 0.236 0.370
LoRA-MCL ($\varepsilon=0.0$) BS 1 0.496 0.707 0.211 0.649 0.174 0.323 0.518 0.702 0.648 0.255 0.428
LoRA-MCL ($\varepsilon=0.0005$) BS 1 0.450 0.678 0.292 0.654 0.195 0.333 0.530 0.710 0.679 0.272 0.448
LoRA-MCL ($\varepsilon=0.05$) BS 1 0.438 0.663 0.326 0.640 0.198 0.332 0.517 0.710 0.611 0.272 0.414
LoRA-MCL ($\varepsilon=0.0$) BS 2 0.459 0.672 0.233 0.650 0.196 0.330 0.529 0.702 0.679 0.262 0.447
LoRA-MCL ($\varepsilon=0.0005$) BS 2 0.408 0.629 0.348 0.658 0.213 0.336 0.539 0.711 0.716 0.276 0.470
LoRA-MCL ($\varepsilon=0.05$) BS 2 0.397 0.616 0.384 0.632 0.208 0.336 0.524 0.710 0.629 0.274 0.424
LoRA-MCL ($\varepsilon=0.0$) BS 5 0.453 0.657 0.254 0.656 0.211 0.331 0.532 0.702 0.697 0.262 0.456
LoRA-MCL ($\varepsilon=0.0005$) BS 5 0.403 0.615 0.372 0.661 0.229 0.339 0.545 0.712 0.744 0.278 0.484
LoRA-MCL ($\varepsilon=0.05$) BS 5 0.392 0.599 0.419 0.640 0.221 0.337 0.530 0.710 0.660 0.274 0.442
Evaluating with Sampling-based decoding. We also report results using sampling-based decoding methods in Tables 7 and 8 for Clotho and AudioCaps. Specifically, we use Top-$k$ sampling with $k=50$, Top-$p$ sampling with $p=0.95$, and Typical sampling with a threshold of $0.95$. For all sampling methods, we apply a repetition penalty of $1.1$, following [29]. In these experiments, the temperature was set to $\eta=1.0$.
We observe that both LoRA-MLE and LoRA-MCL yield significantly higher diversity compared to MAP decoding, albeit at the cost of reduced output quality. This trade-off is evident when comparing results in Tables 5 and 6. While both methods show an overall improvement in quality on the CIDEr and SPIDEr metrics across datasets, performance on other metrics such as METEOR and SPICE remains relatively comparable. In terms of diversity, there does not appear to be a clear trend indicating which method performs best, as the conclusions vary depending on the dataset. These findings suggest the need for further evaluation with different values of epsilon to better understand the quality-diversity trade-off in sampling-based decoding. Additionally, a more in-depth exploration of how the number of generated hypotheses affects sampling quality—and its implications for test-time inference scaling with LoRA-MCL [85]—is left for future work.
Table 7: Results for Clotho with 5 candidates and Sampling-based decoding.
Training Decoding Div1 Div2 mBLEU-4 BLEU-1 BLEU-4 METEOR ROUGE-L sBERT CIDEr-D SPICE SPIDEr
LoRA-MLE ($r=8$) Top-$k$ sampling 0.635 0.814 0.113 0.585 0.073 0.214 0.400 0.590 0.486 0.176 0.315
LoRA-MLE ($r=8$) Typical $p$ sampling 0.645 0.817 0.108 0.589 0.077 0.215 0.402 0.587 0.492 0.177 0.319
LoRA-MLE ($r=8$) Nucleus (Top-$p$) sampling 0.646 0.818 0.106 0.592 0.074 0.214 0.401 0.589 0.489 0.178 0.318
LoRA-MCL ($\varepsilon=0.0005$) Top-$k$ sampling 0.642 0.816 0.110 0.600 0.079 0.214 0.406 0.588 0.500 0.175 0.322
LoRA-MCL ($\varepsilon=0.0005$) Typical $p$ sampling 0.648 0.815 0.111 0.595 0.082 0.215 0.409 0.584 0.506 0.175 0.327
LoRA-MCL ($\varepsilon=0.0005$) Nucleus (Top-$p$) sampling 0.649 0.817 0.107 0.595 0.079 0.214 0.408 0.586 0.504 0.171 0.325
Table 8: Results for AudioCaps with 5 candidates and Sampling-based decoding.
Training Decoding Div1 Div2 mBLEU-4 BLEU-1 BLEU-4 METEOR ROUGE-L sBERT CIDEr-D SPICE SPIDEr
LoRA-MLE ($r=8$) Top-$k$ sampling 0.562 0.810 0.134 0.560 0.115 0.299 0.457 0.699 0.442 0.241 0.318
LoRA-MLE ($r=8$) Nucleus (Top-$p$) sampling 0.576 0.810 0.141 0.560 0.121 0.299 0.460 0.699 0.450 0.241 0.325
LoRA-MLE ($r=8$) Typical sampling 0.567 0.809 0.143 0.550 0.115 0.300 0.452 0.697 0.390 0.241 0.294
LoRA-MCL ($\varepsilon=0.0005$) Top-$k$ sampling 0.583 0.830 0.098 0.582 0.115 0.298 0.470 0.700 0.498 0.239 0.344
LoRA-MCL ($\varepsilon=0.0005$) Nucleus (Top-$p$) sampling 0.597 0.831 0.098 0.575 0.124 0.298 0.468 0.697 0.497 0.237 0.344
LoRA-MCL ($\varepsilon=0.0005$) Typical sampling 0.597 0.831 0.098 0.575 0.124 0.298 0.468 0.697 0.497 0.237 0.344
E.3 Image Captioning Experiments
E.3.1 Experimental setup
Table 9: Quality and Diversity Evaluation on TextCaps with $3$ candidates. For each of the presented metrics, higher is better ()($\uparrow$)( ↑ ) except for mBLEU-4 ()($\downarrow$)( ↓ ). LoRA-MCL is trained with $\varepsilon=0.1$, $r=8$ and $\alpha=32$. LoRA-MLE is trained with $r=24$ and $\alpha=96$. For completeness, we also trained LoRA-MLE with $r=8$ and $\alpha=32$ in the rows marked with .
Training Decoding Beam Div1 Div2 mBLEU-4 BLEU-1 BLEU-4 METEOR ROUGE-L sBERT CIDEr-D SPICE SPIDEr
LoRA-MLE Beam Search 3 0.433 0.509 0.688 0.802 0.318 0.315 0.580 0.670 1.517 0.244 0.873
LoRA-MLE Beam Search 3 0.383 0.421 0.833 0.788 0.315 0.317 0.573 0.670 1.517 0.241 0.874
LoRA-MLE Beam Search 6 0.402 0.457 0.786 0.795 0.338 0.326 0.583 0.671 1.557 0.246 0.895
LoRA-MLE Beam Search 6 0.401 0.456 0.784 0.796 0.339 0.327 0.585 0.672 1.572 0.248 0.903
LoRA-MLE DBS ($\lambda=0.5$) 3 0.489 0.600 0.529 0.818 0.345 0.325 0.596 0.684 1.571 0.253 0.902
LoRA-MLE DBS ($\lambda=0.8$) 3 0.530 0.655 0.437 0.824 0.349 0.327 0.601 0.686 1.590 0.251 0.909
LoRA-MLE DBS ($\lambda=1.0$) 3 0.540 0.669 0.416 0.822 0.348 0.326 0.599 0.685 1.586 0.250 0.906
LoRA-MLE DBS ($\lambda=0.5$) 3 0.485 0.597 0.531 0.821 0.346 0.326 0.596 0.685 1.589 0.255 0.912
LoRA-MLE DBS ($\lambda=0.8$) 3 0.485 0.597 0.531 0.821 0.346 0.326 0.596 0.685 1.589 0.255 0.912
LoRA-MLE DBS ($\lambda=1.0$) 3 0.535 0.665 0.425 0.827 0.357 0.327 0.601 0.686 1.601 0.252 0.915
LoRA-MLE DBS ($\lambda=0.5$) 6 0.455 0.532 0.694 0.813 0.344 0.328 0.595 0.681 1.580 0.253 0.908
LoRA-MLE DBS ($\lambda=0.8$) 6 0.470 0.549 0.671 0.812 0.341 0.328 0.593 0.681 1.573 0.251 0.903
LoRA-MLE DBS ($\lambda=1.0$) 6 0.474 0.553 0.666 0.812 0.340 0.328 0.592 0.680 1.577 0.250 0.904
LoRA-MLE DBS ($\lambda=0.5$) 6 0.451 0.528 0.697 0.808 0.343 0.329 0.594 0.681 1.586 0.253 0.911
LoRA-MLE DBS ($\lambda=0.8$) 6 0.464 0.542 0.681 0.810 0.340 0.329 0.593 0.680 1.583 0.253 0.909
LoRA-MLE DBS ($\lambda=1.0$) 6 0.473 0.551 0.671 0.810 0.341 0.328 0.592 0.680 1.584 0.251 0.908
LoRA-MCL Beam Search 1 0.499 0.599 0.520 0.828 0.344 0.330 0.597 0.690 1.674 0.255 0.955
LoRA-MCL Beam Search 2 0.507 0.618 0.490 0.824 0.360 0.333 0.604 0.687 1.627 0.258 0.932
We used LLaVA 1.6 as the base model, which features 7.1 billion parameters. We trained using bfloat16 precision. We used LoRA adapters applied to the $Q,K,V$, up and down linear projections of each block of the language model. We used a rank $r$, with $r=8$ and $\alpha=32$ unless otherwise stated, with dropout equal to $0.1$. We trained each model with a batch size of $8$, with AdamW optimizer [40] (with $\beta_{1}=0.9$, and $\beta_{2}=0.98$), weight decay of $0.05$, using a cosine scheduler with maximum learning rate of $10^{-4}$, with a warmup ratio of $0.1$. Gradient clipping is used with a maximum gradient norm of $1.0$. We used 1 epoch for training, where we duplicated the image as many times as the number of its captions, such that the model sees exactly one time each caption.
E.4 Additional results
Table 9 presents the results of LoRA-MCL and LoRA-MLE. Almost all the metrics show the same trend in which LoRA-MCL outperforms LoRA-MLE with Beam Search (BS) and Diverse Beam Search (DBS) in terms of quality, although DBS produces more varied outputs. Depending on the rank ($r=8$ or $r=24$), we found that setting $\lambda=1.0$ or $\lambda=0.8$, respectively, yields the best quality scores for DBS. Similar to our experiments on Audio Captioning, increasing the rank of LoRA-MLE results in a slight degradation in performance but improves diversity. Additionally, increasing the number of beams in BS with LoRA-MLE results in slightly improved performance but reduced diversity. In contrast, increasing the number of beams in DBS with LoRA-MLE (with fixed $\lambda$) leads to declines in both quality and diversity. Interestingly, with LoRA-MCL, increasing the number of beams enhances both performance and diversity here.
E.4.1 Artificial multilingual dataset creation
To evaluate the behaviour of LoRA-MCL under a multi-modal distribution, we simulated an artificial bi-modal dataset by automatically translating half of the captions from English to French using T5-small [57], while keeping the prompts in English. More specifically, we randomly sampled half of the images and translated their five associated captions. All the training parameters are the same as those in the experiments on the original TextCaps dataset, except the learning rate, which we set at $2\times 10^{-5}$ (as the maximum value in the scheduler) in both the LoRA-MLE and LoRA-MCL.
During evaluation, to assess which head is considered as the winner (for the head specialization analysis), we selected the one that maximizes the SPIDEr score over the references of the given sample.
E.5 Computation details
We run the experiments mostly on H100 NVIDIA GPUs with $80$ GB of RAM. Trainings and inferences were launched on a single GPU for the Audio Captioning experiments, and up to $8$ GPUs for the Image Captioning experiments. The total computing resources used for this project, including failed experiments, amount to approximately $20{,}000$ GPU hours.

### v2

@@
-Multiple Choice Learning of Low Rank Adapters for Language Modeling
+Multiple Choice Learning of Low-Rank Adapters for Language Modeling
 Abstract
 We propose LoRA-MCL, a training scheme that extends next-token prediction in language models with a method designed to decode diverse, plausible sentence continuations at inference time.
-Traditional language modeling is an intrinsically ill-posed problem: given a context, multiple futures may be equally plausible.
-Our approach leverages Multiple Choice Learning (MCL) and the Winner-Takes-All (WTA) loss to efficiently handle ambiguity through Low-Rank Adaptation (LoRA).
-We provide a theoretical interpretation of applying Multiple Choice Learning to Language Modeling, assuming the data is generated from a mixture of distributions.
-To illustrate the proposed approach, we use data sampled from mixtures of Markov chains.
-We then demonstrate with extensive experiments on real-world visual and audio captioning tasks that our method achieves high diversity and relevance in generated outputs.
-**footnotetext: Equal Contribution.
+Traditional language modeling is an intrinsically ill-posed problem: given a context, multiple “futures” may be equally plausible.
+Our approach leverages Multiple Choice Learning (MCL) and the Winner-Takes-All loss to efficiently handle ambiguity through Low-Rank Adaptation.
+We provide a theoretical interpretation of applying MCL to language modeling, assuming the data is generated from a mixture of distributions.
+We illustrate the proposed approach using mixtures of Markov chains.
+We then demonstrate with experiments on visual and audio captioning, as well as machine translation, that our method achieves high diversity and relevance in generated outputs.
+The accompanying code and a general-purpose package for applying LoRA-MCL to a wide range of language models are made available.
+[^1] https://github.com/Victorletzelter/LoRA-MCL
+Winner-takes-all, Diversity, Ambiguity, Multiple Choice Learning, Language Modeling, Low Rank Adapters
 1 Introduction
 Predicting what a person will say next or describing the content of an audio or visual scene with text is difficult, if not impossible, to do with perfect accuracy.
-When the context is not informative enough, external factors may lead to different scenarios or modes of plausible text continuations yang2018breaking ; dieng2020topic ; mei2022diverse .
-In such ambiguous tasks, the conditional distribution over the space of output sentences given the input context may be multi-modal due to the underlying inherent uncertainty malinin2020uncertainty .
-Initially proposed for text processing, transformer-based language models have quickly become a general framework for modeling streams of tokens, which can also represent, for instance, images and audio signals touvron2023llama ; chu2024qwen2 ; liu2023visual .
-Such models are trained as next-token predictors and allow, by nature, for addressing this uncertainty, by generating plausible output sentences through a wide body of maximum a posteriori (MAP) or sampling-based decoding approaches welleckdecoding .
-While sampling allows exploration and diversity, it may lead to unreliable responses and requires truncation to avoid unexpected answers, partly due to the overestimation of low-probability tokens zhang2021trading ; hewitt2022truncation .
-When seeking reliable and expected answers, maximum a posteriori estimation techniques, like Beam Search beamSearch , look for sentences that maximize the model’s likelihood.
-However, these alternatives have drawbacks as they may lack diversity, be prone to repetition loops keskar2019ctrl , and may sound unnatural holtzmancurious .
-Some approaches, e.g., Diverse Beam Search (DBS) vijayakumar2018diverse , were therefore proposed to artificially increase the diversity at inference, e.g., through a diversity penalty parameter $\lambda$, to find a tradeoff between generation quality and sample diversity sucontrastive .
+When the context is not informative enough, external factors may lead to different scenarios or modes of plausible text continuations (Yang et al., 2018; Dieng et al., 2020; Mei et al., 2022b).
+In such ambiguous tasks, the conditional distribution over the space of output sentences given the input context may be multi-modal due to the underlying inherent uncertainty (Malinin and Gales, 2020).
+Initially proposed for text processing, transformer-based auto-regressive language models (Radford et al., 2018, 2019) have quickly become a general framework for modeling streams of tokens, which can also represent, for instance, images and audio signals (Touvron et al., 2023; Chu et al., 2024; Liu et al., 2023).
+Such models are trained as next-token predictors and allow, by nature, for addressing this uncertainty, by generating plausible output sentences through a wide body of maximum a posteriori (MAP) or sampling-based decoding approaches (Welleck et al., 2024).
+While sampling allows exploration and diversity, it may lead to unreliable responses and requires truncation to avoid unexpected answers, partly due to the overestimation of low-probability tokens (Zhang et al., 2021; Hewitt et al., 2022).
+When seeking reliable and expected answers, MAP estimation techniques, like Beam Search (Lowerre, 1976), look for sentences that maximize the model’s likelihood.
+However, these alternatives have drawbacks as they may lack diversity, be prone to repetition loops (Keskar et al., 2019), and may sound unnatural (Holtzman et al., 2019).
+Some approaches, e.g., Diverse Beam Search (Vijayakumar et al., 2018), were therefore proposed to artificially increase the diversity at inference, e.g., through a diversity penalty parameter $\lambda$, to find a tradeoff between generation quality and sample diversity (Su and Collier, 2023).
 In contrast with these methods, our approach aims to predict diverse sentences reflecting the ambiguity of the input context.
-Multiple Choice Learning (MCL) guzman2012multiple ; lee2016stochastic has emerged as a paradigm for addressing ambiguous tasks.
+Multiple Choice Learning (MCL) (Guzman-Rivera et al., 2012; Lee et al., 2016) has emerged as a paradigm for addressing ambiguous tasks.
 It generally consists of a network with a shared backbone and multiple output heads.
 During training, it utilizes the winner-takes-all loss for adaptively updating the head that performs the best for each example.
-This is a competitive training scheme that specializes each model to subsets of the conditional output distribution rupprecht2017learning ; letzelter2024winner .
-In this paper, we propose incorporating this idea for language model finetuning, leveraging multiple Low-Rank Adapters hu2022lora instead of multiple heads, which may be impractical due to computation requirements and architectural constraints.
+This is a competitive training scheme that specializes each model to subsets of the conditional output distribution (Rupprecht et al., 2017).
+In this paper, we propose incorporating this idea for language model finetuning, leveraging multiple Low-Rank Adapters (Hu et al., 2022) instead of multiple heads, which may be impractical due to computation requirements and architectural constraints.
 Our method natively generates diverse and plausible sequences in a single forward pass, aiming to best approximate the conditional output distribution.
-Our contributions are as follows:
-We propose a new paradigm that adapts MCL for token sequence modeling, with a method, LoRA-MCL, that is particularly suited for efficient finetuning of language models.
-We provide a theoretical analysis of our approach, which applies MCL to language modeling.
-Assuming the sequences are sampled from a mixture of distributions, we explain in particular why LoRA-MCL should enable capturing the modes of the data distribution.
-We then consider the case of Markov chains, for which we validate our claims with a well-designed toy example.
-We undertake an extensive experimental study to validate our approach, on audio and machine vision captioning tasks, demonstrating its wide applicability in challenging real-world scenarios and showing an excellent diversity–quality trade-off.
-The code will be made publicly available.
+Our main contributions are as follows:
+We propose a new paradigm that adapts MCL for token sequence modeling with LoRA-MCL, that is particularly suited for efficient finetuning of language models.
+We provide a theoretical analysis of our approach.
+Assuming the sequences are sampled from a mixture of distributions, we explain why LoRA-MCL should capture the data distribution modes, and validate the claims on Markov chains with a well-designed toy example.
+We conduct extensive experiments on audio and vision captioning, as well as machine translation, showing wide applicability and an excellent diversity–quality trade-off.
 2 Problem setup
@@
 Language modeling aims at learning the law $p(x\,|\,c)=\prod_{t=1}^{T}p(x_{t}\,|\,x_{<t},c)$ using a model $p_{\theta}$ with parameters $\theta$, by minimizing the following negative log-likelihood loss, which is equivalent to the maximum likelihood estimation (MLE):
-$\mathcal{L}(\theta)=-\mathbb{E}_{c,x}[\log p_{\theta}(x\,|\,c)]=\mathbb{E}_{c,x}\Big{[}-\sum_{t=1}^{T}\log p_{\theta}(x_{t}\,|\,x_{<t},c)\Big{]}\;,$ (1)
+$\mathcal{L}(\theta)=\mathbb{E}_{c,x}[-\sum_{t=1}^{T}\log p_{\theta}(x_{t}\,|\,x_{<t},c)]\;,$ (1)
 where $x_{<t}$ denotes the sequence of tokens prior to time $t$.
-Optimizing (1) is referred to as teacher-forcing williams1989learning , where $p_{\theta}$ is fed with target (instead of predicted) tokens $x_{<t}$ during training sutskever2014sequence .
-When using a transformer architecture vaswani2017attention , this is implemented with causal attention modules, which allow for computing the conditional distributions in parallel through all time steps within a single forward pass.
-During inference, decoding methods welleckdecoding proceed to generating sequences $\hat{x}$ from the trained model $p_{\theta}$ in an auto-regressive fashion.
+Optimizing (1) is referred to as teacher-forcing (Williams and Zipser, 1989), where $p_{\theta}$ is fed with target (instead of predicted) tokens $x_{<t}$ during training (Sutskever et al., 2014).
+When using a transformer architecture (Vaswani et al., 2017), this is implemented via causal attention modules, which allow for computing the conditional distributions in parallel through all time steps within a single forward.
+During inference, decoding methods (Welleck et al., 2024) proceed to generating sequences $\hat{x}$ from the trained model $p_{\theta}$ in an auto-regressive fashion.
 First, they start with a conditional distribution for the first token from the context: $p_{\theta}(x_{1}\,|\,c)$, which allows selecting $\hat{x}_{1}$.
-Then, for $t\geq 2$ they predict $p_{\theta}(x_{t}\,|\,\hat{x}_{<t},c)$, and select $\hat{x}_{t}$, until reaching either the sequence length limit or an end-of-sentence (EOS) token.
-The choice of the decoding method used to generate $K$ candidate sequences $\hat{x}^{1},\dots,\hat{x}^{K}$ depends on the purpose of the task, but the general goal is $(i)$ to get highly likely sentences, i.e., ones that maximize $p_{\theta}(\hat{x})$; and $(ii)$ to get sufficiently diverse sentences, as can be measured by $n$-gram similarity ippolito2019comparison .
-Although this is a widely adopted paradigm, we show next how this training and decoding pipeline can be improved: instead of artificially generating diversity at inference time, we aim at learning to predict sequences $\hat{x}^{1},\dots,\hat{x}^{K}$ that cover well the modes of the target distribution $p(x\,|\,c)$.
+Then, for $t\geq 2$ they predict $p_{\theta}(x_{t}\,|\,\hat{x}_{<t},c)$, and select $\hat{x}_{t}$, until reaching either the sequence length limit or an end-of-sentence token.
+The choice of the decoding method to generate $K$ candidate sequences $\hat{x}^{1},\dots,\hat{x}^{K}$ depends on the purpose of the task, but the general goal is $(i)$ to get highly likely sentences, i.e., ones that maximize $p_{\theta}(\hat{x}\,|\,c)$; and $(ii)$ to get diverse sentences, as can be measured by $n$-gram similarity (Ippolito et al., 2019).
+Although this is a widely adopted paradigm, we show next how this training and decoding pipeline can be improved: instead of artificially generating diversity at inference time, we aim at learning to predict sequences that cover well the modes of the target distribution p(x|c)p(x\,|\,c).
 3 Methodology
 3.1 Motivation
-In language modeling, topic models nigam2000text ; blei2003latent ; dieng2020topic are data-generating processes in which the ground-truth probability distribution of sequences is modeled as a mixture of latent components or topics.
+In language modeling, topic models (Nigam et al., 2000; Blei et al., 2003; Dieng et al., 2020) are data-generating processes in which the ground-truth probability distribution of sequences is modeled as a mixture of latent components or topics.
 For example, the sequence “I am eating …” may have multiple plausible continuations, but the likelihood of each depends heavily on contextual factors such as the speaker’s location, which influences their culinary habits.
@@
 In topic models, data generation proceeds by first sampling a topic $z\in\mathcal{Z}$ for each sentence (usually referred to as a document in the literature), and then sampling words (or $n$-grams) from the distribution associated with that topic.
-With this in mind, MLE in (1) may not be suitable yang2018breaking .
-While MLE is effective for estimating the overall distribution $p(x)$, it does not capture the individual components when $p(x)$ is expressed as a mixture, i.e., $p(x)=\sum_{k}p(z_{k})p(x\,|\,z_{k})$.
+With this in mind, MLE in (1) may not be suitable (Yang et al., 2018).
+While MLE is effective for estimating the distribution $p(x)$, it does not capture the components when it is expressed as a mixture, i.e., $p(x)=\sum_{k}p(z_{k})p(x\,|\,z_{k})$.
 In such cases, MLE tends to model the aggregate rather than distinguish topic-specific distributions $p(x\,|\,z_{k})$.
-3.2 Applying Multiple Choice Learning to language modeling
-Our approach is inspired by the multiple choice learning (MCL) literature guzman2012multiple ; lee2016stochastic .
+3.2 Applying MCL to language modeling
+Our approach is inspired by the multiple choice learning (MCL) literature (Guzman-Rivera et al., 2012; Lee et al., 2016).
 We propose the following training scheme, intending to enable the recovery of the different topics $z_{k}$.
@@
   1. 1.
-    For each training sample $(c,x)$ in the batch $\mathcal{B}$: Compute $p(x\,|\,c;\theta_{k})$ for $k\in\{1,\dots,K\}$, and choose the best model $k^{\star}(x,c)=\mathrm{argmax}_{k}\;p(x\,|\,c;\theta_{k})$.
+    For each training sample $(c,x)$ in the batch: Compute $p(x\,|\,c;\theta_{k})$ for $k\in\{1,\dots,K\}$, and choose the best model $k^{\star}(x,c)=\mathrm{argmax}_{k}\;p(x\,|\,c;\theta_{k})$.
   2. 2.
     Compute the winner-takes-all (WTA) loss as:
-    $\mathcal{L}^{\mathrm{WTA}}(\theta_{1},\dots,\theta_{K})=-\mathbb{E}_{c,x}\Big{[}\max_{k=1,\dots,K}\log p(x\,|\,c;\theta_{k})\Big{]}\,,$ (2)
+    $\mathcal{L}^{\mathrm{WTA}}(\theta_{1},\dots,\theta_{K})=-\mathbb{E}_{c,x}\Big[\max_{k=1,\dots,K}\log p(x\,|\,c;\theta_{k})\Big]$ (2)
     where $\log p(x\,|\,c;\theta_{k})=\sum_{t=1}^{T}\log p(x_{t}\,|\,x_{<t},c;\theta_{k})$, and perform an optimization step.
-In practice when using a finite batch size, (2) is approximated with
-$\mathcal{L}^{\mathrm{WTA}}(\theta_{1},\dots,\theta_{K})\simeq-\sum_{k=1}^{K}\sum_{\{(x,c)\in\mathcal{B}\;\,|\,\;k^{\star}(x,c)=k\}}\log p(x\,|\,c;\theta_{k})\;,$ (3)
-where each hypothesis $k$ is trained only on the subset of examples in which it is the winner, according to the assignment $k^{\star}(x,c)$.
-This training procedure, similar to a hard-EM style optimization min2019discrete ; wen2023an , is a competitive training scheme that encourages the different models to explore different areas of the data distribution.
-However, it is subject to two main issues: First, using $K$ models instead of a single one drastically increases the training time and memory cost, which may be intractable for large language models.
+This training procedure, similar to a hard-EM style optimization (Min et al., 2019; Wen et al., 2023), is a competitive training scheme that encourages the different models to explore different areas of the data distribution.
+However, it is subject to two main issues: First, using $K$ models instead of a single one drastically increases the training time and memory cost, which may be intractable for large language models (LLMs).
 Second, the optimization may be subject to collapse, where the same models are chosen as winners through the iterations, leaving the other models untrained.
@@
 3.3 LoRA-MCL method
-Multiple choice learning typically alleviates the high training cost issue of $K$ models by training a single model with several heads lee2016stochastic ; lee2017confident .
+Multiple choice learning typically alleviates the high training cost issue of $K$ models by training a single model with several heads (Lee et al., 2016, 2017).
 However, we argue that such an approach is not well-suited for fine-tuning language models.
-First, heads of most language models are quite large (for example in Qwen2-Audio chu2024qwen2 the lm_head has $d\times|\mathcal{V}|=4096\times 156032\simeq 640\mathrm{M}$ params), and standard MCL would not scale easily with the number of heads.
+First, heads of most language models are quite large (for example, in Qwen2-Audio (Chu et al., 2024) the lm_head has $d\times|\mathcal{V}|=4096\times 156032\simeq 640\mathrm{M}$ parameters), and standard MCL would not scale easily with the number of heads.
 Second, the initialization of the heads poses several challenges.
@@
 A complete re-initialization of the heads is detrimental to performance, as numerous training iterations would be necessary to reach the same level of knowledge as in the pretrained head.
-For these reasons, we consider a Low Rank Adapter (LoRA) approach hu2022lora due to its excellent trade-off between performance and computational requirements, as well as its wide adoption in the context of large model fine-tuning.
+These effects are empirically verified in Apx.
+E.6.3.
+For these reasons, we consider a Low Rank Adapter (LoRA) approach (Hu et al., 2022) due to its excellent trade-off between performance and computational requirements, as well as its wide adoption in the context of large model fine-tuning.
 Let $\theta$ be the parameters of the pretrained base model.
@@
 Let
-$\theta_{k}=\theta\cup\left\{(A_{\ell}^{k},B_{\ell}^{k})\;|\;\ell=1,\dots,L\right\}\;,$ (4)
+$\theta_{k}=\theta\cup\left\{(A_{\ell}^{k},B_{\ell}^{k})\;|\;\ell\in\{1,\dots,L\}\right\}\;,$ (3)
 be the set of parameters that are involved in hypothesis $k$, with $L$ being the total number of layers where LoRA is used.
-Training in the WTA fashion involves computing $p(x\,|\,c;\theta_{k})$ for $k=1,\dots,K$. To avoid collapse into a single head, we may use a (relaxed) winner-takes-all loss as a training objective:
-$\mathcal{L}^{\mathrm{WTA}}(\theta)=-\mathbb{E}_{c,x}\Big{[}\sum_{k=1}^{K}q_{k}\log p(x\,|\,c;\theta_{k})\Big{]}\,,$ (5)
-where $\{q_{k}\}$ is a set of coefficients that sum to $1$ and that give more weight to the winner index $q_{k^{\star}}$ than to the others, but still provides gradient weight to non-winner heads in order to avoid collapse.
-In practice, we use $q_{k^{\star}}=1-\varepsilon$ and $q_{k}=\frac{\varepsilon}{K-1}$ for $k\neq k^{\star}$ as in rupprecht2017learning .
+Training in the WTA fashion involves computing $p(x\,|\,c;\theta_{k})$ for $k=1,\dots,K$.
+To avoid situations where some heads may be under-trained, including the collapse when a single head is trained, we use the relaxation of the winner-takes-all training objective of the form
+$\mathcal{L}^{\mathrm{WTA}}(\theta_{1},\dots,\theta_{K})=-\mathbb{E}_{c,x}\Big[\sum_{k=1}^{K}q_{k}\log p(x\,|\,c;\theta_{k})\Big]\;.$ (4)
+where $\{q_{k}\}$ is a set of positive coefficients that sum to 1.
+These coefficients assign higher weight to the winning head $q_{k^{\star}}$ while still providing nonzero gradient contributions to the other heads, thereby mitigating collapse.
+We experimented with two relaxation techniques.
+First, Relaxed-WTA (Rupprecht et al., 2017) where $q_{k^{\star}}=1-\varepsilon$ and $q_{k}=\frac{\varepsilon}{K-1}$ for $k\neq k^{\star}$, with $\varepsilon>0$ a small constant.
+We also considered the annealed MCL method (Perera et al., 2024), which introduces a temperature parameter $\uptau$:
+$q_{k}(x,c;\uptau)=\frac{p(x\,|\,c;\theta_{k})^{\frac{1}{\uptau}}}{Z_{x,c}(\uptau)}\,,\;Z_{x,c}(\uptau)=\sum_{s=1}^{K}p(x\,|\,c;\theta_{s})^{\frac{1}{\uptau}}\;.$ (5)
+Here the temperature $t\mapsto\uptau(t)$ follows a decreasing schedule, typically $\uptau(t)=\uptau(0)\rho^{t}$ with $\rho<1$ and $\uptau(0)>0$.
+At high temperatures, training is distributed more evenly across all hypotheses, preventing collapse; as $\uptau\to 0$, the method converges to the greedy WTA regime.
+[…]
+Figure 1: LoRA-MCL.
+Components of a linear layer $\ell$ where LoRA is enabled, with context $c$ omitted.
+Frozen base weights $W_{\ell}$ are in blue; trainable LoRA adapters in light red.
+The forward pass (in gray) is computed independently for each hypothesis, where $h^{(1)},\dots,h^{(K)}$ denote the hidden states as in (34).
+Gradients (purple arrows) are stronger for the winning hypothesis ($k^{\star}$).
 3.4 Accelerating LoRA-MCL training with parallelization over the hypotheses
-A naive implementation of MCL for winner selection, as in Section 3.2, may require a loop over the $K$ hypotheses in the batch to determine the winner associated with each index of the batch, which would drastically slow down training.
-To alleviate this issue, we propose the following methodology.
-Let $\mathbf{x}\in\mathbb{R}^{b\times\mathcal{T}\times d}$ denote the input of the transformer architecture, where $b$ is the batch size, $\mathcal{T}\triangleq T+\tau$ is the total sequence length, and $d$ is the number of features.
-We duplicate $\mathbf{x}$, $K$ times along the batch size dimension to get $\mathbf{x}^{\prime}=(\mathbf{x}_{1},\dots,\mathbf{x}_{K})\in\mathbb{R}^{bK\times\mathcal{T}\times d}$.
-LoRA units at layer $\ell$ are then computed as:
-$\left[\begin{array}[]{c}\mathbf{x}_{1}\\ \mathbf{x}_{2}\\ \vdots\\ \mathbf{x}_{K}\end{array}\right]\leftarrow\left[\begin{array}[]{cccc}B^{1}_{\ell}A^{1}_{\ell}&0&0&0\\ 0&B^{2}_{\ell}A^{2}_{\ell}&0&0\\ \vdots&\vdots&\ddots&\vdots\\ 0&0&0&B^{K}_{\ell}A^{K}_{\ell}\end{array}\right]\left[\begin{array}[]{c}\mathbf{x}_{1}\\ \mathbf{x}_{2}\\ \vdots\\ \mathbf{x}_{K}\end{array}\right]+\left[\begin{array}[]{c}f_{\theta}^{\ell}(\mathbf{x}_{1})\\ f_{\theta}^{\ell}(\mathbf{x}_{2})\\ \vdots\\ f_{\theta}^{\ell}(\mathbf{x}_{K})\end{array}\right]\,,$ (6)
-where $f_{\theta}^{\ell}$ is the base model (whose parameters remain frozen during training).
-In practice, this computation is conducted using a Group Convolution operation.111 This can be done by first reshaping $\mathbf{x}^{\prime}$ to shape $(b\times\mathcal{T}\times Kd)$ and applying a nn.Conv1d($Kd$, $Kd$, kernel_size=1, groups=$K$) following the PyTorch layer implementation, reshaping back $\mathbf{x}^{\prime}$ to shape $(Kb\times\mathcal{T}\times d)$ before adding the base model output, and then repeating at the next LoRA unit.
-While this duplication virtually multiplies the batch size by $K$, the memory overhead remains manageable, assuming that $r\ll d$.
-The next section is dedicated to the inner properties of the proposed approach.
+A naive implementation of LoRA-MCL would require looping over the $K$ hypotheses in the batch to evaluate each candidate separately, which would drastically slow down training.
+To avoid this, we process all hypotheses in parallel.
+Specifically, given an input sequence, we duplicate it $K$ times along the batch dimension.
+Each copy is then passed through a LoRA-adapted transformer, but crucially, we implement this in a grouped fashion: instead of running $K$ independent forward passes, we combine them into a single batched operation where each group corresponds to one hypothesis.
+In practice, this can be achieved using a grouped 1D convolution (nn.Conv1d in PyTorch) with $K$ groups, so that each hypothesis uses its own LoRA weights while still sharing the frozen base model.
+This trick effectively multiplies the batch size by $K$ while keeping the memory overhead manageable (since the LoRA rank $r\ll d$).
+It removes the sequential loop, enabling efficient parallel training of all hypotheses.
+Details are in Apx.
+E.5.
 4 Theoretical Analysis
-We justify the use of MCL in the context of language modeling by assuming that the sequence distribution can be expressed as a mixture.
-We contrast this formulation with standard maximum likelihood estimation (1).
-In particular, Section 4.1 establishes a connection between our approach and the Expectation-Maximization (EM) algorithm and derives lower and upper bounds on the optimal achievable test loss.
-We then apply this analysis to the case of Markov chains in Section 4.2, where the dynamic of the method is simulated in a controlled setting in Section 4.3.
+We justify the use of MCL for language modeling by assuming the sequence distribution is a mixture.
+Section 4.1 links our method to EM and derives lower and upper bounds on the optimal achievable test loss.
+We then apply this analysis to the case of Markov chains and simulate the method’s dynamics in a controlled setting.
 4.1 Training dynamics and optimality conditions
-For the next-token prediction loss (1), one can show that: $\mathrm{min}_{\theta}\;\mathcal{L}(\theta)=\mathcal{H}(x\,|\,c)$, where $\mathcal{H}(x\,|\,c)\triangleq-\mathbb{E}_{c,x}[\mathrm{log}\;p(x\,|\,c)]$ is the entropy of $p$ mackay2003information .
-Following the rationale of Section 3.1, let us now assume the data distribution can be written as a mixture.
-In the case of the WTA loss under this setup, we have the following proposition.
-Proposition 1.
-Assume the data-generating process is $p(x\,|\,c)=\sum_{k=1}^{K}p(z_{k}\,|\,c)\;p(x\,|\,z_{k},c)$ (Asm.
-2) with the number of modes equal to the number of hypotheses.
-Also assume perfect model expressiveness (Asm.
-1), and large enough batch size to approximate the true risk well (Asm.
+For the next-token prediction loss in (1), one can show that: $\mathrm{min}_{\theta}\;\mathcal{L}(\theta)=\mathcal{H}(x\,|\,c)$, where $\mathcal{H}(x\,|\,c)\triangleq-\mathbb{E}_{c,x}[\mathrm{log}\;p(x\,|\,c)]$ is the entropy of $p$ (MacKay, 2003).
+Following the rationale of Sec.
+3.1, let us now assume the data distribution can be written as a mixture.
+In the case of the WTA loss, we have the following proposition.
+Proposition 1 (Proof in Apx. B).
+Assume a data-generating process $p(x\,|\,c)=\sum_{k=1}^{K}p(z_{k}\,|\,c)\;p(x\,|\,z_{k},c)$ (Asm.
+2), perfect model expressiveness (Asm.
+1), and large enough batch size to approximate the true risk (Asm.
 3).
@@
   • (i)
-    The Winner-Takes-All two-step optimization in LoRA-MCL acts as a conditional form of the hard-EM algorithm.
+    LoRA-MCL acts as a conditional form of the hard-EM.
   • (ii)
-    Under Assumption 4, and assuming (with one permutation) that $p(x\,|\,z_{k},c)=p(x\,|\,c;\theta_{k})$ for each $k$, $\mathcal{L}^{\mathrm{WTA}}(\theta)=-\mathbb{E}_{x,c}\left[\underset{k=1,\dots,K}{\mathrm{max}}\mathrm{log}\,p(x\,|\,c,z_{k})\right]$.
-In this case, we also have:
-    $\mathcal{L}^{\mathrm{WTA}}(\theta)=\mathcal{H}(x\,|\,c,z)\triangleq\mathbb{E}_{c}\left[\sum_{k=1}^{K}p(z_{k}\,|\,c)\mathcal{H}\big{(}x\,|\,c,z_{k}\big{)}\right]\,,$ (7)
-    where $\mathcal{H}(x\,|\,c,z)$ is the conditional entropy given the random variable $z$.
+    Assuming disjoint components for the data mixture (Asm. 4), and assuming $p(x\,|\,z_{k},c)=p(x\,|\,c;\theta_{k})$ for each $k$, then
+    $\mathcal{L}^{\mathrm{WTA}}(\theta)=\mathcal{H}(x\,|\,c,z)\,$, (6)
+    where $\mathcal{H}(x\,|\,c,z)\triangleq\mathbb{E}_{c}\left[\sum_{k=1}^{K}p(z_{k}\,|\,c)\mathcal{H}\big(x\,|\,c,z_{k}\big)\right]$ is the conditional entropy given the random variable $z$.
   • (iii)
     We have the following inequalities:
-    $\mathrm{min}_{\theta}\,\mathcal{L}(\theta)-\mathrm{log}\,K\overset{(a)}{\leq}\mathrm{min}_{\theta}\,\mathcal{L}^{\mathrm{WTA}}(\theta)\overset{(b)}{\leq}\mathcal{H}(x\,|\,c,z)\overset{(c)}{\leq}\mathrm{min}_{\theta}\,\mathcal{L}(\theta)\;,$ (8)
+    $\begin{split}\min_{\theta}\,\mathcal{L}(\theta)-\log K\overset{(a)}{\leq}\min_{\theta}\,\mathcal{L}^{\mathrm{WTA}}(\theta)\\ \overset{(b)}{\leq}\mathcal{H}(x\,|\,c,z)\overset{(c)}{\leq}\min_{\theta}\,\mathcal{L}(\theta)\;.\end{split}$ (7)
     where $\mathrm{min}_{\theta}\;\mathcal{L}(\theta)=\mathcal{H}(x\,|\,c)$.
-Proof. See Appendix B.∎
-$(i)$ in Proposition 1 describes the relationship between LoRA-MCL and the hard-EM algorithm.
-$(ii)$ provides an expression for the WTA loss as a conditional entropy, $\mathcal{H}(x\,|\,c,z)$, under the assumption of a perfect matching between the hypotheses and the modes.
-Proposition 1 also establishes in $(iii)$ both a lower and an upper bound on the optimal achievable loss for LoRA-MCL, given in $(a)$ and $(b)$, respectively.
+$(i)$ in Prop.
+1 describes the relationship between LoRA-MCL and the hard-EM algorithm.
+$(ii)$ provides an expression for the WTA loss as a conditional entropy, $\mathcal{H}(x\,|\,c,z)$, assuming a perfect matching between the hypotheses and the modes.
+$(iii)$ establishes both a lower and an upper bound on the optimal achievable loss for LoRA-MCL, given in $(a)$ and $(b)$, respectively.
 Note that the gap between these bounds is $\mathcal{H}(x\,|\,c,z)-\mathrm{min}_{\theta}\,\mathcal{L}(\theta)+\mathrm{log}\,K=-\mathbb{E}_{c}[\mathcal{I}(x,z\,|\,c)]+\mathrm{log}\,K$ where $\mathcal{I}$ denotes the mutual information.
@@
 To make the analysis more concrete, we consider the case where the sequence of tokens is generated from Markov chains.
-Formally, given a finite state space $\mathcal{V}$, a homogeneous Markov chain is defined by an initial distribution $\pi$ over $\mathcal{V}$ and a transition matrix $P\in[0,1]^{\mathcal{V}\times\mathcal{V}}$.
-A sequence $(x_{1},\dots,x_{T})$ is sampled from the Markov chain if $x_{1}\sim\pi$ and, for each $t\geq 1$, the conditional distribution of $x_{t+1}$ given $x_{t}$ is $x_{t+1}\,|\,x_{t}\sim P_{x_{t},\cdot}$, where $P_{i,j}=\mathbb{P}(x_{t+1}=j\,|\,x_{t}=i)$.
-In the following, we ignore the initial warm-up phase and we assume that $\pi$ is the stationary distribution of $P$, i.e., such that $\pi P=\pi$.
+Given a finite state space $\mathcal{V}$, a homogeneous Markov chain is defined by an initial distribution $\pi$ over $\mathcal{V}$ and a transition matrix $P\in[0,1]^{\mathcal{V}\times\mathcal{V}}$.
+A sequence $(x_{1},\dots,x_{T})$ is sampled from the Markov chain if $x_{1}\sim\pi$ and, for each $t\geq 1$, $x_{t+1}\,|\,x_{t}\sim P_{x_{t},\cdot}$, where $P_{i,j}=p(x_{t+1}=j\,|\,x_{t}=i)$.
+In the following, we ignore the initial warm-up phase, and we assume that $\pi$ is the stationary distribution of $P$.
 In this case, we will denote $x\sim\mathrm{MC}(P)$.
-While the study of the training dynamics of transformers on Markov Chain data has been investigated in previous works edelman2024evolution ; rajaraman2024analysis ; makkuva2024attention ; zekri2024large , our setup instead considers a mixture of Markov chains gupta2016mixtures ; kausik2023learning .
-Assuming a uniform mixture, the data generating process is $x\sim\frac{1}{K}\sum_{k=1}^{K}\mathrm{MC}(P_{k})$, which is equivalent to:
-$z_{k}\sim\mathcal{U}(1,\dots,K),\quad x\,|\,z_{k}\sim\mathrm{MC}(P_{k})\;.$ (9)
+While the study of the training dynamics of transformers on Markov Chain data has been investigated in previous works (Edelman et al., 2024; Rajaraman et al., 2024; Makkuva et al., 2024; Zekri et al., 2024), our setup instead considers a mixture of Markov chains (Gupta et al., 2016; Kausik et al., 2023).
+Assuming a uniform mixture, the data generating process is $x\sim\frac{1}{K}\sum_{k=1}^{K}\mathrm{MC}(P_{k})$;
+$k\sim\mathcal{U}(1,\dots,K),\quad x\,|\,z_{k}\sim\mathrm{MC}(P_{k})\;.$ (8)
 When training a language model on such sequences, we have the following Corollary from Prop. 1, where the context $c$ is ignored for simplicity.
-Corollary 1.
-As per Assumption 5, let us assume that the data-generating process writes as a uniform mixture of Markov chains of order $n=1$.
-We denote as $\hat{P}(\theta)\triangleq(p(x_{t+1}=j\,|\,x_{t}=i))_{i,j}$ the predicted transition matrix when using a language model with parameters $\theta$.
-Under the same assumptions that in Proposition 1, we have that:
+Corollary 1 (Proof in Apx. C).
+Assume the data-generating process is a uniform mixture of first-order Markov chains.
+Let $\hat{P}(\theta)\triangleq(p_{\theta}(x_{t+1}=j\,|\,x_{t}=i))_{i,j}$ be the predicted transition matrix.
+Under the same Asm.
+as in Prop.
+1:
   • (i)
-    Whenever the maximum likelihood estimator trained with next-token-prediction (12) reaches its optimal loss $\mathcal{L}(\theta)$, we have
+    When the MLE estimator trained with (1) reaches its optimal loss, we have
     $\hat{P}(\theta)_{i,j}=\frac{1}{\sum_{s=1}^{K}(\pi_{s})_{i}}\sum_{k=1}^{K}(\pi_{k})_{i}(P_{k})_{i,j}\;,$ (9)
@@
   • (ii)
-    The inequality (8) holds in this context, where the conditional entropy $\mathcal{H}(x\,|\,z)$ can be computed by weighting the entropy rate of each Markov Chain:
-    $\mathcal{H}(x\,|\,z)=-{\color[rgb]{0,0,0}T}\sum_{k=1}^{K}\sum_{i=1}^{|\mathcal{V}|}(\pi_{k})_{i}\Big{[}\sum_{j=1}^{|\mathcal{V}|}(P_{k})_{i,j}\;\mathrm{log}(P_{k})_{i,j}\Big{]}\,.$ (11)
-Proof. See Appendix C. ∎
-The entropy $\mathcal{H}(x)$, which is theoretical optimal loss of the MLE baseline, can be computed either exactly for short sequences, or approximated, e.g., through Monte-Carlo integration.
-Note that we considered first-order Markov chains in our analysis.
-We expect the properties generalize to higher orders ($n>1$) by using transition matrices in the form $P\in[0,1]^{\mathcal{V}^{n+1}}$ where $P_{i_{1},\dots,i_{n+1}}=p(x_{t+1}=i_{n+1}\,|\,x_{t}=i_{n},\dots,x_{t-n+1}=i_{1})$.
-Please refer to Appendix C for further discussions.
+    The inequalities (7) holds, where $\mathcal{H}(x\,|\,z)$ is a weighted average of the entropy rate of each Markov Chain.
+The entropy $\mathcal{H}(x)$, which is the optimal loss of the MLE baseline, can be computed either exactly for short sequences or approximated, e.g., through Monte-Carlo integration.
+Our analysis considers first-order Markov chains, but we expect the results to extend to higher orders (see Apx.
+C).
 4.3 Illustration with synthetic data
-Figure 1: Comparison of LoRA-MCL with vanilla maximum likelihood estimation (MLE).
-The experimental setup is detailed in Section 4.2.
-(Left) Validation loss over training steps (averaged across three random seeds) for LoRA-MLE (blue) and LoRA-MCL (red).
-The theoretical optimal MLE loss, approximated via Monte Carlo sampling, is given by the entropy $\mathcal{H}(x)$.
-The gray shaded region represents the lower and upper bounds of the theoretical optimal MCL loss, as given by $(a)$ and $(b)$ in (8).
-(Right) Learned transition matrices (top) are compared with the references (bottom).
-MLE converges approximately toward the weighted average $\bar{P}$ defined as the right-hand side of (10).
-In contrast, LoRA-MCL successfully recovers the two underlying modes.
-To illustrate our approach, let us evaluate our algorithm on a synthetic dataset, for which results are given in Figure 1.
+Figure 2: Comparison of MCL with MLE.
+(Left) Validation loss over training steps (averaged across three seeds) for LoRA-MLE (blue) and LoRA-MCL (red).
+The theoretical optimal MLE loss is the entropy $\mathcal{H}(x)$.
+The gray shaded region represents the bounds of the theoretical optimal MCL loss, as given by $(a)$ and $(b)$ in (7).
+(Right) Learned transition matrices (top) versus references (bottom).
+MLE converges approximately toward the weighted average $\bar{P}$ (right-hand side of (9)), whereas LoRA-MCL recovers the two modes.
+To illustrate our approach, let us evaluate our algorithm on a synthetic dataset, for which results are given in Fig.
+2.
 We used $\mathcal{V}=\{1,2\}$, and $(P_{1},P_{2})=(P(p_{1},q_{1}),P(p_{2},q_{2}))$, with $P(p,q)\triangleq\left[\begin{array}[]{cc}1-p&p\\ q&1-q\end{array}\right]\;$ and $p,q\in[0,1]$.
-We sampled data from a mixture of two Markov chains following (9): we generated a sequence by sampling first $k\sim\mathcal{U}\{1,2\}$ and then $P_{k}$.
-Once $P_{k}$ was set, we sampled uniformly the initial state, then sampled the Markov chain according to the transition matrix until reaching the maximum sequence length ($T=32$ here).
-We considered a GPT-2-like architecture radford2019language ; gpt-neo using local-attention suggested by makkuva2024attention to improve convergence on Markov chain data.
-We then trained the model with $1$ and $2$ hypotheses, using LoRA adapters as described above, $1$ hypothesis corresponding to vanilla maximum likelihood estimation.
-Training details are provided in Appendix D.
-Figures 1 and 4 show both the evolution of the losses along training and the predicted transition matrix from the trained models.
+We sampled data from a mixture of two Markov chains following (8): we generated a sequence by sampling first $P_{k}$ with $k\sim\mathcal{U}\{1,2\}$.
+Once $P_{k}$ was set, we sampled the initial state uniformly, then we sampled the Markov chain according to the transition matrix until reaching the maximum sequence length ($T=32$ here).
+We considered a GPT-2-like architecture (Radford et al., 2019; Black et al., 2021) using local-attention suggested by Makkuva et al.
+(2024) to improve convergence on Markov chain data.
+We then trained the model with one and two hypotheses, using LoRA adapters as described above, one hypothesis corresponding to vanilla MLE.
+Training details are provided in Apx.
+D.
+Figures 2 and 7 show both the evolution of the losses along training and the predicted transition matrix from the trained models.
 We see that the optimum is close to being global in this setup.
@@
 5 Empirical evaluation
-We evaluate LoRA-MCL on realistic datasets and large-scale models for audio and image captioning tasks.
-Predicting a textual description for images or audio signals is an ill-posed problem: from an input image or audio clip, multiple descriptions may be plausible.
-We argue this is a particularly relevant real-world case of ill-posed problems, where the conditional output distribution is inherently multi-modal.
-Our goal is to demonstrate that LoRA-MCL provides a competitive approach for capturing these distribution modes when fine-tuning either audio (Sec.
-5.2) or vision-language models (Sec.
-5.3).
-We present hereafter the experimental setup.
-Please refer to Appendix E.2 for further details.
-5.1 Experimental Setup
+We evaluate LoRA-MCL on realistic datasets and large-scale models for audio and image captioning tasks, as well as machine translation.
+Predicting a textual description for images or audio signals is an ill-posed problem: from an input image or audio clip, multiple descriptions may be plausible; this is a real-world case where the conditional output distribution is inherently multi-modal.
+Similarly, Machine Translation is a one-to-many problem (Ott et al., 2018).
+We demonstrate that LoRA-MCL provides a competitive approach for capturing these distribution when fine-tuning either audio, vision-language, or language-only models.
+We describe the experimental setup.
+See Apx.
+E.6 for details.
+5.1 Experimental Setup in captioning
+Figure 3: Quality–diversity trade-off on audio captioning (5 candidates).
+SPIDEr $(\uparrow)$ for quality, mBLEU-4 $(\downarrow)$ for diversity.
+Marker shape stands for the method, color for the decoding method, and size is proportional to forward passes per example at inference.
+LoRA-MLE uses $r\in\{8,8K\}$ for parameter parity.
+LoRA-MCL uses circle markers: Relaxed (black edge) and Annealed (red edge).
 Datasets.
-We experimented on both Clotho-V2 font2013freesound ; drossos2020clotho , and AudioCaps gemmeke2017audio ; kim2019audiocaps datasets for the audio captioning task, while we make use of the TextCaps dataset sidorov2020textcaps for a challenging task for vision language models coined as image captioning with reading comprehension.
-Table 4 of the Appendix describes the datasets sizes.
-Architecture and training details with audio data.
-We used the instructed version of Qwen2-Audio chu2024qwen2 as the base model, which has ${\sim}8.4$ billion parameters and a vocabulary size $|\mathcal{V}|=156{,}032$.
-We used LoRA adapters applied to the $Q,K,V$ linear projections of the attention modules, and the upside and downside projections of the feedforward blocks, for all the layers of the transformers.
+We experimented on both Clotho-V2 (Font et al., 2013; Drossos et al., 2020), and AudioCaps (Gemmeke et al., 2017; Kim et al., 2019) datasets for the audio captioning task, while we make use of the TextCaps dataset (Sidorov et al., 2020) for the task of image captioning with reading comprehension.
+Table 4 describes the datasets sizes.
+Experimental details in audio.
+We used the instructed Qwen2-Audio (Chu et al., 2024) as the base model, which has ${\sim}8.4$ billion parameters and a vocabulary size $|\mathcal{V}|=156{,}032$.
+We used LoRA adapters applied to the $Q,K,V$ linear projections of the attention modules, and the upside and downside projections of the feedforward blocks, across all layers.
 We used a rank $r$ and scaling factor $\alpha$, with $r=\alpha=8$ unless otherwise stated.
-We trained for 1 epoch on AudioCaps and 10 epochs on Clotho with a validation loss computed at the end of each epoch.
-Architecture and training details with visual data.
-We used LLaVA 1.6 liu2023visual with Vicuna-7B zheng2023judging for the LLM, as the base model which features ${\sim}7.1$ billion parameters and a vocabulary size $|\mathcal{V}|=32{,}000$.
-We used the official codebase for the implementation.
-Following standard practices zhou2024empirical , we applied LoRA adapters only for the LLM decoder.
+We trained for 1 and 10 epochs on AudioCaps and Clotho, respectively.
+Experimental details with visual data.
+We used LLaVA 1.6 (Liu et al., 2023), as the base model which features ${\sim}7.1$ billion parameters and a vocabulary size $|\mathcal{V}|=32{,}000$.
+We applied LoRA adapters only for the LLM decoder following (Zhou et al., 2024).
 The adapters were applied to $Q,K,V$, upside and downside projections as in Qwen2-Audio, and we used $r=\frac{\alpha}{4}=8$ unless otherwise stated.
@@
 Metrics.
-We evaluate the approach using various quality metrics.
-First, we compute the test-loss (or perplexity), which evaluates the likelihood of the reference captions under the trained model.
-Second, we use natural language generation quality metrics (BLEU-$n$, ROUGE, METEOR) papineni2002bleu ; lin2004rouge ; banerjee2005meteor based on $n$-gram matching between predicted and target sentences.
-We also leverage the latter proposed CIDEr, SPICE and SPIDEr vedantam2015cider ; anderson2016spice ; liu2017improved metrics, which show higher correlation with human evaluation for captioning tasks.
-Finally, we use Sentence-BERT (sBERT) reimers2019sentence ; zhou2022can for embedding comparison.
-We consider sentence-based oracle evaluation for these metrics (see e.g., (lee2016stochastic, , Sec.
-4.3) or labbe2022my ).
-To evaluate the diversity of the generated captions, we consider: Div-$n$, which compares the number of distinct $n$-grams against the number of $n$-grams for a set of generated captions as in mei2022diverse and Self-BLEU (mBLEU-4), which computes BLEU similarity scores between generated captions zhu2018texygen .
+We evaluate both quality and diversity.
+For quality, we report test-loss and standard Natural Language Generation metrics (BLEU, ROUGE, METEOR) (Papineni et al., 2002; Lin, 2004; Banerjee and Lavie, 2005), and captioning-specific scores (CIDEr, SPICE, and SPIDEr) (Vedantam et al., 2015; Anderson et al., 2016; Liu et al., 2017), which better correlate with human judgments in captioning.
+We also use Sentence-BERT (Reimers, 2019).
+We consider sentence-based oracle evaluation for these metrics (see Lee et al.
+(2016); Labbé et al.
+(2022)).
+For diversity, we used mBLEU-4 (Mei et al., 2022b) measuring similarity across generated captions (Zhu et al., 2018).
 Baselines.
-To validate our approach, we compare it against a maximum likelihood baseline (LoRA-MLE) trained using (1), under the same conditions as the ones considered for our multi-hypothesis model.
+We compare LoRA-MCL against the MLE baseline (LoRA-MLE) trained using (1), under the same conditions as the ones considered for our multi-hypothesis model.
 Specifically, both models use the same LoRA configuration, the same number of trainable parameters (the LoRA rank for the baseline is $K\times$ larger to this end), and the same number of iterations.
+We also considered a Mixture of Low Rank Experts (Muqeeth et al., 2024; Wu et al., 2024; Li et al., 2024) (LoRA-MoE) as a baseline.
+See Apx.
+E.3 for more details on the training methods.
 At inference time, for each decoding method applied to the baseline that returns $K$ sentences, we decode the same number of candidates with LoRA-MCL.
-When evaluating Maximum A Posteriori methods –greedy, beam search beamSearch and diverse beam search vijayakumar2018diverse –, we ensure a consistent computational budget by aligning the number of forward passes.
-For example, if the MLE baseline uses a beam size of $B$, our model uses a beam size of $\frac{B}{K}$ per hypothesis.
+When evaluating MAP methods such as greedy, beam search (BS) (Lowerre, 1976), and diverse beam search (DBS) (Vijayakumar et al., 2018), we ensure a consistent computational budget by aligning the number of forward passes.
+In this case, if LoRA-MLE or LoRA-MoE uses a beam size of $B$, our model uses a beam size of $\frac{B}{K}$ per hypothesis.
+Finally, we experimented Test-Time Augmentation (TTA) (Wang et al., 2019) with LoRA-MLE in Audio Captioning, applying SpecAugment (Park et al., 2019) $K$ times to the input Spectrogram to expect diverse outputs.
 5.2 Audio captioning
-Figure 2: Quality vs.
-Diversity on Audio Captioning with $5$ candidates.
-Quality measured by SPIDEr ()($\uparrow$)( ↑ ) and Diversity by mBLEU-4 ()($\downarrow$)( ↓ ).
-Marker shape stands for the method, and size is proportional to the number of forwards performed per example at inference time.
-Note that LoRA-MLE was trained with two rank values, $r=8$ and $r=8K$, for fair comparison in terms of parameter count.
-Color corresponds to the decoding method (beam search or diverse beam search).
-Values of $\varepsilon$ and $\lambda$ for LoRA-MCL and diverse beam search, resp., are indicated in the plot, with the color shade proportional to the corresponding parameter value to better distinguish the markers.
 Quality vs.
 diversity trade-off.
-Quality–diversity performance is shown in Figure 2.
-While a well-chosen value of $\lambda$ allows Diverse Beam Search applied to LoRA-MLE to be competitive, LoRA-MCL (circles) achieves the best trade-off between quality and diversity, appearing in the top-left corner of the plot.
-Although increasing the beam size generally improves performance for standard beam search, we observe that increasing the beam size within each group negatively impacts Diverse Beam Search.
-In theory, setting $\varepsilon=0$ should maximize both quality and diversity.
-However, using a small positive value of $\varepsilon$ ensures all heads are trained, albeit at the cost of reduced diversity and potentially quality when $\varepsilon$ is too large.
-Full results are given in Tables 5 and 6 of the Appendix.
+Quality–diversity performance is shown in Fig.
+3.
+For readability, only the TTA runs (orange diamonds) with optimal augmentation strength on the quality–diversity are displayed in Fig.
+3.
+It is strong on Clotho, but performs poorly on AudioCaps.
+We notice that a well-chosen value of $\lambda$ allows DBS applied to LoRA-MLE and LoRA-MoE to be competitive.
+LoRA-MCL (circles) achieves the best trade-off between quality and diversity, appearing in the top-left corner of the plot, where the best relaxation technique (annealed or relaxed) depends on the setup.
+Although increasing the beam size generally improves performance for standard beam search, we observe that increasing the beam size within each group in DBS can negatively impact DBS, as observed in Clotho.
+Additional results and comparisons are provided Apx.
+E.6.4.
+Table 1: Test Loss $(\downarrow)$ as a function of $K$. LoRA-MCL is trained using $\varepsilon=0.05$ and $r=8$.
+Training $K$ AudioCaps Clotho
+LoRA-MLE ($r=8$) 1 2.203 2.812
+LoRA-MLE ($r=8\times 3$) 1 2.195 2.868
+LoRA-MLE ($r=8\times 7$) 1 2.182 2.935
+LoRA-MCL 3 2.063 2.663
+LoRA-MCL 5 1.999 2.643
+LoRA-MCL 7 1.932 2.612
 Effect of the number of hypotheses.
-We include in Figures 6 and 5 (given in the appendix) the same plot as in Figure 2, but using 3 and 4 candidates, respectively.
-These results further reinforce the findings obtained with 5 hypotheses.
-In Table 5.2, we report the evolution of the test negative log-likelihood (NLL) loss as a function of the number of hypotheses, showing a monotonically decreasing trend.
-This provides additional evidence for Proposition 1, and shows that LoRA-MCL improves its coverage of the data distribution modes as $K$ increases.
-Table 1: Test Loss ()($\downarrow$)( ↓ ) as a function of the number of hypotheses $K$. LoRA-MCL is trained with $\varepsilon=0.05$ and $r=8$.
-Training $K$ AudioCaps Clotho
-LoRA-MLE ($r=8$) 1 2.152 2.811
-LoRA-MLE ($r=40$) 1 2.124 2.902
-LoRA-MCL 2 2.071 2.700
-LoRA-MCL 3 2.021 2.665
-LoRA-MCL 5 1.927 2.629
-LoRA-MCL 8 1.869 2.592
-Table 2: Quality and Diversity Evaluation on TextCaps with $3$ candidates.
-Best scores are in bold, second-best are underlined.
-For each of the presented metrics, higher is better ()($\uparrow$)( ↑ ) except for mBLEU-4 ()($\downarrow$)( ↓ ).
-LoRA-MCL is trained with $\varepsilon=0.1$, $r=8$ and $\alpha=32$.
-LoRA-MLE is trained with $r=24$ and $\alpha=4\times r=96$, ensuring the same dynamics and number of parameters across models.
-For completeness, we also trained LoRA-MLE with $r=8$ and $\alpha=32$ in the rows marked with .
-Training Decoding Beam mBLEU-4 BLEU-4 METEOR sBERT CIDEr-D SPICE SPIDEr
-LoRA-MLE Beam Search 3 0.688 0.318 0.315 0.670 1.517 0.244 0.873
-LoRA-MLE Beam Search 6 0.786 0.338 0.326 0.671 1.557 0.246 0.895
-LoRA-MLE DBS ($\lambda=0.8$) 3 0.437 0.349 0.327 0.686 1.590 0.251 0.909
-LoRA-MLE DBS ($\lambda=1.0$) 3 0.416 0.348 0.326 0.685 1.586 0.250 0.906
-LoRA-MLE DBS ($\lambda=1.0$) 3 0.425 0.357 0.327 0.686 1.601 0.252 0.915
-LoRA-MLE DBS ($\lambda=0.8$) 6 0.671 0.341 0.328 0.681 1.573 0.251 0.903
-LoRA-MLE DBS ($\lambda=1.0$) 6 0.666 0.340 0.328 0.680 1.577 0.250 0.904
-LoRA-MLE DBS ($\lambda=1.0$) 6 0.671 0.341 0.328 0.680 1.584 0.251 0.908
-LoRA-MCL Beam Search 1 0.520 0.344 0.330 0.690 1.674 0.255 0.955
-LoRA-MCL Beam Search 2 0.490 0.360 0.333 0.687 1.627 0.258 0.932
+Table 1 reports the test negative log-likelihood as a function of $K$.
+The monotonically decreasing trend provides further evidence for Prop.
+1, indicating that LoRA-MCL achieves better coverage of the data distribution modes as $K$ increases.
 5.3 Image description with reading comprehension
-5.3.1 Quality and Diversity Evaluation
-Similarly to Figure 2, Table 2 confirms the trends observed in the audio captioning task.
-At an equal number of forward passes, LoRA-MCL outperforms LoRA-MLE, even using DBS decoding with a $\lambda$ diversity parameter specifically optimized for the task (SPIDEr of $0.955$ against $0.915$).
-However, we noticed that the DBS decoding of LoRA-MLE tends to generate more diverse outputs than the greedy decoding of LoRA-MCL (mBLEU of $0.416$ against $0.520$).
-Interestingly, increasing the number of beams in each group decreases diversity and does not improve the performance of LoRA-MLE, while with LoRA-MCL, increasing the beam size (from 1 to 2) not only slightly increases the performance but also improves the diversity of the generations, indicating promising further inference-time scaling.
-Additional results are provided in Appendix (Table 9).
-5.3.2 Observing hypothesis specialization in bilingual image description
-To highlight the behavior of LoRA-MCL in a realistic case where one can control the modes of the data-generating process, we simulated an artificial bi-modal distribution of the dataset, similarly to the setup of the toy experiment of Section 4.3.
-We did so by translating half of the captions of the data from English to French (using T5-small T5 ), while keeping the prompts in English.
-We trained and evaluated a 2-hypothesis model in the LoRA-MCL setup, as well as a baseline trained in LoRA-MLE.
-We observed a clear specialization of each hypothesis towards a given language (i.e., one hypothesis learned French and the other English): at test-time, the winning head is the first one in $\sim 89\%$ of the French captions and the second one in $\sim 97\%$ of the English captions.
-Table 3 shows performance and diversity scores of both LoRA-MLE and LoRA-MCL on our synthetic test set.
-LoRA-MCL is decoded greedily while LoRA-MLE uses DBS with the $\lambda$ value maximizing its performance (0.8).
-Although the overall performances are similar, LoRA-MCL generates significantly more diverse captions (mBLEU-4 of 0.027 and 0.029 against 0.138 and 0.126 on the French and English subparts, respectively), and outperforms LoRA-MLE in the French part of the test set (SPIDEr of 0.464 against 0.411).
-These results relate to the synthetic data experiments of Sec.
-4.2: in the presence of two modes, LoRA-MLE learns an average of them.
-Here, we believe this is more shifted toward English as the original LLaVA model performs better in English than in French.
-In contrast, LoRA-MCL is expected to learn the two modes separately.
-We believe this is what explains the gain in performance in French, at the cost of a slight reduction in English performance.
-Fig.
-3 qualitatively illustrates the behavior of the models: LoRA-MLE learns a weighted average of the two modes, shifted towards English, and it is sometimes not able to output French captions within the 2 candidates as in the book example.
-We found LoRA-MLE to be particularly prone to errors when outputting French sentences: for instance, in the second generation of examples with an image of a bottle, LoRA-MLE enters a repetition loop.
-On the other hand, LoRA-MCL is less affected by those artifacts on French sentences, and successfully captures the two modes of the distribution, benefiting from the specialization of each hypothesis.
-Table 3: SPIDEr ()($\uparrow$)( ↑ ) & mBLEU-4 ()($\downarrow$)( ↓ ) on different parts of synthetic test set.
+Table 2: SPIDEr $(\uparrow)$ & mBLEU-4 $(\downarrow)$ on different parts of synthetic test set.
 Test subset Training SPIDEr mBLEU-4
@@
 {Un livre papier intitulé Papa Told Me.}
-Figure 3: Observing specialization in bilingual image description. Quantitative (Left) and Qualitative (Right) analysis for LoRA-MLE and LoRA-MCL in the setup of Section 5.3.2.
+Figure 4: Observing specialization in bilingual image description. Quantitative (Left) and Qualitative (Right) analysis for LoRA-MLE and LoRA-MCL in the setup of Sec. 6.
+Table 3: Quality and Diversity Evaluation on TextCaps (3 candidates). Best in bold; second-best underlined. Higher is better $(\uparrow)$ except mBLEU-4 $(\downarrow)$.
+Training Decoding Beam $\mathrm{mBLEU}_{4}$ $\mathrm{CIDEr}_{D}$ $\mathrm{SPICE}$ $\mathrm{SPIDEr}$
+LoRA-MLE BS 3 0.688 1.517 0.244 0.873
+LoRA-MLE BS 6 0.786 1.557 0.246 0.895
+LoRA-MLE DBS ($\lambda=0.8$) 3 0.437 1.590 0.251 0.909
+LoRA-MLE DBS ($\lambda=1.0$) 3 0.416 1.586 0.250 0.906
+LoRA-MLE DBS ($\lambda=0.8$) 6 0.671 1.573 0.251 0.903
+LoRA-MLE DBS ($\lambda=0.8$) 3 0.531 1.589 0.255 0.912
+LoRA-MLE DBS ($\lambda=1.0$) 3 0.425 1.601 0.252 0.915
+LoRA-MoE DBS ($\lambda=0.8$) 3 0.441 1.616 0.254 0.924
+LoRA-MoE DBS ($\lambda=1.0$) 3 0.421 1.622 0.253 0.926
+LoRA-MoE DBS ($\lambda=0.8$) 6 0.678 1.608 0.255 0.922
+LoRA-MCL BS 1 0.520 1.674 0.255 0.955
+LoRA-MCL BS 2 0.490 1.627 0.258 0.932
+Figure 5: Quality vs. Diversity in Machine Translation.
+5.3.1 Quality and Diversity Evaluation
+Evaluation in image captioning Table 3 confirms the trends observed in the audio captioning task.
+At an equal number of forward passes, LoRA-MCL outperforms LoRA-MLE and LoRA-MoE, even using DBS with a $\lambda$ diversity parameter specifically optimized for the task (SPIDEr of $0.955$ vs.
+$0.926$).
+Consistently with audio captioning, increasing the number of beams in each group can decrease diversity and does not improve the performance of LoRA-MLE and LoRA-MoE.
+However, we noticed that the DBS with LoRA-MLE tends to generate more diverse outputs than the greedy decoding of LoRA-MCL (mBLEU of $0.416$ vs.
+$0.520$).
+Combining DBS with LoRA-MCL, inspired by Guzman-Rivera et al.
+(2014), could help address this issue.
+Additional comparisons are provided in Apx.
+E.7.2.
+5.3.2 Observing hypothesis specialization in bilingual image description
+Figure 6: PCA of hypothesis candidate embeddings. PC1 and PC2 correspond to the first two principal components, explaining $31.81\%$ and $15.22\%$ of the variance respectively.
+To highlight the behavior of LoRA-MCL in a realistic case where one can control the modes of the data-generating process, we simulated an artificial bi-modal distribution of the dataset, similarly to the setup of the toy experiment of Sec.
+4.3.
+We did so by translating half of the captions of the data from English to French (using T5-small (Raffel et al., 2020)), while keeping the prompts in English.
+We trained a two-hypothesis LoRA-MCL model and LoRA-MLE baseline.
+We observed a specialization of each hypothesis towards a given language (one hypothesis learned French and the other English): at test-time, the winning head is the first one in $\sim 89\%$ of the French captions and the second one in $\sim 97\%$ of the English captions.
+Table 4 reports quality/diversity on the synthetic test set.
+LoRA-MCL uses greedy decoding, LoRA-MLE DBS with $\lambda=0.8$ (maximizing its performance).
+Overall performance is similar, but LoRA-MCL is notably more diverse (mBLEU-4: FR $0.027$ vs $0.138$, EN $0.029$ vs $0.126$) and outperforms on French (SPIDEr $0.464$ vs $0.411$), with a slight reduction in English performance.
+Consistently with Sec.
+4.2, LoRA-MLE learns an average of the two modes (likely biased to English from LLaVA pretraining), whereas LoRA-MCL separates them.
+Additional experimental evidence is provided in Apx.
+E.7.4.
+Fig.
+4 illustrates the behavior of the models: LoRA-MLE learns a weighted average of the two modes, shifted towards English, and sometimes fails to output French captions within the two candidates, as in the book example.
+We found LoRA-MLE to be prone to errors when outputting French sentences: in the generation examples with an image of a bottle, LoRA-MLE enters a repetition loop.
+On the other hand, LoRA-MCL is less affected by those artifacts on French sentences, and captures the two modes of the distribution, benefiting from hypothesis specialization.
+Fig.
+6 presents a PCA of the captions generated by each head, embedded using a Sentence-BERT model (StyleDistance from Patel et al.
+(2025)).
+The resulting frontiers are clearly visible, demonstrating a strong specialization of each head.
+See Apx.
+E.7.4 for further analysis.
+5.4 Diverse Machine Translation
+We evaluate LoRA-MCL for zero-shot machine translation with LLMs.
+Following Xu et al.
+(2024a), we use a two-stage paradigm: (1) full-parameter fine-tuning on a monolingual corpus, and (2) LoRA fine-tuning on parallel data.
+We build on ALMA-7B, stage-1 fine-tuned, which we fine-tune on parallel data from Xu et al.
+(2024a) (see Apx.
+E.8), comprising WMT’17–20 test sets and Flores-200 dev/test sets, restricted to English–German ($\sim$14k pairs).
+Evaluation uses the newstest2014 subset from Ott et al.
+(2018), containing 500 English sentences with ten German references each.
+Fig.
+5 displays the results, with a legend mirroring the one in Fig.
+3.
+LoRA-MCL (circles) is trained with $K=3$ and $\varepsilon=0.05$, and each model generates $3$ sequences per input.
+We evaluate LoRA-MLE with ranks $r=16$ (diamond) and $r=48$ (square), using BS (blue) with widths $3$, 66, and $9$, and DBS (green) with $\lambda=0.8$.
+We follow the quality–diversity evaluation protocol of Shen et al.
+(2019).
+Scores are reported using both Leave-One-Out BLEU (Loo-BLEU) and Pairwise-BLEU (Shen et al., 2019).
+The results show that LoRA-MCL achieves a balance between quality and diversity, confirming the effectiveness of the method.
 6 Related Work
-Multiple Choice Learning to predict diverse and plausible outputs.
-Multiple Choice Learning guzman2012multiple ; lee2016stochastic is a training paradigm that minimizes the oracle or winner-takes-all loss across a set of models, encouraging specialization rupprecht2017learning ; letzelter2024winner .
-While it has been noted that the collapse issue needs to be addressed with care rupprecht2017learning ; amcl , MCL has demonstrated broad applicability across machine learning tasks lee2017confident ; seo2020trajectory ; garcia2021distillation , typically using a shared backbone and multiple projection heads.
-However, using multiple heads may be impractical in large language models.
-Mixture-of-Experts jacobs1991adaptive ; shazeer2017outrageously can be an effective strategy to manage computational costs in this case.
-However, as primarily thought for model scalability and not diversity, only a small subset of experts are active at each forward pass via routing, making them incompatible with MCL, while still suffering from knowledge redundancy in the experts jiang2024mixtral .
-To the best of our knowledge, we are the first to adapt MCL to the setup of language modeling with next-token prediction, specifically, using internal models instantiated via multiple LoRA hu2022lora .
-Generating Multiple Outputs with Language Models.
+MCL to predict diverse and plausible outputs.
+MCL (Guzman-Rivera et al., 2012; Lee et al., 2016) is a training paradigm that minimizes the WTA loss across a set of models, encouraging specialization (Rupprecht et al., 2017; Letzelter et al., 2024).
+While the collapse issue needs to be addressed (Rupprecht et al., 2017; Perera et al., 2024), MCL has demonstrated broad applicability across tasks (Lee et al., 2017; Seo et al., 2020; Garcia et al., 2021), typically using a shared backbone and multiple heads.
+However, using multiple heads may be impractical in LLMs.
+Mixture-of-Experts (MoE) (Jacobs et al., 1991; Shazeer et al., 2017) offers an alternative for managing computational costs, since only a subset of experts is active at each forward pass.
+In LLMs, however, MoE has primarily been used to improve scalability rather than to encourage diversity, and suffers from redundancy among experts (Jiang et al., 2024).
+While MoE can be adapted to the LoRA setting (Wu et al., 2024; Li et al., 2024), there is no clear consensus on the degree of specialization achieved.
+To our knowledge, this work is the first to adapt MCL to next-token language modeling using multiple LoRA modules.
+Generating Multiple Outputs with LMs.
 Language models are commonly trained via next-token prediction, framed as Maximum Likelihood Estimation.
-This is arguably the most popular method for training large-scale language models shlegerislanguage ; radford2019language ; touvron2023llama , including those that take audio or images as input chu2024qwen2 ; liu2023visual , with much of its success attributed to tokenization makkuva2024attention ; rajaraman2024analysis .
-Generating diverse and plausible sequences at inference remains challenging: $(i)$ Sampling methods holtzmancurious ; fan2018hierarchical ; meister2023locally may be unreliable depending on the chosen parameters; $(ii)$ Exact MAP decoding is intractable due to the exponential search space eikema2020map ; $(iii)$ Strategies like Beam Search often yield repetitive or overly coherent outputs fan2018hierarchical ; holtzmancurious ; keskar2019ctrl .
-Diverse Beam Search mei2022diverse or contrastive decoding sucontrastive inject diversity through test-time parameters (e.g., penalty $\lambda$), but in contrast, our method infers diversity from the input’s inherent ambiguity.
+This is arguably the most popular method for training large-scale language models (Shlegeris et al., 2022; Radford et al., 2019; Touvron et al., 2023), including those that take audio or images as input (Chu et al., 2024; Liu et al., 2023), with much of its success attributed to tokenization (Makkuva et al., 2024; Rajaraman et al., 2024).
+Generating diverse and plausible sequences at inference remains challenging: $(i)$ Sampling methods (Holtzman et al., 2019; Fan et al., 2018; Meister et al., 2023) may be unreliable depending on the chosen parameters; $(ii)$ Exact MAP decoding is intractable due to the exponential search space (Eikema and Aziz, 2020); $(iii)$ Strategies like Beam Search often yield repetitive or overly coherent outputs (Fan et al., 2018; Holtzman et al., 2019; Keskar et al., 2019).
+Diverse Beam Search (Mei et al., 2022b) and Test-time Augmentation (Wang et al., 2019; Kim et al., 2022; Kaya et al., 2025) inject diversity through test-time parameters (e.g., penalty $\lambda$), but in contrast, LoRA-MCL infers diversity from input’s ambiguity.
 Diversity in Audio and Visual Captioning.
-Audio drossos2017automated ; mei2021audio ; mei2022automated and image captioning aneja2018convolutional ; herdade2019image ; hossain2019comprehensive have traditionally relied on dataset- and task-specific models trained with Maximum Likelihood Estimation (MLE).
-A key challenge identified in both domains is the lack of diversity in generated captions, prompting the development of other training objectives specifically designed to address this issue mei2022diverse ; xu2022diversity ; xu2024towards ; zhang2024generating ; wang2019describing ; wang2020diversity ; mahajan2020diverse .
-These methods often require architectural modifications and are typically designed for models trained from scratch.
-However, recent advances have shifted the field toward general-purpose multimodal large language models chu2024qwen2 ; liu2023visual , which offer broader capabilities but pose new challenges to address the diversity-quality trade-off.
-In this work, we demonstrate that LoRA-MCL can be effectively applied at the fine-tuning stage of such models.
+Audio (Drossos et al., 2017; Mei et al., 2021, 2022a) and image captioning (Aneja et al., 2018; Herdade et al., 2019; Hossain et al., 2019) have traditionally relied on MLE-trained, task-specific models.
+A key challenge is the limited diversity of generated captions, leading to training objectives tailored to this issue (Mei et al., 2022b; Xu et al., 2022, 2024b; Zhang et al., 2024; Wang and Chan, 2019; Wang et al., 2020; Mahajan and Roth, 2020).
+These approaches often require architectural changes and models trained from scratch.
+With the rise of general-purpose multimodal LLMs (Chu et al., 2024; Liu et al., 2023), addressing the diversity–quality trade-off remains critical.
+We show that LoRA-MCL effectively tackles this at the fine-tuning stage.
 7 Conclusion
-This paper proposes a new paradigm for training language models to generate diverse plausible predictions.
-This is done with LoRA-MCL, combining the Multiple Choice Learning paradigm and Low-rank Adaptation.
-We provide a theoretical analysis showing that when the target sequence is written as a mixture, LoRA-MCL is expected to capture the modes of the data distribution.
-We validate our analysis with experiments on Markov chains, and on the real-world audio and advanced image captioning tasks.
-Future work includes a careful interpretation of the concepts learned by each model as a function of the ambiguity of the input, e.g., for uncertainty estimation as an alternative to wang2023lora .
+LoRA-MCL combines MCL with LoRA to train language models for diverse, plausible predictions.
+We show that when the target sequence is a mixture, LoRA-MCL can capture the modes of the data distribution.
+We validate this on Markov chains as well as the tasks of audio, image captioning, and machine translation.
+Future work will focus on further interpreting the concepts learned by each model, and investiging recent recent LoRA variants such as Hayou et al.
+(2024); Zhang et al.
+(2025).
 Limitations.
-In LoRA-MCL, setting $\varepsilon$ too high ensures that all hypotheses receive gradients, but it can also overly homogenize the models, thereby reducing diversity.
-While MCL performs well across a broad range of $\varepsilon$ values, identifying the optimal setting can be challenging.
-To do so, we plan to incorporate recent annealing MCL techniques, e.g., amcl , which employ a scheduling mechanism for $\varepsilon$ and may help improve the quality of the solution.
-Note that recent LoRA variants, such as hayou2024lora+ may also be an orthogonal axis of improvement for our method.
-Acknowledgments
-We thank Mickael Chen for his valuable advice, as well as Gilles Puy for his help in the implementation of the parallelization over the hypotheses.
-This work was funded by the French Association for Technological Research (ANRT CIFRE contract 2022-1854), and granted access to the HPC resources of IDRIS (allocation 2024-AD011014345) by GENCI.
+In the relaxed variant, setting $\varepsilon$ too high guarantees that all hypotheses receive gradients, but it may also overly homogenize the models and reduce diversity.
+A similar effect arises in the annealed variant, where the temperature scheduler parameters influence performance.
+Exploring dynamic adjustment of these parameter values depending on the data distribution is left for future work.
 Appendix A Notations and setup
-In the following, $x\triangleq(x_{t})_{t=1}^{T}\in\mathcal{V}^{T}$ be a sequence of $T$ tokens belonging to a finite vocabulary $\mathcal{V}=\{1,\dots,|\mathcal{V}|\}$, and $c\triangleq(c_{t})_{t=1}^{\tau}\in(\mathbb{R}^{d})^{\tau}$ be a sequence of $\tau$ context of embeddings of dimension $d$.
+In the following, $x\triangleq(x_{t})_{t=1}^{T}\in\mathcal{V}^{T}$ be a sequence of $T$ tokens belonging to a finite vocabulary $\mathcal{V}=\{1,\dots,|\mathcal{V}|\}$, and $c\triangleq(c_{t})_{t=1}^{\tau}\in\mathcal{C}$ be a sequence of $\tau$ context embeddings of dimension $d$.
 In the following $\mathcal{X}=\mathcal{V}^{T}$ and $\mathcal{C}=(\mathbb{R}^{d})^{\tau}$.
@@
 In the following, we assume that the model $p_{\theta}$ is perfectly expressive.
-Formally, let $\mathcal{F}_{\theta}\triangleq\{p_{\theta}:\mathcal{C}\rightarrow\mathcal{P}(\mathcal{X})\}$ be the family of conditional distributions realized by the model, where $\mathcal{P}(\mathcal{X})$ is the set of probability distributions on $\mathcal{X}$.
-We assume the family $\mathcal{F}_{\theta}$ is perfectly expressive, that is $\mathcal{F}_{\theta}=\{p^{\star}:\mathcal{C}\rightarrow\mathcal{P}(\mathcal{X})\}$.
-First, note that we have the Proposition 2, a well-known result that is due to the Gibbs inequality (e.g., [42]), for which we provide a proof for completeness.
-Proposition 2 ([42]).
+Formally, let $\mathcal{F}_{\Theta}\triangleq\{p_{\theta}:c\in\mathcal{C}\rightarrow p(\cdot\,|\,c)\in\mathcal{P}(\mathcal{X})\,|\,\theta\in\Theta\}$ be the family of conditional distributions realized by the model, where $\mathcal{P}(\mathcal{X})$ is the set of probability distributions on $\mathcal{X}$.
+We assume the family $\mathcal{F}_{\Theta}$ is perfectly expressive, that is $\mathcal{F}_{\Theta}=\mathcal{P}(\mathcal{X})^{\mathcal{C}}$.
+First, note that we have the Proposition 2, a well-known result that is due to the Gibbs inequality (e.g., MacKay (2003)), for which we provide a proof for completeness.
+Proposition 2 (MacKay, 2003).
 Under Assumption 1, for the next-token prediction loss (10), one can show that
 $\mathrm{min}_{\theta}\;\mathcal{L}(\theta)=\mathcal{H}(x\,|\,c)\,$, (11)
-where $\mathcal{H}(x\,|\,c)\triangleq-\mathbb{E}_{c,x}[\mathrm{log}\;p(x\,|\,c)]$ is the entropy of $p(\cdot\,|\,c)$.
+where $\mathcal{H}(x\,|\,c)\triangleq-\mathbb{E}_{c,x}[\mathrm{log}\;p(x\,|\,c)]$.
 Proof.
@@
 Because $\mathrm{log}$ is a concave function, we have under Jensen’s inequality:
-$-\int_{\mathscr{S}(p)}\mathrm{log}\left[\frac{p_{\theta}(x\,|\,c)}{p(x,c)}\right]p(x,c)\;\mathrm{d}x\mathrm{d}c\geq-\mathrm{log}\left(\int_{\mathscr{S}(p)}\frac{p_{\theta}(x\,|\,c)}{p(x,c)}p(x,c)\;\mathrm{d}x\mathrm{d}c\right)\geq 0\;.$ (14)
+$\int_{\mathscr{S}(p)}\mathrm{log}\left[\frac{p_{\theta}(x\,|\,c)}{p(x,c)}\right]p(x,c)\;\mathrm{d}x\mathrm{d}c\geq-\mathrm{log}\left(\int_{\mathscr{S}(p)}\frac{p_{\theta}(x\,|\,c)}{p(x,c)}p(x,c)\;\mathrm{d}x\mathrm{d}c\right)\geq 0\;.$ (12)
 However, because of the convention, the left-hand side of (12) is also equal to the integral over $\mathcal{X}\times\mathcal{C}$. This shows:
-$-\int_{\mathcal{X}\times\mathcal{C}}p(x,c)\;\mathrm{log}\;p_{\theta}(x\,|\,c)\;\mathrm{d}x\mathrm{d}c\geq-\int_{\mathcal{X}\times\mathcal{C}}p(x,c)\;\mathrm{log}\;p(x\,|\,c)\;\mathrm{d}x\mathrm{d}c=\mathcal{H}(x\,|\,c)\;,$
+$\int_{\mathcal{X}\times\mathcal{C}}p(x,c)\;\mathrm{log}\;p_{\theta}(x\,|\,c)\;\mathrm{d}x\mathrm{d}c\geq-\int_{\mathcal{X}\times\mathcal{C}}p(x,c)\;\mathrm{log}\;p(x\,|\,c)\;\mathrm{d}x\mathrm{d}c=\mathcal{H}(x\,|\,c)\;,$
 where the equality is reached for parameter $\theta$ such that $p_{\theta}=p$, whose existence is guaranteed by Assumption 1. ∎
@@
 Note that we have the equality:
-$-\mathbb{E}_{\alpha}[\mathrm{log}\,\beta(x)]=\mathrm{KL}(\alpha\,\|\,\beta)+\mathcal{H}(\alpha)\;,$ (15)
+$\mathbb{E}_{\alpha}[\mathrm{log}\,\beta(x)]=\mathrm{KL}(\alpha\,\|\,\beta)+\mathcal{H}(\alpha)\;,$ (13)
 where the left-hand side is usually referred as the Cross-Entropy, and $\mathcal{H}(\alpha)\triangleq-\mathbb{E}_{\alpha}[\mathrm{log}\,\alpha(x)]$ is the entropy of $\alpha$.
@@
 Let us define:
-$D(\theta,q)\triangleq\int_{\mathcal{X}}\sum_{k=1}^{K}q(k\,|\,x)\mathrm{log}\,p(x,z_{k};\theta)\;\mathrm{d}p(x),\leavevmode\nobreak\ D(\theta)\triangleq\int_{\mathcal{X}}\max_{k=1,\dots,K}\;\mathrm{log}\,p(x,z_{k};\theta)\;\mathrm{d}p(x)\;,$ (18)
+$D(\theta,q)\triangleq\int_{\mathcal{X}}\sum_{k=1}^{K}q(k\,|\,x)\mathrm{log}\,p(x,z_{k};\theta)\;\mathrm{d}p(x),\penalty 10000\ D(\theta)\triangleq\int_{\mathcal{X}}\max_{k=1,\dots,K}\;\mathrm{log}\,p(x,z_{k};\theta)\;\mathrm{d}p(x)\;,$ (16)
 where $q$ is a discrete distribution over $\{1,\dots,K\}$ with exactly one non-zero component that controls the assignment of each $x$ to a fixed $k^{\star}$.
@@
 Indeed, the expectation step comes down to computing $q(\theta^{(t)})$.
-For for the Maximization step, we have: $D\left(\theta^{(t+1)},q(\theta^{(t)})\right)\geq D\left(\theta^{(t)},q(\theta^{(t)})\right)$, by definition.
+For the Maximization step, we have: $D\left(\theta^{(t+1)},q(\theta^{(t)})\right)\geq D\left(\theta^{(t)},q(\theta^{(t)})\right)$, by definition.
 At the next expectation step, we have: $D\left(\theta^{(t+1)},q(\theta^{(t+1)})\right)\geq D\left(\theta^{(t+1)},q(\theta^{(t)})\right)$, because $q(\theta^{(t+1)})$ computes the best assignment given the parameters $\theta^{(t+1)}$.
 This shows that $D(\theta^{(t+1})\geq D(\theta^{(t)})$.
-First note that the main difference compared to the vanilla form of the (hard) EM algorithm, is that the goal is to fit here a conditional distribution $p(x\,|\,c)$ given pair $(c,x)\sim p(c,x)$.
-Furthermore, step 2 performs a gradient step (of the neural network weights) instead of a full maximization.
+The main difference compared to the vanilla form of the (hard) EM algorithm is that $(i)$ the goal here is to fit a conditional distribution p(x|c)p(x\,|\,c) given pairs $(c,x)\sim p(c,x)$, and $(ii)$ step 2 performs a gradient update (of the neural network weights) instead of a full maximization.
 Note that under Assumption 2 the complete data log-likelihood writes as $\mathrm{log}\,p(x\,|\,c;\theta)=\mathrm{log}\left[\;\sum_{k=1}^{K}p(x,z_{k}\,|\,c;\theta)\right]$.
@@
 Indeed, under Assumption 3, each training step of LoRA-MCL writes as the optimization of
-$\mathcal{L}^{\mathrm{WTA}}(\theta)=-\int_{\mathcal{X}\times\mathcal{C}}\max_{k=1,\dots,K}\mathrm{log}\;p(x\,|\,c;\theta_{k})\mathrm{d}p(c,x)\;,$ (19)
-which is expected to decrease at each training iteration.
-Note that because the loss is bounded from below (by 00), the sequence of real numbers $\{\mathcal{L}^{\mathrm{WTA}}(\theta^{(t)})\}_{t\geq 0}$ is therefore expected to converge.
+$\mathcal{L}^{\mathrm{WTA}}(\theta)=-\int_{\mathcal{X}\times\mathcal{C}}\max_{k=1,\dots,K}\mathrm{log}\;p(x\,|\,c;\theta_{k})\mathrm{d}p(c,x)\;.$ (17)
+The loss is generally expected to decrease across training iterations, although strict monotonicity is not guaranteed without additional assumptions on the learning rate and the smoothness of the gradient of the loss (Bertsekas, 1997).
+Since the loss is bounded below (by 0), the sequence of loss values $\{\mathcal{L}^{\mathrm{WTA}}(\theta^{(t)})\}_{t\geq 0}$ is therefore expected to converge in practice.
 To conclude, we can view $(\theta_{1},\dots,\theta_{K})$ as the parameters involved in the estimation of the modes of the conditional distribution with $p(x\,|\,c)=\sum_{k=1}^{K}p(x\,|\,c;\theta_{k})p(\theta_{k}\,|\,c)$.
-Note that the current form of the algorithm does not estimate the weight of each mode $p(\theta_{k}\,|\,c)$, further work could include incorporating scoring heads to estimate $p(\theta_{k}\,|\,c)$ each $k$ as in [35].
+Note that the current form of the algorithm does not estimate the weight of each mode $p(\theta_{k}\,|\,c)$, further work could include incorporating scoring heads to estimate $p(\theta_{k}\,|\,c)$ each $k$ as in Letzelter et al.
+(2023).
 ∎
-We then expect to be able to recover the distributions (with one permutation) $\{p(x\,|\,c,z_{k})\}$ from estimated $\{p(x\,|\,c;\theta_{k})\}$, assuming identifiability of the data generating mixture, which we expect to be made easier if the components are enough separated (See e.g., [59, Par.
-2.5] or [47, Sec.
-2.2]).
-Proof of (ii) Let us assume that (with one permutation) $p(x\,|\,z_{k},c)=p(x\,|\,c;\theta_{k})$ for $k\in\{1,\dots,K\}$. This is possible thanks to Assumption 1. Let us show that (16).
-Let us define:
+We then expect to be able to recover the distributions (with one permutation) $\{p(x\,|\,c,z_{k})\}$ from estimated $\{p(x\,|\,c;\theta_{k})\}$, assuming identifiability of the data generating mixture, which we expect to be made easier if the components are enough separated (e.g., Redner and Walker (1984) Par.
+2.5 or McLachlan et al.
+(2019) Sec.
+2.2).
+Proof of (ii) Let us assume that (with one permutation) $p(x\,|\,z_{k},c)=p(x\,|\,c;\theta_{k})$ for $k\in\{1,\dots,K\}$. This is possible thanks to Assumption 1. Let us show that (14). Let us define:
 $\mathcal{X}_{k}(c,\theta)\triangleq\Big\{x\in\mathcal{X}\,|\,\mathrm{log}\;p(x\,|\,c,\theta_{k})\geq\mathrm{log}\;p(x\,|\,c,\theta_{s})\;\forall s\in\{1,\dots,K\}\Big\}\;.$ (18)
@@
 Now let us leverage Assumption 1 to choose $\tilde{\theta}_{k}$ such that $p(x\,|\,c,\tilde{\theta}_{k})=p(x\,|\,c,z_{k})$ for each $k\in\{1,\dots,K\}$.
-In this case, $\mathrm{max}_{k}\,p(x\,|\,c,\tilde{\theta}_{k})\geq p(x\,|\,c,z)$ for each $z\in\{z_{1},\dots,z_{K}\}$, and $-\mathbb{E}_{z}\mathbb{E}_{x\,|\,z}\left[\mathrm{log}\frac{\mathrm{max}_{k}\,p(x\,|\,c,\tilde{\theta}_{k})}{p(x\,|\,c,z)}\right]\leq 0$.
+In this case, $\mathrm{max}_{k}\,p(x\,|\,c,\tilde{\theta}_{k})\geq p(x\,|\,c,z)$ for each $z\in\{z_{1},\dots,z_{K}\}$, and $\mathbb{E}_{z}\mathbb{E}_{x\,|\,z}\left[\mathrm{log}\frac{\mathrm{max}_{k}\,p(x\,|\,c,\tilde{\theta}_{k})}{p(x\,|\,c,z)}\right]\leq 0$.
 Then,
@@
 Assumption 5 (Markov Chain).
-We assume that the data-generating process writes as a uniform mixture of Markov chains of order $n\in\mathbb{N}\setminus\{0\}$, that is, for each $t$ and each $k$, $p(x_{t}\,|\,x_{<t},c,z_{k})=p(x_{t}\,|\,x_{t-1},\dots,x_{t-n},c,z_{k})$.
+We assume that the data-generating process can be written as a uniform mixture of Markov chains of order $n\in\mathbb{N}\setminus\{0\}$, that is, for each $t$ and each $k$, $p(x_{t}\,|\,x_{<t},c,z_{k})=p(x_{t}\,|\,x_{t-1},\dots,x_{t-n},c,z_{k})$.
 Corollary 2.
 As per Assumption 5, let us assume that the data-generating process writes as a uniform mixture of Markov chains of order $n=1$.
-We denote as $\hat{P}(\theta)\triangleq(p(x_{t+1}=j\,|\,x_{t}=i))_{i,j}$ the predicted transition matrix when using a language model with parameters $\theta$.
+Let $\hat{P}(\theta)\triangleq(p(x_{t+1}=j\,|\,x_{t}=i))_{i,j}$ be the predicted transition matrix when using a language model with parameters $\theta$.
 Under the same assumptions that in Proposition 1, we have:
@@
     The inequality (7) holds in this context, where the conditional entropy $\mathcal{H}(x\,|\,z)$ can be computed by a weighted sum the entropy rate of each of the $K$ Markov Chains:
-    $\mathcal{H}(x\,|\,z)=-{\color[rgb]{0,0,0}T}\sum_{k=1}^{K}\sum_{i=1}^{|\mathcal{V}|}(\pi_{k})_{i}\Big{[}\sum_{j=1}^{|\mathcal{V}|}(P_{k})_{i,j}\;\mathrm{log}(P_{k})_{i,j}\Big{]}\,.$
+    $\mathcal{H}(x\,|\,z)=-{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}T}\sum_{k=1}^{K}\sum_{i=1}^{|\mathcal{V}|}(\pi_{k})_{i}\Big[\sum_{j=1}^{|\mathcal{V}|}(P_{k})_{i,j}\;\mathrm{log}(P_{k})_{i,j}\Big]\,.$ (19)
     The entropy $\mathcal{H}(x)$, which is the lower bound of the MLE baseline, can be computed either exactly for short sequences, or approximated, e.g., through Monte-Carlo integration.
@@
 $\mathcal{H}(x)=\sum_{t=1}^{T}\mathcal{H}(x_{t}\,|\,x_{<t})\,.$
-Because we assumed stationary Markov chains, we have that $\mathcal{H}(x_{t}\,|\,x_{t-1})$ doesn’t depend on $t$ and $\mathcal{H}(x_{t}\,|\,x_{t-1})=\mathcal{H}(x_{2}\,|\,x_{1})=-\sum_{i=1}^{|\mathcal{V}|}(\pi_{k})_{i}\Big{[}\sum_{j=1}^{|\mathcal{V}|}(P_{k})_{i,j}\;\mathrm{log}(P_{k})_{i,j}\Big{]}\,$ (see [7] page 66, Theorem 4.2.4).
+Because we assumed stationary Markov chains, we have that $\mathcal{H}(x_{t}\,|\,x_{t-1})$ doesn’t depend on $t$ and $\mathcal{H}(x_{t}\,|\,x_{t-1})=\mathcal{H}(x_{2}\,|\,x_{1})=-\sum_{i=1}^{|\mathcal{V}|}(\pi_{k})_{i}\Big[\sum_{j=1}^{|\mathcal{V}|}(P_{k})_{i,j}\;\mathrm{log}(P_{k})_{i,j}\Big]\,$ (see Cover (1999) page 66, Theorem 4.2.4).
 ∎
@@
 While the overall behavior remains consistent with Figure 2, we observe a distinct transition in the MLE loss.
-We interpret this as evidence that MLE increasingly incorporates contextual information as training progresses (see also [44]).
+We interpret this as evidence that MLE increasingly incorporates contextual information as training progresses (see also (Makkuva et al., 2024)).
 The results in Figure 2 were obtained using the setup described in this section.
 The same illustration with different transition matrices is shown in Figure 7.
-Note that in both figures, the loss and theoretical quantities are normalized by $T-1$, since the first token is excluded from the computation.
+Note that in both figures, the loss and theoretical quantities are normalized by T1T-1, since the first token is excluded from the computation.
 Dataset.
@@
 Figure 7 uses $(p_{1},q_{1})=(0.7,0.8)$ and $(p_{2},q_{2})=(0.8,0.25)$.
-The first state of the sequences sampled according to the stationary distribution of the Markov chain given by $\pi_{k}=\frac{1}{p_{k}+q_{k}}(q_{k},p_{k})$ (see e.g., [44]).
+The first state of the sequences sampled according to the stationary distribution of the Markov chain given by $\pi_{k}=\frac{1}{p_{k}+q_{k}}(q_{k},p_{k})$ (see e.g., (Makkuva et al., 2024)).
 The sequences have a fixed length of $T=32$.
 Architecture.
-We considered a GPT-2-like architecture [56] using the GPT-Neo implementation [4] using local-attention suggested by [44] to improve convergence on Markov chain data (with window size of $5$).
-The model has a hidden size of $64$, $2$ layers of transformer blocks with $2$-heads attention.
+We considered a GPT-2-like architecture (Radford et al., 2019) using the GPT-Neo implementation (Black et al., 2021) using local-attention suggested by Makkuva et al.
+(2024) to improve convergence on Markov chain data (with window size of $5$).
+The model has a hidden size of $64$, 22 layers of transformer blocks with $2$-heads attention.
 LoRA adapters for the MLE baseline have rank $r=64$, $\alpha=64$, and dropout disabled.
 To align the numbers of parameters when $K=2$, we used $r=32$ (and $\alpha=32$) for LoRA-MCL.
-The models have a total of $65{,}536$ trainable parameters over a total of $230{,}912$.
-Note that aligning the ranks of LoRA-MCL and LoRA-MCL leads to the same conclusions.
+The models have a total of 65,53665{,}536 trainable parameters over a total of 230,912230{,}912.
+Note that aligning the ranks of LoRA-MLE and LoRA-MCL leads to the same conclusions.
 Weights of the base model were kept frozen (both for LoRA-MLE and LoRA-MCL) to mimic the dynamics on larger language models.
 Training details.
-We used a cosine scheduler with learning rate of $41e-$ 4, weight decay of $31e-$ 3 with AdamW optimizer, with $(\beta_{1},\beta_{2})=(0.9,0.95)$ as in [44].
+We used a cosine scheduler with learning rate of $10^{-4}$, weight decay of $10^{-3}$ with AdamW optimizer, with $(\beta_{1},\beta_{2})=(0.9,0.95)$ as in Makkuva et al.
+(2024).
 We used a batch size of $128$, and trained for $500$ iterations, with validation loss computed every $10$ steps.
-Only the first $450$ iterations are plotted in the Figures 1 and 4.
+Only the first $450$ iterations are plotted in the Figures 2 and 7.
 For the LoRA-MCL runs, we trained with vanilla Winner-Takes-All update.
@@
 It is expressed as $\mathcal{H}(x)$ where $x\sim\frac{1}{K}\sum_{k=1}^{K}p(x\,|\,z_{k})$.
-It can be approximated by Monte-Carlo sampling with samples $(z_{s},x_{s})\sim p(x,z)$ by first sampling modes $z_{s}\sim\mathcal{U}\{1,\dots,K\}$, then $x_{s}\,|\,z_{s}\sim p(x\,|\,z_{s})$, and by computing
-    $-\frac{1}{N}\sum_{s=1}^{N}\sum_{t=2}^{T}\mathrm{log}\,(P_{s})_{x_{s,t},x_{s,t+1}}\,.$ (22)
+It can be approximated by Monte-Carlo sampling with samples $(z_{s},x_{s})$ by first sampling modes $z_{s}\sim\mathcal{U}\{1,\dots,K\}$, then $x_{s}\,|\,z_{s}\sim p(x\,|\,z_{s})$, and by computing
+    $\frac{1}{N}\sum_{s=1}^{N}\sum_{t=2}^{T}\mathrm{log}\,(P_{s})_{x_{s,t},x_{s,t+1}}\,.$ (21)
     Here, $(P_{s})_{x_{s,t},x_{s,t+1}}$ denotes the entry of $P_{s}$ at row $x_{s,t}$ and column $x_{s,t+1}$, and $N$ is the total number of samples.
-We used $N=50{,}000$ here.
+We used N=50, [^0] N=50{,}000 here.
 Note that the term corresponding to $t=1$ in (21) is discarded, as the first token is typically excluded from the loss computation.
-For the illustration, (22) was normalized by $T-1$, since the first token is discarded.
+For the illustration, (21) was normalized by T1T-1, since the first token is discarded.
   • Theoretical Optimal Loss of the MCL model.
 It is computed by subtracting $\mathrm{log}K$ to the Theoretical optimal loss (21) of the MLE model.
-To verify $(iii)$ in Proposition 1, we also computed the mixture of entropy rates given by (11), using a normalization factor of $T-1$.
-Appendix E Experimental details on the Audio and Image Captioning tasks
+To verify $(iii)$ in Proposition 1, we also computed the mixture of entropy rates given by (19), using a normalization factor of T1T-1.
+Appendix E Experimental details Captioning and Translation tasks
 E.1 Setup
@@
 Audio Datasets.
-We conducted experiments on two audio captioning datasets: Clotho-V2 [14, 10] and AudioCaps [16, 30].
+We conducted experiments on two audio captioning datasets: Clotho-V2 (Font et al., 2013; Drossos et al., 2020) and AudioCaps (Gemmeke et al., 2017; Kim et al., 2019).
 For both datasets, we used the official training, validation, and test splits.
@@
 Preprocessing of audio data.
-Following the implementation of [31], for Clotho and AudioCaps raw audio files are resampled from $44.1\,\mathrm{kHz}$ and $32\,\mathrm{kHz}$ respectively to $16\,\mathrm{kHz}$.
+Following the implementation of (Labb et al., 2024), for Clotho and AudioCaps raw audio files are resampled from $44.1\,\mathrm{kHz}$ and $32\,\mathrm{kHz}$ respectively to $16\,\mathrm{kHz}$.
 They are then cropped to a maximum length of 30 seconds for Clotho and 10 seconds for AudioCaps.
 The data is then fed to the Qwen-2-Audio pipeline, which includes conversion of the raw waveform into a $128$-channel mel-spectrogram, with a window size of $25$ ms and a hop size of $10$ ms.
+During training, the language model processes sequences formatted as:
+“<Audio bos token><Audio tokens><Audio eos token>Generate the caption in English:<reference text><text eos>”,
+where <Audio bos token> and <Audio eos token> denote the beginning and end of the audio sequence, <Audio tokens> correspond to the audio features, <reference text> to the target caption (uniformly sampled from the available reference captions during training), and <text eos> to the end-of-sentence token.
+The loss is computed from the index of the first <reference text> token and includes the <text eos> token.
+During inference, the same formatting is used (except that the part <reference text><text eos> is discarded).
+This procedure follows the official Qwen-2-Audio documentation in the Transformers library.
 Preprocessing of image data. Our image pre-processing pipeline follows the recipe of LLaVA (i.e., resizing to $(336,336)$ and normalization using CLIP mean and standard deviation).
-For both modalities, we used the HuggingFace transformers [78] and PEFT [46] Python libraries as part of the implementation.
-Note that for each of the experiments, we set the repetition penalty [29] to $1.1$ for decoding.
+For both modalities, we used the HuggingFace transformers (Wolf et al., 2019) and PEFT (Mangrulkar et al., 2022) Python libraries as part of the implementation.
+Note that for each of the experiments, we set the repetition penalty (Keskar et al., 2019) to $1.1$ for decoding.
 Table 4: Statistics of the audio and image captioning datasets. Num. samples includes {train, validation, test} sets, except for AudioCaps, where we used only {train, validation} sets.
 Dataset Num. samples Duration (h) Num. Captions Modality
-AudioCaps [30] 48,286 134.1 54 K Audio
-Clotho [10] 5,929 37.0 30 K Audio
-TextCaps [65] 25,119 N/A 126 K Image
-E.2 Audio Captioning Experiments
-E.2.1 Experimental setup
+AudioCaps (Kim et al., 2019) 48,286 134.1 54 K Audio
+Clotho (Drossos et al., 2020) 5,929 37.0 30 K Audio
+TextCaps (Sidorov et al., 2020) 25,119 N/A 126 K Image
+E.2 Metrics
+In the following, we describe how to assess a language model in generating sequences $\hat{x}^{1},\dots,\hat{x}^{K}$ conditioned on a context $c$ (for instance, an audio recording paired with a captioning prompt in the case of Audio Captioning) using a given decoding method.
+In the case of LoRA-MCL, we denote by $\theta_{1},\dots,\theta_{K}$ the parameters of the language models corresponding to each hypothesis.
+We assume access to a set of $R\geq 1$ references $x^{1},\dots,x^{R}$ for each context, which can be regarded as samples from the ground-truth conditional distribution p(x|c)p(x\,|\,c).
+The evaluation is performed on a dataset of $N$ pairs $(c_{i},\{x_{i}^{1},\dots,x_{i}^{R}\})_{i=1}^{N}$.
+E.2.1 Negative log-likelihood
+Test NLL and Perplexity ($\downarrow$).
+A standard way to evaluate a language model is through its test loss (e.g., (Xiong et al., 2024)), which measures the average likelihood of the reference sentences under the trained model.
+When considering multiple language models, the oracle NLL is defined by averaging, for each reference, the best NLL across the $K$ hypotheses:
+$\mathrm{NLL}\triangleq-\frac{1}{NR}\sum_{i=1}^{N}\sum_{x\in\{x_{i}^{1},\dots,x_{i}^{R}\}}\underset{k=1,\dots,K}{\mathrm{max}}\;\frac{1}{T(x)}\sum_{t=1}^{T(x)}\log p(x_{t}\,|\,x_{<t},c_{i},\theta_{k})\;,$ (22)
+where $T(x)$ denotes the length of the reference sequence $x$ (in number of tokens), and $\mathrm{log}$ refers to the natural logarithm.
+In the context of LLMs, perplexity (Jelinek et al., 1977) is usually defined as $\mathrm{PPL}=\exp(\mathrm{NLL})$.
+It can be interpreted as the effective number of equally likely tokens among which the model is uncertain when predicting the next token.
+In particular, if the model always predicts a uniform distribution over the vocabulary $\mathcal{V}$ then $\mathrm{PPL}=|\mathcal{V}|$.
+E.2.2 Natural language generation quality metrics
+Originally developed within the human translation community, BLEU (Bilingual Evaluation Understudy) (Papineni et al., 2002), ROUGE (Recall-Oriented Understudy for Gisting Evaluation) (Lin, 2004), and METEOR (Metric for Evaluation of Translation with Explicit Ordering) (Banerjee and Lavie, 2005) were introduced to measure the closeness between a machine translation and a professional human translation.
+For consistency, all sequences are tokenized using the Penn Treebank Tokenizer (PTB Tokenizer).
+Implementations rely on the AAC Metrics and COCO Caption libraries.
+An $n$-gram refers to a group of $n$ consecutive tokens in a tokenized sequence.
+When comparing candidate and reference sentences, we use the $F_{\beta}$ score, which balances precision and recall:
+$F_{\beta}(\mathrm{precision},\mathrm{recall})\triangleq\frac{(1+\beta^{2})\mathrm{precision}\cdot\mathrm{recall}}{\beta^{2}\mathrm{precision}+\mathrm{recall}}\;.$ (23)
+Each metric $\mathcal{M}$ is defined as a function of a candidate $\hat{x}^{k}$ and a set of references $X_{i}=\{x_{i}^{1},\dots,x_{i}^{R}\}$.
+However, these metrics do not natively handle multiple candidates (see e.g., Lee et al.
+(2016)[Sec.,4.3]; Labbé et al.
+(2022)).
+We therefore adopt a sentence-based oracle evaluation, where the final score is computed as
+$\mathcal{M}_{\mathrm{Oracle}}=\frac{1}{N}\sum_{i=1}^{N}\underset{k=1,\dots,K}{\mathrm{max}}\mathcal{M}(\hat{x}^{k}_{i},X_{i})\;.$ (24)
+BLEU ($\uparrow$).
+For a given $n$, the modified $n$-gram precision $p_{n}$ (see Section 2.1 in Papineni et al.
+(2002)), measures the fraction of the unigrams in a predicted caption that also appear in the reference captions, while clipping the numerator by the maximum number of times a word occurs in any single reference translation.
+$\mathrm{BLEU}_{n}$ combines the n-grams precision up to length $n$ computing a geometric mean of $p_{s}$ for $s=1,\dots,n$.
+It also applies a multiplicative brevity penalty $\mathrm{BP}$ to penalty too short predicted captions compared to the reference sequences lengths.
+In this case, we have $\mathcal{M}=\mathrm{BLEU}_{n}$ and:
+$\mathrm{BLEU}_{n}(\hat{x},X)\triangleq\mathrm{BP}(\hat{x},X)\times\left(\prod_{s=1}^{n}p_{s}(\hat{x},X)\right)^{\frac{1}{n}}\;,$ (25)
+where $\mathrm{BP}(\cdot,\cdot)$ and $p_{s}(\cdot,\cdot)$ are functions of the candidates and the set of references (please refer to Section 2.3 of (Papineni et al., 2002), and the documentation for details).
+ROUGE ($\uparrow$).
+The ROUGE family of metrics (Lin, 2004) was originally introduced for the automatic evaluation of summaries, following BLEU, and is based on measuring $n$-gram co-occurrence between candidate and reference sequences.
+In this work, we use $\mathrm{ROUGE}_{L}$, which relies on the Longest Common Subsequence (LCS) between a candidate and a reference.
+Formally, we set $\mathcal{M}=\mathrm{ROUGE}_{L}$ with
+$\mathrm{ROUGE}_{L}(\hat{x},X)\triangleq\underset{x^{r}\in X}{\mathrm{max}}\;F_{\beta}(P_{\mathrm{LCS}}(\hat{x},x^{r}),R_{\mathrm{LCS}}(\hat{x},x^{r}))\;,$ (26)
+where the precision and recall are defined as:
+$P_{\mathrm{LCS}}(\hat{x},x^{r})=\frac{\mathrm{LCS}(\hat{x},x^{r})}{\mathrm{length}(\hat{x})},\;R_{\mathrm{LCS}}(\hat{x},x^{r})=\frac{\mathrm{LCS}(\hat{x},x^{r})}{\mathrm{length}(x^{r})}\;,$
+where $\beta=1.2$ by default.
+METEOR ($\uparrow$).
+Unlike BLEU, METEOR (Banerjee and Lavie, 2005) explicitly incorporates recall, measures word-level matches between a candidate and the reference, and accounts for grammaticality through word order.
+Here we set $\mathcal{M}=\mathrm{METEOR}$, defined as an $F_{\beta}$ score between precision and recall, with an additional fragmentation penalty that lowers the score when matching words are not in the correct order:
+$\mathrm{METEOR}(\hat{x},X)\triangleq\underset{x^{r}\in X}{\mathrm{max}}\left(1-\mathrm{Penalty}(\hat{x},x^{r})\right)F_{\beta}(P(\hat{x},x^{r}),R(\hat{x},x^{r}))$ (27)
+where the precision are recall of the unigram matches between a candidate and a reference are given by
+$P(\hat{x},x^{r})=\frac{\mathrm{matches}(\hat{x},x^{r})}{\mathrm{length}(\hat{x})},\;R(\hat{x},x^{r})=\frac{\mathrm{matches}(\hat{x},x^{r})}{\mathrm{length}(x^{r})}\;,$
+with $\beta=\frac{1}{3}$ by default.
+The penalty depends on the number of chunks, i.e., groups of consecutive matches in the correct order, and penalizes disordered matches (See Section 2.2 of Banerjee and Lavie (2005)).
+E.2.3 Captioning evaluation.
+Introduced specifically for image captioning, the more recent metrics CIDEr (Consensus-based Image Description Evaluation) (Vedantam et al., 2015), SPICE (Semantic Propositional Image Caption Evaluation) (Anderson et al., 2016), and SPIDEr (Liu et al., 2017) have demonstrated stronger correlation with human judgment.
+As in the previous section, we employ sentence-based oracle evaluation as defined in (24).
+CIDEr ($\uparrow$).
+CIDEr (Vedantam et al., 2015) assigns weights to $n$-grams in the candidate and reference captions using TF–IDF.
+An $n$-gram receives higher weight if $(i)$ its term frequency (TF) is high, i.e., it appears often in the sequence, and $(ii)$ it is informative, i.e., it occurs infrequently across the set of reference captions in the corpus.
+For each sequence $x$, we construct a vector $g^{n}(x)$ of dimension equal to the number of $n$-grams of length $n$, where the $k$-th component $g^{n}(x)]_{k}$ is the TF–IDF weight of the $k$-th $n$-gram in $x$.
+CIDEr then computes the average cosine similarity between the candidate and each reference:
+$\operatorname{CIDEr}_{n}(\hat{x},X)\triangleq\frac{1}{|X|}\sum_{x_{r}\in X}\frac{g^{n}(\hat{x})\cdot g^{n}(x^{r})}{\left\|g^{n}(\hat{x})\right\|\left\|g^{n}(x^{r})\right\|}\;,\quad\operatorname{CIDEr}\triangleq\frac{1}{4}\sum_{n=1}^{4}\operatorname{CIDEr}_{n}\;,$ (28)
+where $\cdot$ denotes the euclidean dot product.
+SPICE ($\uparrow$).
+SPICE (Anderson et al., 2016) was designed to capture semantic adequacy by focusing on objects, attributes, and relations rather than surface $n$-gram overlap.
+Given a set of object classes $C$, relations $R$, and attributes $A$, each caption $x$ is parsed into a scene graph $T(x)=O(x)\cup E(x)\cup K(x)$ where $O(x)\subseteq C$ is the set of objects, $E(x)\subseteq O(x)\times R\times O(x)$ encodes relations between objects, and $K(x)\subseteq O(x)\times A$ represents attributes associated with objects.
+SPICE is defined as an $F_{1}$ score over the tuples in the semantic graphs:
+$\operatorname{SPICE}(\hat{x},X)\triangleq F_{1}(P(\hat{x},X),R(\hat{x},X))\;,$ (29)
+with precision and recall given by
+$P(\hat{x},X)=\frac{\mathrm{matches}(T(\hat{x}),T(X))}{|T(\hat{x})|}\quad R(\hat{x},X)=\frac{\mathrm{matches}(T(\hat{x}),T(X))}{|T(X)|}\;,$
+Here, $\mathrm{matches}(T(\hat{x}),T(X))$ counts the number of matching tuples between the candidate and the reference semantic graphs.
+SPIDEr ($\uparrow$).
+SPIDEr (Liu et al., 2017) combines the strengths of CIDEr (capturing consensus through $n$-gram overlap) and SPICE (capturing semantic adequacy through scene graphs).
+It is defined as the simple average of the two metrics:
+$\mathrm{SPIDEr}(\hat{x},X)\triangleq\frac{\mathrm{CIDEr}(\hat{x},X)+\mathrm{SPICE}(\hat{x},X)}{2}\;.$ (30)
+sBERT Similarity ($\uparrow$).
+BERTScore (Zhang et al., 2020) leverages contextual embeddings from a pretrained BERT model (Devlin, 2018) to compute token-level similarity.
+This allows it to $(i)$ better match paraphrases and $(ii)$ capture long-range dependencies while penalizing semantic changes.
+sBERT Similarity (Zhou et al., 2022) considers Sentence BERT (Reimers, 2019) (by default paraphrase-TinyBERT-L6-v2) as the pretrained model due to its capability to compare semantics with a single embedding per sequence.
+Denote their contextual embeddings by $e(x),e(\hat{x})\in\mathbb{R}^{d}$.
+For multiple references, it is defined as an average cosine similarity:
+$\mathrm{sBERT}(\hat{x},X)\triangleq\frac{1}{R}\sum_{x^{r}\in X}\frac{e(\hat{x})^{\top}e(x^{r})}{\left\|e(\hat{x})\right\|\left\|e(x^{r})\right\|}\;.$ (31)
+E.2.4 Diversity Evaluation
+Quality evaluation measures how well candidate captions match the references.
+By contrast, diversity evaluation considers only the set of generated candidates, irrespective of the references.
+Below, we present the diversity metrics used in our setup.
+Note that the oracle formulation in (24) does not apply here.
+Div-$n$ ($\uparrow$).
+Div-$n$ is defined as the ratio between the number of distinct $n$-grams in the $K$ generated captions $\hat{x}^{1},\dots,\hat{x}^{K}$ and the total number of $n$-grams across those captions.
+Higher values indicate greater lexical diversity.
+mBLEU-$n$ ($\downarrow$).
+Mutual BLEU (mBLEU-$n$) is computed by treating each generated caption $\hat{x}^{k}$ as a candidate and evaluating its BLEU score against the remaining captions $\{x^{s}\;|\;s\neq k\}$.
+The final score is the average across all $K$ captions:
+$\mathrm{mBLEU}_{n}(\hat{x}^{1},\dots,\hat{x}^{K})=\frac{1}{K}\sum_{k=1}^{K}\mathrm{BLEU}_{n}(\hat{x}^{k},\{\hat{x}^{s}\;|\;s\neq k\})\;.$ (32)
+Lower $\mathrm{mBLEU}_{n}$ values indicate greater diversity among the generated captions.
+E.3 Training methods
+We describe after the specificity of each of the used training methods.
+LoRA-MLE.
+Through the article, we refer to LoRA-MLE as the training method that optimizes (1), where the LoRA adapters are trained, and the rest of the model is frozen.
+This corresponds exactly to the case where $K=1$ is LoRA-MCL (as described hereafter).
+We used the default initialization of Low-Rank adapters in PEFT library, that In Low-Rank adapters, $A$ is initialized with Kaiming Uniform (He et al., 2015) initialization $A\sim\mathcal{U}[-\frac{1}{\sqrt{d}},\frac{1}{\sqrt{d}}]$ where $d$ is the number of input features and $B$ is initialized with zeros.
+LoRA-MoE.
+We denote by LoRA-MoE the use of multiple adapters within LoRA modules at each layer where LoRA is applied, trained in the style of a Mixture of Experts (MoE) (Jacobs et al., 1991; Shazeer et al., 2017).
+Let the hidden state be $h\in\mathbb{R}^{\mathcal{B}\times T\times d}$, where $\mathcal{B}$ is the batch size, $T$ the sequence length, and $d$ the feature dimension.
+Under the soft MoE formulation of (Muqeeth et al., 2024), the computation at layer $\ell$ is given by
+$h\leftarrow f_{\theta}^{\ell}(h)+\sum_{k}[\gamma(h)]_{k}hA_{\ell}^{k}B_{\ell}^{k}\;,$
+where $W^{\ell}$ is the frozen base model at layer $\ell$, and the router $\gamma:\mathbb{R}^{d}\rightarrow\Delta^{K-1}$ is applied independently to each token embedding $h_{b,t}\in\mathbb{R}^{d}$, producing a $K$-dimensional vector of mixing weights.
+In practice, we implement $\gamma$ as a linear projection followed by a softmax.
+In our experiments, we initialized the linear layer weights to zero so that each expert contributed equally at the start of training.
+In large language models, MoE typically employs hard (discrete) routing, e.g., top-$k$ selection from the router to control computational cost.
+In the context of LoRA, however, this tradeoff is less critical since LoRA computations are lightweight relative to the base model.
+Here, we adopt soft routing, which keeps the expert block fully differentiable and avoids reliance on gradient approximation or additional load-balancing losses (Wu et al., 2024; Li et al., 2024).
+LoRA-MCL.
+This is the method introduced in this paper.
+At each LoRA-enabled layer $\ell$, a family of $K$ LoRA adapters $(A_{k}^{\ell},B_{k}^{\ell})_{k=1}^{K}$ is trained, while the rest of the model remains frozen.
+The training objective is
+$\mathcal{L}^{\mathrm{WTA}}(\theta)=-\mathbb{E}_{c,x}\Big[\sum_{k=1}^{K}q_{k}(x,c)\log p(x\,|\,c;\theta_{k})\Big]\,$,
+where the coefficients $q_{k}$ depend on the chosen WTA mode.
+Let $k^{\star}(x,c)=\mathrm{argmax}_{k}\;p(x\,|\,c;\theta_{k})$ be the index of the winning hypothesis for input $c$ and target $x$.
+In the vanilla WTA mode, we set $q_{k}(x,c)=\mathbf{1}[k=k^{\star}(x,c)]$, which directly optimizes the Oracle NLL Loss (22).
+However, this formulation risks collapse, where some hypotheses are rarely selected and thus under-trained.
+To mitigate this, Rupprecht et al.
+(2017) proposed the relaxed WTA mode:
+$q_{k}(x,c)=(1-\varepsilon)\mathbf{1}\left[k=k^{\star}(x,c)\right]+\frac{\varepsilon}{K-1}\mathbf{1}\left[k\neq k^{\star}(x,c)\right]\;.$
+which gives higher weight to the winning hypothesis while still providing a small gradient to the others, controlled by $\varepsilon>0$.
+Subsequent methods extend this idea by making $q_{k}$ a function of the training step $t$, thereby adjusting the contribution of non-winning hypotheses during learning (Makansi et al., 2019; Narayanan et al., 2021; Nehme et al., 2024; Perera et al., 2024).
+For example, in the annealed MCL method (Perera et al., 2024), a temperature parameter $\uptau$ is introduced and we have:
+$q_{k}(x,c;\uptau)=\frac{p(x\,|\,c;\theta_{k})^{\frac{1}{\uptau}}}{Z_{x,c}(\uptau)}\,,\quad\quad Z_{x,c}=\sum_{s=1}^{K}p(x\,|\,c;\theta_{s})^{\frac{1}{\uptau}}\;,$ (33)
+where the temperature $t\mapsto\uptau(t)$ follows a decreasing schedule, typically $\uptau(t)=\uptau(0)\rho^{t}$ with $\rho<1$ and $\uptau(0)>0$ where $t$ in the training step.
+At high temperatures, training is distributed more evenly across all hypotheses, which helps to prevent collapse.
+As $\uptau\to 0$, the method converges to the greedy WTA setup, which can maximize (oracle) performance provided that all hypotheses have been sufficiently trained.
+In the Audio Captioning experiments, Annealed MCL was trained with $\uptau(0)=1.0$, $\rho=0.999$, and we switched back to vanilla WTA when the temperature reached $10^{-6}$.
+E.4 Decoding methods
+Formally, Maximum-A-Posteriori decoding in the context of language modeling consist, given (fixed) parameters $\theta$ of finding sequences $\hat{x}=(\hat{x}_{1},\dots,\hat{x}_{T})\in\mathcal{V}^{T}$ that maximizes $p_{\theta}(\hat{x}\,|\,c)=p_{\theta}(\hat{x}_{1}\,|\,c)\prod_{t=2}^{T}p_{\theta}(\hat{x}_{t}\,|\,\hat{x}_{<t},c)$ given a context $c\in\mathcal{C}$.
+Because an exhaustive search of the most likely sequence given $c$ and $\theta$ would be intractable, we used the following heuristics.
+Greedy & Beam Search.
+Given a beam size (or beam width) $B$, beam search (Lowerre, 1976) proceeds as follows.
+First, compute the $B$ most likely tokens $\hat{x}_{1}^{1},\dots,\hat{x}_{1}^{B}$ from the distribution $p_{\theta}(\hat{x}_{1}\,|\,c)$.
+Next, run a forward pass for each candidate $\hat{x}_{1}^{i}$ through the language model to obtain $B$ distributions $p_{\theta}(\cdot\,|\,\hat{x}_{1}^{k},c)$ (for $k=1,\dots,B$), each over the vocabulary $\mathcal{V}$, yielding $B\times|\mathcal{V}|$ candidate probabilities.
+From these $B\times|\mathcal{V}|$ values, select the top-$B$ tokens to form the next candidates $\hat{x}_{2}^{1},\dots,\hat{x}_{2}^{B}$, while keeping track of the preceding tokens in each beam.
+Repeat this procedure for $t=1,\dots,T$ to produce $B$ sequences $\hat{x}^{1},\dots,\hat{x}^{B}\in\mathcal{V}^{T}$.
+Greedy search is the special case $B=1$, i.e., at each step only the top-1 token is chosen: $\hat{x}_{t+1}^{1}=\mathrm{argmax}_{x_{t+1}}p_{\theta}(x_{t+1}\,|\,\hat{x}_{<t})$.
+While this procedure is quite effective and reliable in practice, Beam Search is known to yield low diversity, returning candidates that differ only slightly near the ends of their decoding paths when asked for multiple outputs (Vijayakumar et al., 2018).
+Diverse Beam Search.
+Diverse Beam Search (Vijayakumar et al., 2018) is an alternative to Beam Search in which the beam set of size $B$ is partitioned into $G$ disjoint groups of equal size $B^{\prime}=B/G$.
+The goal is to maximize the likelihood within each group while encouraging dissimilarity across groups.
+Let the set of hypotheses for group $g$ at time $t$ be $X_{t}^{g}\triangleq\{\,\hat{x}_{1:t}^{\,b+(g-1)B^{\prime}}\,|\,b=1,\dots,B^{\prime}\,\}$.
+Dissimilarity is measured through a dissimilarity function $\Delta(\cdot,\cdot)$ that measures dissimilarity between a sequence $\hat{x}_{1:t}$ and a group $X_{t}^{g}$.
+In the following, we denote by $\mathrm{Top}\text{-}B^{\prime}(S,f)$ the operator returning the subset of $S$ containing the $B^{\prime}$ highest-scoring elements under $f:x\in S\mapsto\mathbb{R}$.
+Let us also denote by $\mathrm{Expand}(X_{t-1}^{g},\mathcal{V})=\{[x,v]\,|\,v\in\mathcal{V},x\in X_{t-1}^{g}\}$ denote the sequence obtained by appending token $v\in\mathcal{V}$ to a hypothesis in $X_{t-1}^{g}$ at decoding step $t$, with $X_{0}^{g}=\{\emptyset\}$.
+Then, for group $g=1$ a beam search with size $B^{\prime}$ is performed, i.e., $X_{t}^{1}=\mathrm{Top}\text{-}{B^{\prime}}\big(\mathrm{Expand}(X_{t-1}^{1},\mathcal{V})\,,x\in\mathcal{V}^{t}\mapsto\log p_{\theta}(x\,|\,c)\big),$ and for groups $g\in\{2,\dots,G\}$,
+$X_{t}^{g}=\mathrm{Top}\text{-}{B^{\prime}}\big(\mathrm{Expand}(X_{t-1}^{g},\mathcal{V})\,,x\in\mathcal{V}^{t}\mapsto\log p_{\theta}(x\,|\,c)+\lambda\sum_{h=1}^{g-1}\Delta(x,X_{t}^{h})\big)\;.$
+This trades off sequence likelihood and dissimilarity to the previous groups, with diversity penalty $\lambda>0$.
+In practice, the dissimilarity decomposes as $\Delta(\hat{x}_{1:t},X_{t}^{h})=\sum_{u_{1:t}\in X_{t}^{h}}\delta(\hat{x}_{1:t},u_{1:t}),$ where $\delta$ is a pairwise dissimilarity.
+In the Transformers implementation, $\delta$ uses a Hamming penalty at the current step, $\delta(\hat{x}_{1:t},u_{1:t})=\mathds{1}\!\big[\hat{x}_{t}\neq u_{t}\big],$ so tokens already chosen at position $t$ by earlier groups are penalized.
+Test-time Augmentation.
+TTA involves applying to the context $c$ a random perturbation $\mathcal{T}_{\phi}:\mathcal{C}\rightarrow\mathcal{C}$, parameterized by $\phi$ that controls the augmentation strength.
+We generate $K$ perturbed contexts $\tilde{c}_{1},\dots,\tilde{c}_{K}\sim\mathcal{T}_{\phi}(c)$ and perform MAP generation via Beam Search with beam size $\tfrac{B}{K}$ for each perturbed context.
+In our audio captioning experiments, we implement TTA using SpecAugment (Park et al., 2019), where $\phi$ consists of two parameters: the percentage of time and frequency bands masked in the input spectrogram.
+Remarks on the decoding setup for each training method.
+Beam Search was used as the decoding strategy for all three training methods described in Section E.3.
+To ensure a fair comparison under a fixed computational budget (i.e., a comparable number of forward passes), we adjust the beam size depending on the method:
+  • LoRA-MLE with BS or DBS uses a beam of size $B$, where $B\geq K$ if $K$ sequences are to be returned. For comparability, LoRA-MLE with TTA uses a beam size of $B/K$.
+  • LoRA-MCL decodes each of the $K$ hypotheses with a beam of size $B/K$, and each hypothesis returns a single sequence.
+  • LoRA-MoE was evaluated with two decoding strategies:
+    1. 1.
+      MLE Decoding, where LoRA-MoE is treated as a single-hypothesis model with LoRA-MoE layers integrated into the base architecture.
+In this case, MAP decoding with beam size $B$ yields up to $K$ sequences (as in LoRA-MLE).
+    2. 2.
+      Stochastic Router Decoding, following Zuo et al.
+(2022), where the expert index is sampled randomly from the router’s mixing-weight distribution at each layer.
+Each forward pass uses a beam size of $B/K$, returning one sequence.
+Each of the above approaches is presented in the main submission results, except for Stochastic Router, which was omitted for conciseness due to its poor performance.
+For completeness, its results are provided in Tables 8, 10, and 12.
+E.5 Parallelization over the hypotheses in LoRA-MCL
+A naive implementation of MCL for winner selection, as in Section 3.2, may require a loop over the $K$ hypotheses in the batch to determine the winner associated with each index of the batch, which would drastically slow down training.
+To alleviate this issue, we propose the following methodology.
+Let $x\in\mathbb{R}^{\mathcal{B}\times T\times d}$ denote the input of the transformer architecture, where $\mathcal{B}$ is the batch size, $T$ is the total sequence length, and $d$ is the number of features.
+We duplicate $x$, $K$ times along the batch size dimension to get $h=(h^{(1)},\dots,h^{(K)})\in\mathbb{R}^{\mathcal{B}K\times T\times d}$.
+Computation at LoRA layer $\ell$ writes as:
+$\begin{bmatrix}h^{(1)}&\cdots&h^{(K)}\end{bmatrix}\leftarrow\begin{bmatrix}h^{(1)}&\cdots&h^{(K)}\end{bmatrix}\begin{bmatrix}A^{1}_{\ell}B^{1}_{\ell}&0&0&0\\ 0&A^{2}_{\ell}B^{2}_{\ell}&0&0\\ \vdots&\vdots&\ddots&\vdots\\ 0&0&0&A^{K}_{\ell}B^{K}_{\ell}\end{bmatrix}+\begin{bmatrix}h^{(1)}W^{\ell}&\cdots&h^{(K)}W^{\ell}\end{bmatrix}\,$, (34)
+where $W_{\ell}\in\mathbb{R}^{d\times d}$ is the base model (whose parameters remain frozen during training).
+In practice, this computation is conducted using a Group Convolution operation.
+[^2] This can be done by first reshaping $h$ to shape $(\mathcal{B}\times Kd\times T)$ and applying a $n$.Conv1d($Kd$, KdKd, kernel_size=1, groups=$K$) following the PyTorch layer implementation, reshaping back $h$ to shape $(K\mathcal{B}\times T\times d)$ before adding the base model output, and then repeating at the next LoRA unit.
+While this duplication virtually multiplies the batch size by $K$, the memory overhead remains manageable, assuming that $r\ll d$.
+E.6 Audio Captioning Experiments
+E.6.1 Experimental setup
 Architecture and training.
-We used the Instructed version of Qwen-2-Audio [6] as the base model, which features $\sim 8.4$ billion parameters.
+We used the Instructed version of Qwen-2-Audio (Chu et al., 2024) as the base model, which features $\sim 8.4$ billion parameters.
 We trained using bfloat16 precision.
 We used LoRA adapters applied to the $Q,K,V$ linear projections of the attention modules, and the upside and downside projections of the feedforward blocks, for all the transformer blocks, which include both the audio encoder and language model decoder.
-We used a rank $r$, with $r=\alpha=8$ unless otherwise stated, with dropout equal to $0.1$.
-We trained with a batch size of $2$, with AdamW optimizer [40] (with $\beta_{1}=0.9$, and $\beta_{2}=0.98$), weight decay of $0.05$, using a cosine scheduler with minimum learning rate of $10^{-6}$ and maximum learning rate of $10^{-5}$, with a warmup ratio of $0.1$.
+We used a rank $r$, with $r=\alpha=8$ unless otherwise stated, with dropout equal to $0.1$ (enabled during training, and disabled during inference).
+We trained with a batch size of $2$, with AdamW optimizer (Loshchilov and Hutter, 2017) (with $\beta_{1}=0.9$, and $\beta_{2}=0.98$), weight decay of $0.05$, using a cosine scheduler with minimum learning rate of $10^{-6}$ and maximum learning rate of $10^{-5}$, with a warmup ratio of $0.1$.
 Gradient clipping is used with a maximum gradient norm of $1.0$.
-The validation loss was computed once every epoch.
-E.2.2 Additional results
-Figure 5: Quality vs. Diversity on Audio Captioning with $4$ candidates.
-Figure 6: Quality vs. Diversity on Audio Captioning with $3$ candidates.
-Varying the number $K$ of hypotheses.
-Figures 5 and 6 present the same scatter plots as Figure 2, but with 4 and 3 candidates respectively.
-The observed trends remain consistent: LoRA-MCL continues to demonstrate an excellent balance between quality and diversity.
-Notably, the case with $\varepsilon=0$ performs particularly well, which we attribute to improved stability in vanilla WTA when the number of heads is reduced.
-Note that the detailed results with $K=5$ associated with Figure 2 are given in Tables 5 and 6 for Clotho and AudioCaps, respectively.
-Table 5: Results on Clotho-V2 with 5 candidates and MAP Decoding. ‘BS’ and ‘DBS’ stand for beam search and diverse beam search respectively.
-Training Decoding Beam Div1 Div2 mBLEU-4 BLEU-1 BLEU-4 METEOR ROUGE-L sBERT CIDEr-D SPICE SPIDEr
-LoRA-MLE ($r=8$) BS 5 0.298 0.360 0.827 0.651 0.133 0.229 0.443 0.575 0.621 0.175 0.391
-LoRA-MLE ($r=40$) BS 5 0.300 0.365 0.819 0.619 0.105 0.213 0.422 0.559 0.556 0.166 0.354
-LoRA-MLE ($r=8$) BS 10 0.310 0.386 0.787 0.650 0.139 0.233 0.450 0.576 0.634 0.181 0.400
-LoRA-MLE ($r=40$) BS 10 0.314 0.393 0.780 0.623 0.113 0.219 0.426 0.561 0.564 0.170 0.360
-LoRA-MLE ($r=8$) BS 25 0.316 0.402 0.764 0.653 0.143 0.236 0.453 0.580 0.639 0.184 0.404
-LoRA-MLE ($r=40$) BS 25 0.324 0.411 0.750 0.629 0.121 0.222 0.433 0.563 0.580 0.171 0.368
-LoRA-MLE ($r=8$) DBS ($\lambda=0.5$) 5 0.421 0.560 0.524 0.679 0.145 0.241 0.467 0.597 0.674 0.193 0.421
-LoRA-MLE ($r=40$) DBS ($\lambda=0.5$) 5 0.422 0.558 0.524 0.652 0.122 0.224 0.445 0.582 0.601 0.176 0.378
-LoRA-MLE ($r=8$) DBS ($\lambda=0.8$) 5 0.451 0.606 0.446 0.683 0.149 0.242 0.470 0.600 0.678 0.194 0.423
-LoRA-MLE ($r=40$) DBS ($\lambda=0.8$) 5 0.454 0.607 0.444 0.659 0.125 0.227 0.450 0.585 0.612 0.179 0.384
-LoRA-MLE ($r=8$) DBS ($\lambda=1.0$) 5 0.466 0.627 0.413 0.685 0.147 0.243 0.471 0.602 0.683 0.195 0.426
-LoRA-MLE ($r=40$) DBS ($\lambda=1.0$) 5 0.470 0.628 0.414 0.660 0.124 0.228 0.451 0.588 0.614 0.180 0.386
-LoRA-MLE ($r=8$) DBS ($\lambda=0.5$) 10 0.376 0.475 0.722 0.667 0.139 0.238 0.459 0.593 0.656 0.190 0.413
-LoRA-MLE ($r=40$) DBS ($\lambda=0.5$) 10 0.377 0.480 0.716 0.642 0.115 0.223 0.439 0.575 0.589 0.179 0.374
-LoRA-MLE ($r=8$) DBS ($\lambda=0.8$) 10 0.386 0.491 0.708 0.668 0.139 0.238 0.459 0.595 0.652 0.191 0.412
-LoRA-MLE ($r=40$) DBS ($\lambda=0.8$) 10 0.390 0.495 0.704 0.640 0.115 0.223 0.437 0.577 0.585 0.179 0.371
-LoRA-MLE ($r=8$) DBS ($\lambda=1.0$) 10 0.392 0.495 0.704 0.664 0.138 0.240 0.455 0.596 0.655 0.192 0.413
-LoRA-MLE ($r=40$) DBS ($\lambda=1.0$) 10 0.395 0.498 0.704 0.641 0.112 0.224 0.437 0.578 0.583 0.182 0.372
-LoRA-MLE ($r=8$) DBS ($\lambda=0.5$) 25 0.304 0.369 0.813 0.650 0.134 0.229 0.445 0.574 0.625 0.176 0.393
-LoRA-MLE ($r=40$) DBS ($\lambda=0.5$) 25 0.305 0.370 0.812 0.619 0.106 0.214 0.423 0.558 0.555 0.166 0.354
-LoRA-MLE ($r=8$) DBS ($\lambda=0.8$) 25 0.302 0.366 0.818 0.649 0.133 0.228 0.443 0.574 0.620 0.176 0.391
-LoRA-MLE ($r=40$) DBS ($\lambda=0.8$) 25 0.304 0.369 0.814 0.619 0.106 0.214 0.423 0.558 0.556 0.166 0.354
-LoRA-MLE ($r=8$) DBS ($\lambda=1.0$) 25 0.302 0.365 0.819 0.648 0.133 0.229 0.444 0.574 0.620 0.177 0.391
-LoRA-MLE ($r=40$) DBS ($\lambda=1.0$) 25 0.305 0.369 0.814 0.619 0.105 0.214 0.423 0.558 0.556 0.167 0.355
-LoRA-MCL ($\varepsilon=0.0$) BS 1 0.490 0.633 0.389 0.662 0.127 0.231 0.453 0.597 0.643 0.186 0.401
-LoRA-MCL ($\varepsilon=0.0005$) BS 1 0.464 0.609 0.433 0.672 0.132 0.233 0.459 0.597 0.649 0.185 0.404
-LoRA-MCL ($\varepsilon=0.05$) BS 1 0.471 0.623 0.405 0.677 0.141 0.238 0.463 0.600 0.672 0.191 0.419
-LoRA-MCL ($\varepsilon=0.0$) BS 2 0.477 0.622 0.407 0.675 0.147 0.237 0.468 0.598 0.670 0.191 0.417
-LoRA-MCL ($\varepsilon=0.0005$) BS 2 0.449 0.594 0.453 0.687 0.151 0.243 0.471 0.597 0.679 0.193 0.424
-LoRA-MCL ($\varepsilon=0.05$) BS 2 0.448 0.598 0.450 0.691 0.157 0.246 0.478 0.600 0.700 0.197 0.436
-LoRA-MCL ($\varepsilon=0.0$) BS 5 0.473 0.618 0.420 0.680 0.153 0.242 0.471 0.599 0.689 0.194 0.429
-LoRA-MCL ($\varepsilon=0.0005$) BS 5 0.443 0.585 0.474 0.688 0.154 0.246 0.475 0.598 0.684 0.197 0.429
-LoRA-MCL ($\varepsilon=0.05$) BS 5 0.445 0.592 0.465 0.691 0.165 0.249 0.482 0.601 0.708 0.199 0.442
-Table 6: Results for AudioCaps with 5 candidates and MAP Decoding.
-Training Decoding Beam Div1 Div2 mBLEU-4 BLEU-1 BLEU-4 METEOR ROUGE-L sBERT CIDEr-D SPICE SPIDEr
-LoRA-MLE ($r=8$) BS 5 0.238 0.303 0.881 0.601 0.145 0.304 0.490 0.682 0.582 0.239 0.400
-LoRA-MLE ($r=40$) BS 5 0.232 0.298 0.888 0.576 0.134 0.296 0.472 0.679 0.510 0.234 0.362
-LoRA-MLE ($r=8$) BS 10 0.252 0.331 0.849 0.606 0.156 0.309 0.497 0.686 0.608 0.242 0.414
-LoRA-MLE ($r=40$) BS 10 0.246 0.324 0.861 0.581 0.147 0.300 0.479 0.681 0.520 0.238 0.369
-LoRA-MLE ($r=8$) BS 25 0.263 0.352 0.825 0.604 0.163 0.310 0.500 0.687 0.616 0.245 0.418
-LoRA-MLE ($r=40$) BS 25 0.258 0.350 0.830 0.583 0.156 0.304 0.483 0.685 0.531 0.240 0.373
-LoRA-MLE ($r=8$) DBS ($\lambda=0.2$) 5 0.312 0.454 0.684 0.620 0.166 0.317 0.508 0.697 0.622 0.257 0.421
-LoRA-MLE ($r=40$) DBS ($\lambda=0.2$) 5 0.310 0.462 0.676 0.595 0.157 0.314 0.492 0.697 0.534 0.255 0.376
-LoRA-MLE ($r=8$) DBS ($\lambda=0.5$) 5 0.358 0.545 0.531 0.636 0.188 0.327 0.521 0.707 0.664 0.270 0.445
-LoRA-MLE ($r=40$) DBS ($\lambda=0.5$) 5 0.354 0.551 0.528 0.610 0.183 0.324 0.506 0.707 0.577 0.267 0.400
-LoRA-MLE ($r=8$) DBS ($\lambda=0.8$) 5 0.377 0.576 0.483 0.640 0.193 0.330 0.521 0.708 0.672 0.270 0.447
-LoRA-MLE ($r=40$) DBS ($\lambda=0.8$) 5 0.374 0.585 0.473 0.614 0.185 0.324 0.509 0.708 0.588 0.269 0.405
-LoRA-MLE ($r=8$) DBS ($\lambda=1.0$) 5 0.386 0.590 0.460 0.638 0.190 0.329 0.521 0.709 0.668 0.269 0.444
-LoRA-MLE ($r=40$) DBS ($\lambda=1.0$) 5 0.382 0.598 0.452 0.612 0.181 0.323 0.505 0.709 0.581 0.266 0.401
-LoRA-MLE ($r=8$) DBS ($\lambda=0.5$) 10 0.314 0.439 0.777 0.624 0.180 0.318 0.510 0.699 0.636 0.262 0.431
-LoRA-MLE ($r=40$) DBS ($\lambda=0.5$) 10 0.309 0.440 0.780 0.605 0.170 0.315 0.496 0.699 0.558 0.260 0.390
-LoRA-MLE ($r=8$) DBS ($\lambda=0.8$) 10 0.320 0.448 0.770 0.624 0.176 0.318 0.507 0.700 0.626 0.263 0.426
-LoRA-MLE ($r=40$) DBS ($\lambda=0.8$) 10 0.316 0.451 0.771 0.604 0.167 0.313 0.496 0.698 0.556 0.258 0.388
-LoRA-MLE ($r=8$) DBS ($\lambda=1.0$) 10 0.325 0.455 0.764 0.621 0.175 0.317 0.506 0.702 0.632 0.262 0.427
-LoRA-MLE ($r=40$) DBS ($\lambda=1.0$) 10 0.322 0.456 0.769 0.602 0.165 0.313 0.495 0.699 0.553 0.258 0.386
-LoRA-MLE ($r=8$) DBS ($\lambda=0.5$) 25 0.247 0.319 0.864 0.604 0.152 0.306 0.493 0.682 0.597 0.243 0.408
-LoRA-MLE ($r=40$) DBS ($\lambda=0.5$) 25 0.241 0.315 0.872 0.582 0.141 0.298 0.477 0.683 0.523 0.236 0.370
-LoRA-MLE ($r=8$) DBS ($\lambda=0.8$) 25 0.246 0.316 0.867 0.603 0.148 0.305 0.492 0.682 0.592 0.242 0.406
-LoRA-MLE ($r=40$) DBS ($\lambda=0.8$) 25 0.240 0.311 0.876 0.581 0.138 0.298 0.475 0.682 0.523 0.235 0.369
-LoRA-MLE ($r=8$) DBS ($\lambda=1.0$) 25 0.245 0.314 0.870 0.603 0.150 0.306 0.493 0.682 0.592 0.242 0.406
-LoRA-MLE ($r=40$) DBS ($\lambda=1.0$) 25 0.239 0.310 0.878 0.581 0.140 0.298 0.476 0.682 0.525 0.236 0.370
-LoRA-MCL ($\varepsilon=0.0$) BS 1 0.496 0.707 0.211 0.649 0.174 0.323 0.518 0.702 0.648 0.255 0.428
-LoRA-MCL ($\varepsilon=0.0005$) BS 1 0.450 0.678 0.292 0.654 0.195 0.333 0.530 0.710 0.679 0.272 0.448
-LoRA-MCL ($\varepsilon=0.05$) BS 1 0.438 0.663 0.326 0.640 0.198 0.332 0.517 0.710 0.611 0.272 0.414
-LoRA-MCL ($\varepsilon=0.0$) BS 2 0.459 0.672 0.233 0.650 0.196 0.330 0.529 0.702 0.679 0.262 0.447
-LoRA-MCL ($\varepsilon=0.0005$) BS 2 0.408 0.629 0.348 0.658 0.213 0.336 0.539 0.711 0.716 0.276 0.470
-LoRA-MCL ($\varepsilon=0.05$) BS 2 0.397 0.616 0.384 0.632 0.208 0.336 0.524 0.710 0.629 0.274 0.424
-LoRA-MCL ($\varepsilon=0.0$) BS 5 0.453 0.657 0.254 0.656 0.211 0.331 0.532 0.702 0.697 0.262 0.456
-LoRA-MCL ($\varepsilon=0.0005$) BS 5 0.403 0.615 0.372 0.661 0.229 0.339 0.545 0.712 0.744 0.278 0.484
-LoRA-MCL ($\varepsilon=0.05$) BS 5 0.392 0.599 0.419 0.640 0.221 0.337 0.530 0.710 0.660 0.274 0.442
+The validation loss was computed once at the end of the epoch.
+E.6.2 Ablation on the relaxation parameters in LoRA-MCL
+Effect of $\varepsilon$. Let $\ell_{k}=-\log p\left(x\,|\,c,\theta_{k}\right)$ denote the NLL of hypothesis $k$ for a pair $(c,x)$. Recall that the relaxed MCL loss for such a pair is
+$(1-\varepsilon)\ell_{k}^{\star}+\frac{\varepsilon}{K-1}\sum_{k=1,k\neq k^{\star}}^{K}\ell_{k}=\left(1-\frac{K\varepsilon}{K-1}\right)\ell_{k}^{\star}+\frac{\varepsilon}{K-1}\sum_{k=1}^{K}\ell_{k}\;,$
+where $k^{\star}(x,c)=\mathrm{argmax}_{k}\;p(x\,|\,c;\theta_{k})$ is the winner hypothesis.
+We see that the first term is a force that pushes the hypothesis $k$ toward the winner hypothesis, while the second term assigns equal weight to each hypothesis, pulling them toward the barycenter of the conditional distribution (See Figure 1 in (Perera et al., 2024)).
+The first term vanishes when $\varepsilon=\frac{K-1}{K}$, which provides an upper bound on the value of $\varepsilon$.
+In practice, when choosing $\varepsilon$ for a new task, the smaller the better unless a collapse is observed; we recommend trying $\varepsilon\in[0,0.1]$.
+We run LoRA-MCL (relaxed) with $\varepsilon\in\{0.0005,0.05,0.1,0.3,0.5,0.8\}$ on AudioCaps (AC) and Clotho (CL) in Table 5 with $K=5$ (LoRA-MCL “relaxed” rows).
+We observe that outside the “small” $\varepsilon$ regime ($0.0005,0.05$), increasing $\varepsilon$ significantly degrades the diversity (e.g., mBLEU 4 from $0.410$ with $\varepsilon=0.0005$ on AudioCaps to $0.963$ with $\varepsilon=0.8$ ).
+For quality (as measured by Oracle SPIDEr), performance also tends to degrade outside the small $\varepsilon$ regime.
+This occurs because large $\varepsilon$ diminishes the benefits of MCL training, causing the model to behave increasingly like LoRA-MLE, where the heads eventually become uniformised.
+Effect of the temperature scheduler.
+LoRA-MCL annealed is based on equations (4) and (5).
+The annealed method relies on the theoretical foundation of deterministic annealing (e.g., Rose et al.
+(2002); Perera et al.
+(2024)), which predicts phase transition phenomena in which the hypotheses “split” at specific temperature levels, namely critical temperatures, to explore hierarchically different modes of the conditional distribution.
+Given a fixed number of training steps, the temperature schedule needs to be neither too slow, otherwise the final training step may still occur above the critical temperature, nor too fast, in which cause the model does not fully benefit from annealing and collapse may still occur.
+We provide an ablation on the decay rate in Table 5 (see the annealed rows) for $\rho\in\{0.9,0.995,0.999,0.9999\}$.
+Performance remains good accross most of this range, except for $\rho=0.9999$ where we observe a degradation of diversity and performance.
+We attribute this to the final temperature remaining above the critical temperature.
+Table 5: Ablation on the relaxation parameters in LoRA-MCL.
+$\varepsilon$ is the relaxation parameter, and $\rho$ in the temperature scheduler $\uptau(t)=\rho^{t}\uptau_{0}$, with $\uptau_{0}=1.0$.
+Evaluation on AudioCaps (AC) and Clotho (CL).
+Training Decoding Beam $\mathrm{mBLEU_{4}}$ (AC) $\mathrm{mBLEU_{4}}$ (CL) $\mathrm{SPIDEr}$ (AC) $\mathrm{SPIDEr}$ (CL)
+LoRA-MLE ($r=8$) DBS ($\lambda=0.8$) 5 0.448 0.446 0.662 0.423
+LoRA-MCL ($\varepsilon=0.0005$) BS 5 0.410 0.488 0.706 0.436
+LoRA-MCL ($\varepsilon=0.05$) BS 5 0.491 0.478 0.728 0.434
+LoRA-MCL ($\varepsilon=0.1$) BS 5 0.502 0.526 0.738 0.435
+LoRA-MCL ($\varepsilon=0.3$) BS 5 0.624 0.643 0.700 0.421
+LoRA-MCL ($\varepsilon=0.5$) BS 5 0.766 0.790 0.634 0.390
+LoRA-MCL ($\varepsilon=0.8$) BS 5 0.963 0.952 0.508 0.329
+LoRA-MCL ($\rho=0.9$) BS 5 0.458 0.432 0.701 0.429
+LoRA-MCL ($\rho=0.995$) BS 5 0.422 0.490 0.729 0.438
+LoRA-MCL ($\rho=0.999$) BS 5 0.423 0.478 0.716 0.443
+LoRA-MCL ($\rho=0.9999$) BS 5 0.830 0.947 0.610 0.331
+E.6.3 Comparison of LoRA-MCL with multi-head fine-tuning
+In this section, we compare LoRA-MCL with multi-head MCL.
+Using Qwen-2-Audio for audio captioning, we duplicated the Language model head (“LMHead”, 640M params) $K$ times, froze the rest of the model, and applied MCL training (without LoRA) with both relaxed and annealed variants.
+As an additional variant, showing that parameter count alone is not the only issue, we also duplicated an internal module (the multimodal projector “MMProj”, $\sim$ 5.2M parameters) and trained it with the same scheme.
+As discussed in Section 3.3, multi-head initialization is not trivial.
+Copying the pretrained head yields no initialization diversity, while random reinitialization discards pretrained knowledge.
+Results are displayed in Table 6.
+We find that $(i)$ Random initialization (Init “Random”) degrades quality for both MMProj and LMHead, and in the LMHead case, produces largely unintelligible captions with artificially high diversity.
+$(ii)$ Copying the parameters of the pretrained model (Init “Copy”) produces coherent outputs, but does not meaningfully benefit from MCL training due to amplified collapse risk, yielding significantly lower quality in practice.
+These findings show that multi-head approaches are not competitive with LoRA-MCL in a comparable training setup in the setting of Large language models.
+Table 6: Comparison of LoRA-MCL with Multi-head fine-tuning.
+Comparison is done with $K=5$ hypotheses in AudioCaps, with the same experimental setup as in the paper.
+The “Init” column refer to the initialization technique for the trainable parameters.
+In the Multi-head versions, the trainable weights are either the Language model head (LMHead) or the Multimodal Projector (MMProj).
+The latter are trained with the relaxed variant with $\varepsilon=0.05$.
+Training Init $K$ Decoding Beam $\mathrm{mBLEU_{4}}$ $\mathrm{SPIDEr}$
+Multi-head (LMHead) Copy 5 BS 5 0.489 0.561
+Multi-head (LMHead) Random 5 BS 5 0.001 0.002
+Multi-head (MMProj) Copy 5 BS 5 0.530 0.394
+Multi-head (MMProj) Random 5 BS 5 0.256 0.140
+LoRA-MCL ($\varepsilon=0.05$) LoRA 5 BS 5 0.491 0.728
+LoRA-MCL (annealed) LoRA 5 BS 5 0.423 0.716
+E.6.4 Additional results
 Evaluating with Sampling-based decoding.
-We also report results using sampling-based decoding methods in Tables 7 and 8 for Clotho and AudioCaps.
+We also report results using sampling-based decoding methods in Tables 9 and 11 for Clotho and AudioCaps.
 Specifically, we use Top-$k$ sampling with $k=50$, Top-$p$ sampling with $p=0.95$, and Typical sampling with a threshold of $0.95$.
-For all sampling methods, we apply a repetition penalty of $1.1$, following [29].
+For all sampling methods, we apply a repetition penalty of $1.1$, following (Keskar et al., 2019).
 In these experiments, the temperature was set to $\eta=1.0$.
-We observe that both LoRA-MLE and LoRA-MCL yield significantly higher diversity compared to MAP decoding, albeit at the cost of reduced output quality.
-This trade-off is evident when comparing results in Tables 5 and 6.
-While both methods show an overall improvement in quality on the CIDEr and SPIDEr metrics across datasets, performance on other metrics such as METEOR and SPICE remains relatively comparable.
-In terms of diversity, there does not appear to be a clear trend indicating which method performs best, as the conclusions vary depending on the dataset.
-These findings suggest the need for further evaluation with different values of epsilon to better understand the quality-diversity trade-off in sampling-based decoding.
-Additionally, a more in-depth exploration of how the number of generated hypotheses affects sampling quality—and its implications for test-time inference scaling with LoRA-MCL [85]—is left for future work.
-Table 7: Results for Clotho with 5 candidates and Sampling-based decoding.
-Training Decoding Div1 Div2 mBLEU-4 BLEU-1 BLEU-4 METEOR ROUGE-L sBERT CIDEr-D SPICE SPIDEr
-LoRA-MLE ($r=8$) Top-$k$ sampling 0.635 0.814 0.113 0.585 0.073 0.214 0.400 0.590 0.486 0.176 0.315
-LoRA-MLE ($r=8$) Typical $p$ sampling 0.645 0.817 0.108 0.589 0.077 0.215 0.402 0.587 0.492 0.177 0.319
-LoRA-MLE ($r=8$) Nucleus (Top-$p$) sampling 0.646 0.818 0.106 0.592 0.074 0.214 0.401 0.589 0.489 0.178 0.318
-LoRA-MCL ($\varepsilon=0.0005$) Top-$k$ sampling 0.642 0.816 0.110 0.600 0.079 0.214 0.406 0.588 0.500 0.175 0.322
-LoRA-MCL ($\varepsilon=0.0005$) Typical $p$ sampling 0.648 0.815 0.111 0.595 0.082 0.215 0.409 0.584 0.506 0.175 0.327
-LoRA-MCL ($\varepsilon=0.0005$) Nucleus (Top-$p$) sampling 0.649 0.817 0.107 0.595 0.079 0.214 0.408 0.586 0.504 0.171 0.325
-Table 8: Results for AudioCaps with 5 candidates and Sampling-based decoding.
-Training Decoding Div1 Div2 mBLEU-4 BLEU-1 BLEU-4 METEOR ROUGE-L sBERT CIDEr-D SPICE SPIDEr
-LoRA-MLE ($r=8$) Top-$k$ sampling 0.562 0.810 0.134 0.560 0.115 0.299 0.457 0.699 0.442 0.241 0.318
-LoRA-MLE ($r=8$) Nucleus (Top-$p$) sampling 0.576 0.810 0.141 0.560 0.121 0.299 0.460 0.699 0.450 0.241 0.325
-LoRA-MLE ($r=8$) Typical sampling 0.567 0.809 0.143 0.550 0.115 0.300 0.452 0.697 0.390 0.241 0.294
-LoRA-MCL ($\varepsilon=0.0005$) Top-$k$ sampling 0.583 0.830 0.098 0.582 0.115 0.298 0.470 0.700 0.498 0.239 0.344
-LoRA-MCL ($\varepsilon=0.0005$) Nucleus (Top-$p$) sampling 0.597 0.831 0.098 0.575 0.124 0.298 0.468 0.697 0.497 0.237 0.344
-LoRA-MCL ($\varepsilon=0.0005$) Typical sampling 0.597 0.831 0.098 0.575 0.124 0.298 0.468 0.697 0.497 0.237 0.344
-E.3 Image Captioning Experiments
-E.3.1 Experimental setup
-Table 9: Quality and Diversity Evaluation on TextCaps with $3$ candidates.
-For each of the presented metrics, higher is better ()($\uparrow$)( ↑ ) except for mBLEU-4 ()($\downarrow$)( ↓ ).
+We observe that both LoRA-MLE and LoRA-MCL yield significantly higher diversity than MAP decoding, albeit at the cost of reduced output quality.
+This trade-off is evident in Tables 8 and 10.
+On AudioCaps, LoRA-MCL shows a slight improvement in both quality and diversity compared to LoRA-MLE.
+On Clotho, it provides a small gain in diversity, while quality slightly favors LoRA-MLE.
+These findings highlight the need for further evaluation with different annealing schedules to better characterize the quality–diversity trade-off in sampling-based decoding.
+Moreover, a deeper study of how the number of generated hypotheses influences sampling quality and its implications for test-time inference scaling with LoRA-MCL (Zhao et al., 2025) is left for future work.
+Further comparison with prior work.
+We compare against Zhu et al.
+(2025) a diffusion-based method using retrieval-guided Langevin dynamics (DAC-RLD), which we identify as the strongest published method for diverse audio captioning on Clotho and AudioCaps with publicly available code and checkpoints.
+A VAE-based method from Zhang et al.
+(2024) is also open-sourced (Clotho only) but underperforms DAC-RLD in both quality and diversity according to the reported scores.
+Adversarial training has also been explored (Mei et al., 2022b; Xu et al., 2022), though these methods are outperformed by DAC-RLD, and do not release code.
+We evaluate the pretrained DAC-RLD checkpoints using our setup: 5 candidate captions generated per audio (instead of 50 followed by Minimum Bayes Risk decoding (Tromble et al., 2008) in the original work), oracle sentence-level quality metrics, and mBLEU-4 for diversity.
+For DAC-RLD, we run both with Beam Search and Nucleus sampling, as in the original work.
+Results (Table 7) show that DAC-RLD achieves diversity comparable to LoRA-MCL with nucleus sampling (slightly better or worse depending on the dataset) but at the cost of substantially lower SPIDEr scores than LoRA-MCL on both Clotho and AudioCaps.
+Table 7: Comparison against DAC-RLD (Zhu et al., 2025) in Diverse Audio Captioning. Evaluation on AudioCaps (AC) and Clotho (CL).
+Training Decoding Beam $\mathrm{mBLEU}_{4}$ (AC) $\mathrm{mBLEU}_{4}$ (CL) SPIDEr (AC) SPIDEr (CL)
+DAC-RLD Nucleus ($p=0.95)$ 1 0.157 0.150 0.435 0.244
+DAC-RLD BS 5 0.239 0.215 0.505 0.287
+LoRA-MCL (annealed) Nucleus ($p=0.95)$ 1 0.163 0.098 0.569 0.325
+LoRA-MCL ($\varepsilon=0.05$) BS 5 0.491 0.478 0.728 0.434
+LoRA-MCL ($\rho=0.999$) BS 5 0.423 0.478 0.716 0.443
+Table 8: Results for Clotho with 5 hypotheses and MAP Decoding.
+‘BS’, ‘DBS’, ‘SR’ and ‘TTA’ stand for beam search, diverse beam search, stochastic router, and test-time augmentation respectively.
+We refer to the TTA parameters as $\phi_{i}$, where the strength that increases with $i$; $\phi_{1}=(0.2,0.3)$, $\phi_{2}=(0.4,0.6)$, and $\phi_{3}=(0.6,0.9)$ as the time and frequency proportion mask with SpecAugment (See Apx.
+E.4).
+Training Decoding Beam $\mathrm{Div}_{2}$ $\mathrm{mBLEU}_{4}$ $\mathrm{BLEU}_{1}$ $\mathrm{BLEU}_{4}$ $\mathrm{METEOR}$ $\mathrm{ROUGE}_{L}$ $\mathrm{sBERT}$ $\mathrm{CIDEr}_{D}$ $\mathrm{SPICE}$ $\mathrm{SPIDEr}$
+LoRA-MLE ($r=8$) BS 5 0.365 0.822 0.656 0.137 0.228 0.445 0.575 0.626 0.174 0.394
+LoRA-MLE ($r=40$) BS 5 0.367 0.818 0.643 0.115 0.226 0.435 0.570 0.595 0.172 0.376
+LoRA-MLE ($r=8$) BS 10 0.391 0.783 0.661 0.142 0.236 0.453 0.578 0.646 0.181 0.406
+LoRA-MLE ($r=40$) BS 10 0.397 0.777 0.647 0.127 0.232 0.445 0.573 0.615 0.177 0.388
+LoRA-MLE ($r=8$) BS 25 0.405 0.759 0.658 0.144 0.236 0.455 0.579 0.648 0.182 0.407
+LoRA-MLE ($r=40$) BS 25 0.415 0.746 0.648 0.131 0.234 0.447 0.578 0.625 0.181 0.395
+LoRA-MLE ($r=8$) DBS ($\lambda=0.8$) 5 0.605 0.446 0.686 0.147 0.241 0.471 0.601 0.678 0.193 0.423
+LoRA-MLE ($r=40$) DBS ($\lambda=0.8$) 5 0.613 0.440 0.677 0.140 0.239 0.463 0.599 0.659 0.194 0.414
+LoRA-MLE ($r=8$) DBS ($\lambda=1.0$) 5 0.625 0.417 0.685 0.148 0.242 0.470 0.602 0.681 0.194 0.425
+LoRA-MLE ($r=40$) DBS ($\lambda=1.0$) 5 0.634 0.407 0.678 0.142 0.239 0.463 0.600 0.660 0.193 0.414
+LoRA-MLE ($r=8$) DBS ($\lambda=0.8$) 10 0.487 0.712 0.670 0.143 0.238 0.460 0.594 0.656 0.191 0.414
+LoRA-MLE ($r=40$) DBS ($\lambda=0.8$) 10 0.499 0.694 0.661 0.132 0.235 0.448 0.590 0.634 0.187 0.401
+LoRA-MLE ($r=8$) DBS ($\lambda=1.0$) 10 0.491 0.708 0.671 0.143 0.238 0.456 0.595 0.662 0.192 0.417
+LoRA-MLE ($r=40$) DBS ($\lambda=1.0$) 10 0.501 0.696 0.659 0.131 0.234 0.444 0.591 0.629 0.188 0.397
+LoRA-MLE ($r=8$) DBS ($\lambda=0.8$) 25 0.368 0.818 0.653 0.135 0.228 0.444 0.575 0.626 0.176 0.394
+LoRA-MLE ($r=40$) DBS ($\lambda=0.8$) 25 0.374 0.811 0.644 0.115 0.226 0.435 0.571 0.594 0.174 0.377
+LoRA-MLE ($r=8$) DBS ($\lambda=1.0$) 25 0.368 0.818 0.652 0.134 0.228 0.443 0.574 0.622 0.174 0.392
+LoRA-MLE ($r=40$) DBS ($\lambda=1.0$) 25 0.372 0.812 0.643 0.115 0.226 0.435 0.571 0.592 0.174 0.376
+LoRA-MLE ($r=8$) TTA BS ($\phi_{1}$) 1 0.443 0.699 0.652 0.114 0.225 0.440 0.581 0.608 0.174 0.383
+LoRA-MLE ($r=8$) TTA BS ($\phi_{2}$) 1 0.540 0.558 0.663 0.130 0.230 0.453 0.588 0.634 0.179 0.397
+LoRA-MLE ($r=8$) TTA BS ($\phi_{3}$) 1 0.642 0.404 0.653 0.118 0.225 0.441 0.581 0.596 0.174 0.376
+LoRA-MLE ($r=8$) TTA BS ($\phi_{1}$) 5 0.412 0.745 0.655 0.137 0.232 0.452 0.580 0.637 0.181 0.402
+LoRA-MLE ($r=8$) TTA BS ($\phi_{2}$) 5 0.516 0.593 0.681 0.156 0.243 0.470 0.591 0.685 0.194 0.430
+LoRA-MLE ($r=8$) TTA BS ($\phi_{3}$) 5 0.619 0.445 0.675 0.148 0.236 0.463 0.586 0.648 0.186 0.407
+LoRA-MoE BS 5 0.363 0.823 0.650 0.131 0.228 0.443 0.571 0.614 0.173 0.387
+LoRA-MoE BS 10 0.395 0.783 0.659 0.137 0.236 0.452 0.579 0.643 0.181 0.405
+LoRA-MoE BS 25 0.408 0.757 0.657 0.140 0.237 0.454 0.579 0.643 0.184 0.405
+LoRA-MoE DBS ($\lambda=0.8$) 5 0.611 0.441 0.682 0.143 0.241 0.468 0.599 0.668 0.194 0.418
+LoRA-MoE DBS ($\lambda=1.0$) 5 0.630 0.410 0.682 0.147 0.241 0.467 0.600 0.675 0.195 0.422
+LoRA-MoE DBS ($\lambda=0.8$) 10 0.490 0.706 0.670 0.142 0.238 0.459 0.593 0.662 0.192 0.416
+LoRA-MoE DBS ($\lambda=1.0$) 10 0.494 0.705 0.668 0.138 0.237 0.456 0.592 0.650 0.191 0.410
+LoRA-MoE DBS ($\lambda=0.8$) 25 0.370 0.814 0.650 0.128 0.228 0.445 0.572 0.613 0.175 0.386
+LoRA-MoE DBS ($\lambda=1.0$) 25 0.370 0.814 0.651 0.129 0.229 0.445 0.573 0.620 0.174 0.390
+LoRA-MoE SR BS 1 0.308 0.878 0.604 0.090 0.203 0.406 0.554 0.515 0.153 0.330
+LoRA-MoE SR BS 2 0.317 0.868 0.629 0.116 0.217 0.427 0.564 0.576 0.163 0.364
+LoRA-MoE SR BS 5 0.309 0.880 0.620 0.109 0.217 0.424 0.562 0.561 0.162 0.357
+LoRA-MCL ($\varepsilon=0.0005$) BS 1 0.605 0.440 0.675 0.135 0.233 0.458 0.597 0.654 0.187 0.408
+LoRA-MCL ($\varepsilon=0.05$) BS 1 0.617 0.415 0.680 0.130 0.239 0.463 0.601 0.662 0.190 0.414
+LoRA-MCL (annealed) BS 1 0.621 0.415 0.673 0.131 0.237 0.458 0.599 0.665 0.189 0.415
+LoRA-MCL ($\varepsilon=0.0005$) BS 2 0.591 0.461 0.688 0.154 0.242 0.472 0.598 0.688 0.192 0.428
+LoRA-MCL ($\varepsilon=0.05$) BS 2 0.593 0.452 0.692 0.160 0.247 0.481 0.599 0.694 0.193 0.431
+LoRA-MCL (annealed) BS 2 0.595 0.456 0.687 0.158 0.245 0.472 0.599 0.698 0.196 0.435
+LoRA-MCL ($\varepsilon=0.0005$) BS 5 0.581 0.488 0.687 0.156 0.244 0.476 0.599 0.700 0.196 0.436
+LoRA-MCL ($\varepsilon=0.05$) BS 5 0.581 0.478 0.689 0.162 0.249 0.480 0.599 0.697 0.196 0.434
+LoRA-MCL (annealed) BS 5 0.584 0.478 0.689 0.168 0.246 0.477 0.600 0.711 0.199 0.443
+Table 9: Results for Clotho with 5 hypotheses and Sampling-based Decoding.
+Training Decoding $\mathrm{Div}_{2}$ $\mathrm{mBLEU}_{4}$ $\mathrm{BLEU}_{4}$ $\mathrm{METEOR}$ $\mathrm{ROUGE}_{L}$ $\mathrm{sBERT}$ $\mathrm{CIDEr}_{D}$ $\mathrm{SPICE}$ $\mathrm{SPIDEr}$
+LoRA-MLE ($r=8$) Top-$k$ sampling 0.813 0.109 0.066 0.215 0.402 0.593 0.497 0.176 0.323
+LoRA-MLE ($r=8$) Nucleus (Top-$p$) sampling 0.812 0.112 0.073 0.212 0.403 0.586 0.516 0.173 0.330
+LoRA-MLE ($r=8$) Typical $p$ sampling 0.812 0.111 0.073 0.212 0.403 0.586 0.516 0.173 0.331
+LoRA-MCL (annealed) Top-$k$ sampling 0.819 0.097 0.074 0.212 0.403 0.597 0.507 0.175 0.327
+LoRA-MCL (annealed) Nucleus (Top-$p$) sampling 0.820 0.098 0.074 0.212 0.403 0.589 0.507 0.171 0.325
+LoRA-MCL (annealed) Typical $p$ sampling 0.819 0.103 0.072 0.211 0.402 0.590 0.509 0.172 0.327
+Table 10: Results for AudioCaps with 5 hypotheses and MAP Decoding.
+Training Decoding Beam $\mathrm{Div}_{2}$ $\mathrm{mBLEU}_{4}$ $\mathrm{BLEU}_{4}$ $\mathrm{METEOR}$ $\mathrm{ROUGE}_{L}$ $\mathrm{sBERT}$ $\mathrm{CIDEr}_{D}$ $\mathrm{SPICE}$ $\mathrm{SPIDEr}$
+LoRA-MLE ($r=8$) BS 5 0.395 0.764 0.267 0.377 0.606 0.700 1.121 0.250 0.668
+LoRA-MLE ($r=40$) BS 5 0.392 0.773 0.280 0.382 0.606 0.701 1.144 0.251 0.681
+LoRA-MLE ($r=8$) BS 10 0.410 0.746 0.286 0.385 0.610 0.704 1.157 0.256 0.689
+LoRA-MLE ($r=40$) BS 10 0.407 0.747 0.298 0.382 0.610 0.704 1.151 0.255 0.686
+LoRA-MLE ($r=8$) BS 25 0.417 0.732 0.281 0.383 0.609 0.706 1.144 0.258 0.683
+LoRA-MLE ($r=40$) BS 25 0.415 0.735 0.289 0.384 0.611 0.706 1.152 0.260 0.688
+LoRA-MLE ($r=8$) DBS ($\lambda=0.8$) 5 0.553 0.448 0.268 0.378 0.610 0.708 1.117 0.248 0.662
+LoRA-MLE ($r=40$) DBS ($\lambda=0.8$) 5 0.557 0.444 0.263 0.380 0.612 0.707 1.130 0.248 0.669
+LoRA-MLE ($r=8$) DBS ($\lambda=1.0$) 5 0.580 0.403 0.275 0.376 0.612 0.708 1.128 0.249 0.667
+LoRA-MLE ($r=40$) DBS ($\lambda=1.0$) 5 0.580 0.403 0.268 0.378 0.614 0.709 1.138 0.250 0.672
+LoRA-MLE ($r=8$) DBS ($\lambda=0.8$) 10 0.481 0.684 0.274 0.373 0.606 0.709 1.114 0.259 0.665
+LoRA-MLE ($r=40$) DBS ($\lambda=0.8$) 10 0.481 0.683 0.278 0.380 0.610 0.710 1.124 0.257 0.670
+LoRA-MLE ($r=8$) DBS ($\lambda=1.0$) 10 0.486 0.678 0.276 0.372 0.607 0.710 1.115 0.259 0.665
+LoRA-MLE ($r=40$) DBS ($\lambda=1.0$) 10 0.492 0.675 0.276 0.376 0.609 0.709 1.118 0.256 0.666
+LoRA-MLE ($r=8$) DBS ($\lambda=0.8$) 25 0.402 0.756 0.273 0.376 0.607 0.703 1.129 0.251 0.673
+LoRA-MLE ($r=40$) DBS ($\lambda=0.8$) 25 0.399 0.763 0.282 0.381 0.605 0.702 1.139 0.253 0.680
+LoRA-MLE ($r=8$) DBS ($\lambda=1.0$) 25 0.402 0.757 0.271 0.377 0.608 0.703 1.128 0.251 0.672
+LoRA-MLE ($r=40$) DBS ($\lambda=1.0$) 25 0.398 0.765 0.282 0.381 0.606 0.701 1.136 0.252 0.678
+LoRA-MLE ($r=8$) TTA BS ($\phi_{1}$) 1 0.385 0.731 0.198 0.343 0.574 0.693 0.969 0.226 0.584
+LoRA-MLE ($r=8$) TTA BS ($\phi_{2}$) 1 0.479 0.588 0.207 0.352 0.586 0.695 0.992 0.231 0.597
+LoRA-MLE ($r=8$) TTA BS ($\phi_{3}$) 1 0.576 0.438 0.185 0.332 0.567 0.683 0.916 0.218 0.550
+LoRA-MLE ($r=8$) TTA BS ($\phi_{1}$) 5 0.395 0.733 0.241 0.358 0.592 0.700 1.057 0.242 0.636
+LoRA-MLE ($r=8$) TTA BS ($\phi_{2}$) 5 0.491 0.590 0.260 0.360 0.597 0.703 1.047 0.246 0.630
+LoRA-MLE ($r=8$) TTA BS ($\phi_{3}$) 5 0.596 0.435 0.251 0.347 0.586 0.693 1.005 0.235 0.605
+LoRA-MoE BS 5 0.396 0.766 0.274 0.381 0.608 0.702 1.129 0.252 0.674
+LoRA-MoE BS 10 0.411 0.746 0.288 0.385 0.613 0.703 1.161 0.256 0.692
+LoRA-MoE BS 25 0.415 0.736 0.287 0.385 0.612 0.705 1.154 0.258 0.689
+LoRA-MoE DBS ($\lambda=0.8$) 5 0.555 0.443 0.275 0.379 0.611 0.707 1.136 0.248 0.670
+LoRA-MoE DBS ($\lambda=1.0$) 5 0.578 0.409 0.268 0.377 0.609 0.708 1.130 0.247 0.667
+LoRA-MoE DBS ($\lambda=0.8$) 10 0.484 0.677 0.279 0.374 0.608 0.709 1.122 0.256 0.667
+LoRA-MoE DBS ($\lambda=1.0$) 10 0.488 0.675 0.283 0.374 0.608 0.710 1.119 0.257 0.666
+LoRA-MoE DBS ($\lambda=0.8$) 25 0.402 0.757 0.275 0.382 0.609 0.701 1.135 0.253 0.676
+LoRA-MoE DBS ($\lambda=1.0$) 25 0.401 0.759 0.275 0.382 0.609 0.701 1.133 0.253 0.675
+LoRA-MoE SR BS 1 0.256 0.909 0.165 0.318 0.538 0.670 0.865 0.200 0.526
+LoRA-MoE SR BS 2 0.269 0.895 0.195 0.326 0.549 0.675 0.905 0.210 0.551
+LoRA-MoE SR BS 5 0.261 0.907 0.196 0.331 0.550 0.678 0.925 0.216 0.563
+LoRA-MCL ($\varepsilon=0.0005$) BS 1 0.580 0.374 0.259 0.377 0.613 0.711 1.119 0.248 0.662
+LoRA-MCL ($\varepsilon=0.05$) BS 1 0.551 0.427 0.273 0.382 0.622 0.712 1.181 0.253 0.695
+LoRA-MCL (annealed) BS 1 0.570 0.397 0.275 0.379 0.615 0.713 1.149 0.255 0.680
+LoRA-MCL ($\varepsilon=0.0005$) BS 2 0.575 0.401 0.277 0.385 0.626 0.712 1.152 0.258 0.684
+LoRA-MCL ($\varepsilon=0.05$) BS 2 0.544 0.464 0.298 0.394 0.631 0.712 1.209 0.260 0.714
+LoRA-MCL (annealed) BS 2 0.574 0.411 0.309 0.392 0.632 0.714 1.205 0.264 0.712
+LoRA-MCL ($\varepsilon=0.0005$) BS 5 0.579 0.410 0.306 0.392 0.631 0.712 1.190 0.263 0.706
+LoRA-MCL ($\varepsilon=0.05$) BS 5 0.542 0.491 0.309 0.399 0.636 0.715 1.237 0.265 0.728
+LoRA-MCL (annealed) BS 5 0.575 0.423 0.315 0.398 0.636 0.714 1.211 0.268 0.716
+Table 11: Results for AudioCaps with 5 hypotheses and Sampling-based Decoding.
+Training Decoding $\mathrm{Div}_{2}$ $\mathrm{mBLEU}_{4}$ $\mathrm{BLEU}_{4}$ $\mathrm{METEOR}$ $\mathrm{ROUGE}_{L}$ $\mathrm{sBERT}$ $\mathrm{CIDEr}_{D}$ $\mathrm{SPICE}$ $\mathrm{SPIDEr}$
+LoRA-MLE ($r=8$) Top-$k$ sampling 0.711 0.173 0.169 0.335 0.549 0.701 0.906 0.229 0.543
+LoRA-MLE ($r=8$) Nucleus (Top-$p$) sampling 0.697 0.193 0.180 0.344 0.564 0.704 0.940 0.234 0.563
+LoRA-MLE ($r=8$) Typical $p$ sampling 0.699 0.189 0.182 0.343 0.564 0.704 0.945 0.233 0.566
+LoRA-MCL (annealed) Top-$k$ sampling 0.735 0.149 0.175 0.342 0.568 0.705 0.938 0.233 0.563
+LoRA-MCL (annealed) Nucleus (Top-$p$) sampling 0.724 0.163 0.182 0.340 0.566 0.702 0.950 0.235 0.569
+LoRA-MCL (annealed) Typical $p$ sampling 0.724 0.160 0.186 0.341 0.568 0.703 0.953 0.235 0.570
+E.6.5 Qualitative Examples
+We provide some qualitative examples of the predictions on AudioCaps. Here, LoRA-MCL uses $\varepsilon=0.05$, Beam Search with $B=5$ and $K=5$, LoRA-MoE uses Beam search with $B=25$.
+Example 1. References:
+  • A large truck driving by as an emergency siren wails and truck horn honks
+  • A wailing siren fades, a motor sputters, then the siren resumes accompanied by blaring horns
+  • An emergency siren ringing with car horn honking
+  • A fire truck engine runs and the siren is blowing but stops, traffic is present, the fire truck horn honks twice, and the siren begins again
+  • A fire engine with a siren fading then another loud siren
+LoRA-MoE.
+An emergency vehicle passes by and blows its horn followed by the siren of an emergency vehicle}\}
+[…]
+[…]
+[…]
+[…]
+LoRA-MCL.
+An emergency vehicle siren is going off}\}
+[…]
+[…]
+[…]
+[…]
+Example 2. References:
+  • A man speaks as birds chirp and dogs bark
+  • A man is speaking as birds are squawking, and a dog barks
+  • A man talks while several animals make noises in the background
+  • A man speaking followed by dogs barking alongside chimps screaming and birds chirping
+  • A man speaking as monkeys scream and dogs bark followed by birds cawing in the distance
+LoRA-MoE.
+A man is speaking and dogs are barking}\}
+[…]
+[…]
+[…]
+[…]
+LoRA-MCL.
+A man speaking and dogs barking}\}
+[…]
+[…]
+[…]
+[…]
+Example 3. References:
+  • An engine running and wind with various speech in the background
+  • A motorboat engine operating as a crowd of people talk followed by metal creaking and a man speaking
+  • A large motor is running smoothly, water is splashing, people are talking in the background, and an adult male speaks in the distance
+  • A ship engine running as a crowd of people talk followed by a ship hull creaking as wind blows heavily into a microphone
+  • Outdoor noise from a water vehicle as people are talking
+LoRA-MoE.
+Humming of an idling engine followed by a horn honking}\}
+[…]
+[…]
+[…]
+[…]
+LoRA-MCL.
+A boat motor is running and people are talking in the background}\}
+[…]
+[…]
+[…]
+[…]
+E.7 Image Captioning Experiments
+E.7.1 Experimental setup
+Table 12: Quality and Diversity Evaluation on TextCaps with 33 candidates.
+For each of the presented metrics, higher is better $(\uparrow)$ except for mBLEU-4 $(\downarrow)$.
 LoRA-MCL is trained with $\varepsilon=0.1$, $r=8$ and $\alpha=32$.
@@
 For completeness, we also trained LoRA-MLE with $r=8$ and $\alpha=32$ in the rows marked with .
-Training Decoding Beam Div1 Div2 mBLEU-4 BLEU-1 BLEU-4 METEOR ROUGE-L sBERT CIDEr-D SPICE SPIDEr
-LoRA-MLE Beam Search 3 0.433 0.509 0.688 0.802 0.318 0.315 0.580 0.670 1.517 0.244 0.873
-LoRA-MLE Beam Search 3 0.383 0.421 0.833 0.788 0.315 0.317 0.573 0.670 1.517 0.241 0.874
-LoRA-MLE Beam Search 6 0.402 0.457 0.786 0.795 0.338 0.326 0.583 0.671 1.557 0.246 0.895
-LoRA-MLE Beam Search 6 0.401 0.456 0.784 0.796 0.339 0.327 0.585 0.672 1.572 0.248 0.903
-LoRA-MLE DBS ($\lambda=0.5$) 3 0.489 0.600 0.529 0.818 0.345 0.325 0.596 0.684 1.571 0.253 0.902
-LoRA-MLE DBS ($\lambda=0.8$) 3 0.530 0.655 0.437 0.824 0.349 0.327 0.601 0.686 1.590 0.251 0.909
-LoRA-MLE DBS ($\lambda=1.0$) 3 0.540 0.669 0.416 0.822 0.348 0.326 0.599 0.685 1.586 0.250 0.906
-LoRA-MLE DBS ($\lambda=0.5$) 3 0.485 0.597 0.531 0.821 0.346 0.326 0.596 0.685 1.589 0.255 0.912
-LoRA-MLE DBS ($\lambda=0.8$) 3 0.485 0.597 0.531 0.821 0.346 0.326 0.596 0.685 1.589 0.255 0.912
-LoRA-MLE DBS ($\lambda=1.0$) 3 0.535 0.665 0.425 0.827 0.357 0.327 0.601 0.686 1.601 0.252 0.915
-LoRA-MLE DBS ($\lambda=0.5$) 6 0.455 0.532 0.694 0.813 0.344 0.328 0.595 0.681 1.580 0.253 0.908
-LoRA-MLE DBS ($\lambda=0.8$) 6 0.470 0.549 0.671 0.812 0.341 0.328 0.593 0.681 1.573 0.251 0.903
-LoRA-MLE DBS ($\lambda=1.0$) 6 0.474 0.553 0.666 0.812 0.340 0.328 0.592 0.680 1.577 0.250 0.904
-LoRA-MLE DBS ($\lambda=0.5$) 6 0.451 0.528 0.697 0.808 0.343 0.329 0.594 0.681 1.586 0.253 0.911
-LoRA-MLE DBS ($\lambda=0.8$) 6 0.464 0.542 0.681 0.810 0.340 0.329 0.593 0.680 1.583 0.253 0.909
-LoRA-MLE DBS ($\lambda=1.0$) 6 0.473 0.551 0.671 0.810 0.341 0.328 0.592 0.680 1.584 0.251 0.908
-LoRA-MCL Beam Search 1 0.499 0.599 0.520 0.828 0.344 0.330 0.597 0.690 1.674 0.255 0.955
-LoRA-MCL Beam Search 2 0.507 0.618 0.490 0.824 0.360 0.333 0.604 0.687 1.627 0.258 0.932
-We used LLaVA 1.6 as the base model, which features 7.1 billion parameters.
+PaliGemma-3B rows are shown in gray, as this model was fully fine-tuned and is therefore not directly comparable to LoRA-based methods.
+Training Decoding Beam $\mathrm{Div}_{2}$ $\mathrm{mBLEU}_{4}$ $\mathrm{BLEU}_{1}$ $\mathrm{BLEU}_{4}$ $\mathrm{METEOR}$ $\mathrm{ROUGE}_{L}$ $\mathrm{sBERT}$ $\mathrm{CIDEr}_{D}$ $\mathrm{SPICE}$ $\mathrm{SPIDEr}$
+LoRA-MLE BS 3 0.509 0.688 0.802 0.318 0.315 0.580 0.670 1.517 0.244 0.873
+LoRA-MLE BS 6 0.457 0.786 0.795 0.338 0.326 0.583 0.671 1.557 0.246 0.895
+LoRA-MLE BS 3 0.421 0.833 0.788 0.315 0.317 0.573 0.670 1.517 0.241 0.874
+LoRA-MLE BS 6 0.456 0.784 0.796 0.339 0.327 0.585 0.672 1.572 0.248 0.903
+LoRA-MLE DBS ($\lambda=0.5$) 3 0.600 0.529 0.818 0.345 0.325 0.596 0.684 1.571 0.253 0.902
+LoRA-MLE DBS ($\lambda=0.8$) 3 0.655 0.437 0.824 0.349 0.327 0.601 0.686 1.590 0.251 0.909
+LoRA-MLE DBS ($\lambda=1.0$) 3 0.669 0.416 0.822 0.348 0.326 0.599 0.685 1.586 0.250 0.906
+LoRA-MLE DBS ($\lambda=0.5$) 6 0.532 0.694 0.813 0.344 0.328 0.595 0.681 1.580 0.253 0.908
+LoRA-MLE DBS ($\lambda=0.8$) 6 0.549 0.671 0.812 0.341 0.328 0.593 0.681 1.573 0.251 0.903
+LoRA-MLE DBS ($\lambda=1.0$) 6 0.553 0.666 0.812 0.340 0.328 0.592 0.680 1.577 0.250 0.904
+LoRA-MLE DBS ($\lambda=0.5$) 3 0.597 0.531 0.821 0.346 0.326 0.596 0.685 1.589 0.255 0.912
+LoRA-MLE DBS ($\lambda=0.8$) 3 0.597 0.531 0.821 0.346 0.326 0.596 0.685 1.589 0.255 0.912
+LoRA-MLE DBS ($\lambda=1.0$) 3 0.665 0.425 0.827 0.357 0.327 0.601 0.686 1.601 0.252 0.915
+LoRA-MLE DBS ($\lambda=0.5$) 6 0.528 0.697 0.808 0.343 0.329 0.594 0.681 1.586 0.253 0.911
+LoRA-MLE DBS ($\lambda=0.8$) 6 0.542 0.681 0.810 0.340 0.329 0.593 0.680 1.583 0.253 0.909
+LoRA-MLE DBS ($\lambda=1.0$) 6 0.551 0.671 0.810 0.341 0.328 0.592 0.680 1.584 0.251 0.908
+LoRA-MoE DBS ($\lambda=0.5$) 3 0.600 0.529 0.822 0.347 0.327 0.598 0.684 1.610 0.258 0.923
+LoRA-MoE DBS ($\lambda=0.8$) 3 0.654 0.441 0.828 0.353 0.327 0.603 0.685 1.616 0.254 0.924
+LoRA-MoE DBS ($\lambda=1.0$) 3 0.666 0.421 0.828 0.354 0.328 0.602 0.685 1.622 0.253 0.926
+LoRA-MoE DBS ($\lambda=0.5$) 6 0.530 0.698 0.814 0.348 0.331 0.595 0.681 1.607 0.257 0.923
+LoRA-MoE DBS ($\lambda=0.8$) 6 0.545 0.678 0.813 0.349 0.330 0.596 0.680 1.608 0.255 0.922
+LoRA-MoE DBS ($\lambda=1.0$) 6 0.551 0.670 0.813 0.346 0.330 0.596 0.679 1.607 0.254 0.921
+LoRA-MoE SR BS 1 0.556 0.597 0.809 0.315 0.311 0.576 0.678 1.541 0.243 0.883
+PaliGemma-3B (ft) DBS ($\lambda=0.5$) 3 0.595 0.532 0.832 0.369 0.330 0.608 0.688 1.658 0.255 0.947
+PaliGemma-3B (ft) DBS ($\lambda=0.8$) 3 0.635 0.467 0.835 0.374 0.332 0.611 0.689 1.661 0.257 0.949
+PaliGemma-3B (ft) DBS ($\lambda=1.0$) 3 0.655 0.439 0.833 0.368 0.332 0.609 0.687 1.654 0.256 0.944
+LoRA-MCL BS 1 0.599 0.520 0.828 0.344 0.330 0.597 0.690 1.674 0.255 0.955
+LoRA-MCL BS 2 0.618 0.490 0.824 0.360 0.333 0.604 0.687 1.627 0.258 0.932
+We used LLaVA 1.6 with Vicuna-7B (Zheng et al., 2023) for the LLM, as the base model, which features 7.1 billion parameters.
+We used the official codebase for the implementation.
 We trained using bfloat16 precision.
 We used LoRA adapters applied to the $Q,K,V$, up and down linear projections of each block of the language model.
-We used a rank $r$, with $r=8$ and $\alpha=32$ unless otherwise stated, with dropout equal to $0.1$.
-We trained each model with a batch size of $8$, with AdamW optimizer [40] (with $\beta_{1}=0.9$, and $\beta_{2}=0.98$), weight decay of $0.05$, using a cosine scheduler with maximum learning rate of $10^{-4}$, with a warmup ratio of $0.1$.
+Fine-tuning hyperparameters follow those from the recipe provided by the authors, except that we used smaller batch sizes and a lower LoRA rank due to compute constraints.
+We used a rank $r$, with $r=8$ and $\alpha=32$ unless otherwise stated, with dropout equal to $0.1$ (enabled during training, and disabled during inference).
+We trained each model with a batch size of $8$, with Adam optimizer (Diederik and Jimmy, 2014) (with $\beta_{1}=0.9$, and $\beta_{2}=0.999$), using a cosine scheduler with maximum learning rate of $2\times 10^{-4}$, with a warmup ratio of $0.03$.
+Maximum sequence length is set to $2048$.
 Gradient clipping is used with a maximum gradient norm of $1.0$.
 We used 1 epoch for training, where we duplicated the image as many times as the number of its captions, such that the model sees exactly one time each caption.
-E.4 Additional results
-Table 9 presents the results of LoRA-MCL and LoRA-MLE.
-Almost all the metrics show the same trend in which LoRA-MCL outperforms LoRA-MLE with Beam Search (BS) and Diverse Beam Search (DBS) in terms of quality, although DBS produces more varied outputs.
-Depending on the rank ($r=8$ or $r=24$), we found that setting $\lambda=1.0$ or $\lambda=0.8$, respectively, yields the best quality scores for DBS.
-Similar to our experiments on Audio Captioning, increasing the rank of LoRA-MLE results in a slight degradation in performance but improves diversity.
+E.7.2 Additional results
+Comparing with PaliGemma-3B.
+Several works have explored improving caption diversity in image captioning (Wang and Chan, 2019; Wang et al., 2020; Mahajan and Roth, 2020).
+Among those evaluating on TextCaps, Zhang et al.
+(2022) and Xu et al.
+(2021) report results, but their performance is far below that of recent VLMs, such as PaliGemma-3B (Beyer et al., 2024).
+For example, the reported best corpus-level CIDEr on the TextCaps validation set are: 76.6 for Zhang et al.
+(2022), 95.5 for Xu et al.
+(2021), 127.48 for PaliGemma-3B (224x224).
+We therefore compare with PaliGemma-3B, which provides a publicly available fine-tuned version on Hugging Face.
+We evaluate it using DBS returning $K=3$ hypotheses (Table 12).
+PaliGemma is a strong baseline, arguably the state-of-the-art open-weight model on TextCaps.
+Note, however, that the comparison is not entirely fair, as PaliGemma undergoes full-weight fine-tuning (instead of LoRA).
+Despite this, our method matches and even slightly improves its performance (Oracle SPIDEr of $0.955$ vs.
+$0.949$).
+We also expect that applying LoRA-MCL to PaliGemma could further improve the results.
+The full results presented in Table 12 show that LoRA-MCL tends to outperform LoRA-MLE with Beam Search (BS) and Diverse Beam Search (DBS) in terms of quality, although DBS produces more varied outputs.
+Depending on the rank ($r=8$ or $r=24$), we found that setting $\lambda$ to $1.0$ or $\lambda=0.8$, respectively, yields the best quality scores for DBS.
+Similar to our experiments on Audio Captioning, increasing the rank of LoRA-MLE results in a slight degradation in performance while improving diversity.
 Additionally, increasing the number of beams in BS with LoRA-MLE results in slightly improved performance but reduced diversity.
-In contrast, increasing the number of beams in DBS with LoRA-MLE (with fixed $\lambda$) leads to declines in both quality and diversity.
+In contrast, increasing the number of beams in DBS with LoRA-MLE (with $\lambda$ fixed) leads to declines in both quality and diversity.
 Interestingly, with LoRA-MCL, increasing the number of beams enhances both performance and diversity here.
-E.4.1 Artificial multilingual dataset creation
-To evaluate the behaviour of LoRA-MCL under a multi-modal distribution, we simulated an artificial bi-modal dataset by automatically translating half of the captions from English to French using T5-small [57], while keeping the prompts in English.
+E.7.3 Artificial multilingual dataset creation
+To evaluate the behaviour of LoRA-MCL under a multi-modal distribution, we simulated an artificial bi-modal dataset by automatically translating half of the captions from English to French using T5-small (Raffel et al., 2020), while keeping the prompts in English.
 More specifically, we randomly sampled half of the images and translated their five associated captions.
@@
 During evaluation, to assess which head is considered as the winner (for the head specialization analysis), we selected the one that maximizes the SPIDEr score over the references of the given sample.
-E.5 Computation details
+E.7.4 Specialization of the hypotheses
+Understanding what each hypotheses learn is difficult.
+The founding paper of MCL by Lee et al.
+(2016) already introduced the notion of specialization of the hypotheses in the context of classification (see Lee et al.
+(2016) in Figure 4).
+We provide additional insights with the French/English controlled experiment of Section 6, where we observed some “unsupervised” specialization of the hypotheses; one of the hypothesis learned English, and the other learned French.
+We tried to visualize the hypotheses predictions to understand what the hypotheses have learned in Figure 6.
+To quantify specialization, we embed every generated caption on the evaluation set (for hypotheses $k\in\{1,\dots,K\}$ and each example $i\in\{1,\dots,N\}$) using a Sentence-BERT model (StyleDistance (Patel et al., 2025)).
+We trained a linear Support Vector Machine (SVM) (Cortes and Vapnik, 1995) on this space to predict which head produced each caption, using a $70/30$ train-test split on captions.
+In the French/English experiment, the SVM achieves $99.8\%$ test accuracy for the two-head LoRA-MCL model of Section 6, confirming specialization.
+Embeddings of the hypotheses are visualized using a Principal Components Analysis (PCA) with two components in Figure 6.
+Outside this controlled setup, we repeated this analysis on captions from the LoRA-MCL trained on AudioCaps with $K=5$.
+The SVM reaches $63.7\%$ test-accuracy, over three times better than random choice, indicating a clear specialization of the heads.
+In contrast, applying the same procedure to captions from the LoRA-MLE baseline (with DBS decoding, $\lambda=0.8$, and beam size = 5) yielded $23.2\%$ accuracy on the test set after SVM fitting, close to random choice and showing no evidence of specialization, likewise for LoRA-MoE (yielding $19.8\%$ accuracy in the same setup).
+We believe that further understanding of how specialization emerges represents a promising direction for future work.
+E.7.5 Qualitative Examples
+In this section, we show some qualitative examples (image–predicted caption pairs) that highlight the behavior of LoRA-MCL compared to a baseline model (LoRA-MLE with $r=24$).
+When the scene contains numerous objects or pieces of information, LoRA-MCL (through its different heads) covers a wider variety of descriptions than diverse beam search decoding.
+We performed inference with LoRA-MCL using greedy decoding, while we used diverse beam search ($\lambda=0.8$ and using $3$ beams) for the baseline model.
+LoRA-MLE. LoRA-MCL.
+{A display of Hugo Victor products on a table.} {A display of Hugo Victor products is set up in a store.}
+{A display for Hug Victor with a sign that says Hug Victor.} {A store display for Hugo Victor with a shelf full of items.}
+{a store display for Hugo Victor with many items} {A store display for Hugo Victor features a variety of plates, bowls, and books.}
+LoRA-MLE. LoRA-MCL.
+{a person getting food from a vending machine with the number 680 on it} {A girl is putting food into a vending machine that says FEEDO.}
+{A woman is getting food from a vending machine that says "Fresh" on it.} {A woman is putting money into a vending machine that says 100.}
+{A girl puts a sandwich into a vending machine that says "Fresh" on it.} {A girl is getting a hot dog from a vending machine.}
+LoRA-MLE. LoRA-MCL.
+{ A Scotch brand glue stick is on a table.} {A bottle of Scotch tape sits on a table.}
+{A can of Scotch branded glue is on a table.} {A bottle of Scotch tape sits on a table next to a piece of paper.}
+{a scotch tape that is on the ground} {A Scotch tape is on the table next to a piece of paper.}
+E.8 Diverse Machine Translation
+E.8.1 Experimental setup
+We fine-tuned ALMA-7B (based on LLaMA-2-7B (Touvron et al., 2023)) for two epochs on the human-written data collected by Xu et al.
+(2024a), which includes the WMT’17–WMT’20 test sets and the development and test sets from Flores-200 (Costa-Jussà et al., 2022).
+Our fine-tuning setup follows Xu et al.
+(2024a), except that we employ bf16 precision instead of fp16.
+We use $r=\tfrac{\alpha}{2}=16$, a maximum sequence length of $512$, an effective batch size of $256$, a warm-up ratio of $0.01$ and a peak learning rate of $2\times 10^{-5}$ with an inverse square-root scheduler.
+We follow the quality–diversity evaluation protocol of Shen et al.
+(2019), computing BLEU scores with the SacreBLEU (Post, 2018) library, which uses the commonly used Moses 13a Tokenizer (standard in Machine Translation) instead of the Penn Treebank Tokenizer we used for Audio and Image Captioning.
+The reported metrics are Leave-One-Out BLEU (Loo-BLEU) and Pairwise-BLEU (see Section 4 of Shen et al.
+(2019)), using the official codebase.
+Leave-One-Out BLEU (denoted simply as BLEU in (Shen et al., 2019)) measures the overall quality of the hypothesis set by computing the corpus-level BLEU for each hypothesis against the reference set (higher is better).
+Pairwise-BLEU considers only the predicted hypotheses, computing BLEU scores for each pair of candidates (lower indicates greater diversity).
+We define hereafter formally these two metrics, as done in (Shen et al., 2019).
+Let $\operatorname{BLEU}\left\{\left(\left[x^{1},\cdots,x^{R}\right],\hat{x}^{k}\right)\right\}_{\{k\in[\![1,K]\!]\}}$ denote the corpus-level BLEU when hypothesis $\hat{x}^{k}$ is evaluated against references $x_{1},\cdots,x_{R}$ as references for each example in the test-set and for each $k\in[\![1,K]\!]$.
+We also define $x^{-r}]=\{x^{1},\dots,x^{r-1},x^{r+1},\dots,x^{R}\}$.
+Leave-One-Out BLEU. Corresponds to an average of each corpus-level BLEU when using hypothesis $k$ against all possible sets of M1M-1 references
+$\mathrm{Loo{-}BLEU}=\frac{1}{R}\sum_{r=1}^{R}\operatorname{BLEU}\left\{\left(\left[x^{-r}\right],\hat{x}^{k}\right)\right\}_{\{k\in[\![1,K]\!]\}}\;.$
+Pairwise-BLEU. Similar to $\mathrm{mBLEU}_{4}$, Pairwise-BLEU considers only the set of candidates as
+$\mathrm{Pairwise{-}BLEU}=\operatorname{BLEU}\left\{\left(\left[\hat{x}^{j}\right],\hat{x}^{k}\right)\right\}_{\{j\in[\![1,K]\!],k\in[\![1,K]\!],j\neq k\}}\;.$
+We used it instead of $\mathrm{mBLEU}_{4}$ for consistency with the machine translation community.
+E.8.2 Qualitative Examples
+We provide some qualitative examples of the predictions on newstest2014. Here, LoRA-MCL uses $\varepsilon=0.05$, $B=3$ and $K=3$, and LoRA-MLE ($r=48$, DBS with $B=6$ and $\lambda=0.8$).
+[…]
+  • Wie Reuters im Juli erstmals berichtete, ist das Sitzlayout die treibende Kraft hinter der Auseinandersetzung um die neuen Jets.
+  • Wie Reuters im Juli erstmals berichtete, ist die Sitzanordnung genau das, was den Kampf zwischen den neuesten Jets antreibt.
+  • Wie Reuters erstmals im Juli berichtete, ist die Sitzanordnung genau das, was den Kampf zwischen den neuesten Jets antreibt.
+  • Wie Reuters im Juli berichtete, ist es genau das Thema der Sitzanordnung, das den Kampf zwischen den neuesten Flugzeugmodellen antreibt.
+  • Wie Reuters erstmals im Juli berichtet hatte, ist das Sitzkonzept genau der Punkt, der das Ringen zwischen den neuesten Jets beleb.
+  • Wie Reuters bereits erstmals im Juli berichtete, ist es genau das Layout der Sitze, was den Konkurrenzkampf zwischen den neuesten Jets anschürt.
+  • Wie Reuters erstmals im Juli berichtete, ist es das Sitz @-@ Layout, das den Wettbewerb zwischen den aktuellen Flugzeugen antreib.
+  • Wie Reuters zuerst im Januar berichtete, ist der Aufbau der Sitzreihen genau das, was den Kampf zwischen den neuesten Jets antreibt.
+LoRA-MLE.
+Wie Reuters zuerst im Juli berichtete, ist die Sitzanordnung genau das, was den Kampf zwischen den neuesten Jets antreibt .}\}
+[…]
+[…]
+LoRA-MCL.
+Als Reuters im Juli berichtete, ist die Sitzanordnung genau das, was den Kampf zwischen den neuesten Jets ausmacht.}\}
+[…]
+[…]
+Example 2. Input english sentence: $\{$“I was vocal about looking into a whole bunch of things, and this was one of them,”Daley told the paper.}\} References:
+  • “Ich äußerte eine ganze Reihe von Dingen, die man in Erwägung ziehen sollte, und das war eines davon”, erklärte Daley dem Magazin.
+Ich habe mich freimütig darüber geäußert eine ganze Reihe an Sachen ausprobieren zu wollen und das war eine davon, teilte Daley der Zeitung mit.
+  • “Ich war lautstark dabei, in eine ganze Reihe von Dingen zu schauen, und das war eines von ihnen”, Daley sagte der Zeitung.
+  • “Ich äußerte mich lautstark darüber, dass ich mir eine ganze Reihe von Dingen ansehen wollte. Und das war eines davon", sagte Daley der Zeitung..
+  • “Ich verkündete lautstark, mir verschiedene Dinge ansehen zu wollen, und dies war eines davon”Daley hat der Zeitung davon erzählt.
+  • “Ich war entschlossen, mich mit einer ganzen Reihe von Dingen zu befassen, und dies war eins davon”, sagte Daley der Zeitung.
+  • “Ich habe mündlich geäußert, mich mit einer ganzen Menge Dinge befasst zu haben und das hier gehörte dazu”, erzählte Daley der Zeitung.
+  • “Ich habe es deutlich gemacht, dass ich eine ganze Reihe von Sachen unter die Lupe nehmen werde, und das war eine von ihnen”, erzählte Daley der Zeitung.
+  • “Ich habe mich dazu geäußert, mir mehrere Dinge ansehen zu wollen, und das war eines davon”, sagte Daley der Zeitung.
+  • “Ich hatte geäußert, eine ganze Reihe an Dingen zu untersuchen, und dies war eine von ihnen”, sagte Daley der Zeitung.
+LoRA-MLE.
+„Ich war laut darüber, dass wir uns mit einer ganzen Reihe von Dingen beschäftigen sollten, und dies war eine davon“, sagte Daley der Zeitung.}\}
+$\{$„Ich war laut darüber, dass wir uns mit einer ganzen Reihe von Dingen beschäftigen sollten, und das war eine davon“, sagte Daley der Zeitung.}\}
+$\{$“Ich habe mich laut darüber geäußert, dass ich mich in eine ganze Reihe von Dingen einlesen wollte, und das war eine davon”, sagte Daley der Zeitung.}\}
+LoRA-MCL.
+“Ich war laut darüber, dass ich mich in eine ganze Reihe von Dingen einlesen wollte, und das war eine davon,” Daley sagte der Zeitung.}\}
+$\{$“Ich war laut über die Untersuchung einer ganzen Reihe von Dingen, und dies war eines von ihnen”, sagte Daley der Zeitung.}\}
+[…]
+E.9 Computation details
 We run the experiments mostly on H100 NVIDIA GPUs with $80$ GB of RAM.
-Trainings and inferences were launched on a single GPU for the Audio Captioning experiments, and up to $8$ GPUs for the Image Captioning experiments.
-The total computing resources used for this project, including failed experiments, amount to approximately $20{,}000$ GPU hours.
+Training and inference were launched on a single GPU for the Audio Captioning experiments, and up to $8$ GPUs for the Image Captioning experiments.
+The total computing resources used for this project, including failed experiments, amount to approximately 25,00025{,}000 GPU hours.
+E.10 Use of Large Language Models
+We used LLMs assistants in the preparation of this work.
+They helped to polish the writing (improving clarity, grammar, and style without altering the content) and serving as a coding assistant (visualization, debugging).